### v1

# Zilber’s Trichotomy in Hausdorff Geometric Structures

Abstract.
We give a new axiomatic treatment of the Zilber trichotomy, and use it to complete the proof of the trichotomy for relics of algebraically closed fields, i.e., reducts of the ACF-induced structure on ACF-definable sets. More precisely, we introduce a class of geometric structures equipped with a Hausdorff topology, called Hausdorff geometric structures. Natural examples include the complex field; algebraically closed valued fields; o-minimal expansions of real closed fields; and characteristic zero Henselian fields (in particular $p$-adically closed fields). We then study the Zilber trichotomy for relics of Hausdorff geometric structures, showing that under additional assumptions, every non-locally modular strongly minimal relic on a real sort interprets a one-dimensional group. Combined with recent results, this allows us to prove the trichotomy for strongly minimal relics on the real sorts of algebraically closed valued fields. Finally, we make progress on the imaginary sorts, reducing the trichotomy for all ACVF relics (in all sorts) to a conjectural technical condition that we prove in characteristic $(0,0)$.
The first author was partially supported by the Fields Institute for Research in Mathematical Sciences; NSF grant DMS 1800692; ISF grant No. 555/21; a BGU Kreitman foundation fellowship; and a UMD Brin postdoc. The second author was supported by ISF grants No. 555/21. The third author was partially supported by the Fondation Sciences Mathématiques de Paris
1. Introduction
The main goal of this paper was to complete the proof of Zilber’s Restricted Trichotomy Conjecture for algebraically closed fields. In characteristic 00 this conjecture was proven by the first author in [Cas23]. In order to adapt key analytic ideas from [Cas23] to positive characteristic, we opted, rather than working with formal schemes (as in [HS24]), to work in algebraically closed valued fields (ACVF). In this setting – in complete models – the theory of analytic functions provides suitable analogues of the characteristic 0 statements we need. Our main result is, thus:
Theorem 1.
Let $\mathcal{K}$ be an algebraically closed valued field (ACVF). Let $\mathcal{M}$ be a definable strongly minimal $\mathcal{K}$-relic. If $\mathcal{K}$ is not locally modular then $\mathcal{K}$ interprets a field $\mathcal{K}$-definably isomorphic to $K$.
Let us briefly explain the terminologies used above. Recall that a $\mathcal{K}$-relic (see [CH24]) is a structure $\mathcal{M}=(M,...)$ such that the universe $M$, and all $\mathcal{M}$-definable subsets of powers of $M$, are interpretable sets in $\mathcal{K}$. If $M\subset K^{n}$ for some $n$, we call $\mathcal{M}$ a definable $\mathcal{K}$-relic.
A structure $\mathcal{M}=(M,...)$ is strongly minimal if every definable subset of $M$ is either finite or co-finite (uniformly in definable families). Non-local modularity is a necessary non-triviality condition (explained below) for $\mathcal{M}$ to interpret a field.
By elimination of imaginaries, every ACF-relic is isomorphic to a definable ACF-relic. Moreover, clearly, every definable ACF-relic is also a definable ACVF-relic. In particular, as a special case, we obtain Zilber’s restricted trichotomy for ACF:
Corollary.
Let $\mathcal{K}$ be an algebraically closed field. Any strongly minimal $\mathcal{K}$-relic is either locally modular or interprets a field $\mathcal{K}$-definably isomorphic to $K$. In particular, an arbitrary $\mathcal{K}$-relic is either 1-based or interprets such a field.
Following Hrushovski’s refutation [Hru93] of Zilber’s full trichotomy conjecture (suggesting that every non-locally modular strongly minimal structure interprets an algebraically closed field), and the seminal work of Hrushovski and Zilber [HZ96] (proving the conjecture in the abstract setting of Zariski Geometries), the conjecture’s status became somewhat unclear. On the one hand, no alternative formulation emerged that could withstand Hrushovski’s technique for constructing counter-examples. On the other hand, Hrushovski’s applications of special instances of the conjecture (ultimately, based on variants of [HZ96]) in the solution of Diophantine problems (see [Bou98] for a survey of those) singled out Zilber’s trichotomy as a powerful principle, with a tendency to hold in geometric settings.
Lacking a unifying conjecture, research focused on restricting attention to strongly minimal relics of various theories of a geometric nature. In the late 1980s, Zilber and Rabinovich studied ACF-relics (see [Rab93] and references therein), subsuming works of Martin [Mar88] and Marker-Pillay [MP90]. In 2006 Peterzil suggested an analogous conjecture for o-minimal relics, and in [KR16] Kowalski and Randriambololona conjectured an analogue for ACVF.
The present work contributes to all of these conjectures. It gives a positive solution to the conjecture restricted to definable ACVF-relics, and thus to ACF-relics. In residue characteristic 0, the result also covers general (i.e. interpretable) ACVF-relics. Moreover, in a forthcoming paper, the first author uses the results of the present work to solve Peterzil’s conjecture for high-dimensional relics (which, combined with the main result of [HS24], gives Zilber’s trichotomy restricted to strongly minimal relics of Compact Complex Manifolds – a result that may be viewed as generalizing Moosa’s non-standard Riemann existence theorem, [Moo04]).
As the proof of Theorem 1 emerged, it became clear that the arguments could be written in greater generality. This led us to introduce an axiomatic framework – which we call Hausdorff geometric structures – and which allows for uniform treatments of crucial steps in various proofs of instances of the trichotomy. Our main abstract result (Theorem 8.1) states:
Theorem 2.
Let $(\mathcal{K},\tau)$ be a Hausdorff geometric structure. Assume that $(\mathcal{K},\tau)$ has ramification purity, has definable slopes satisfying TIMI, and is either differentiable or has the open mapping property. Let $\mathcal{M}=(M,...)$ be a non-locally modular strongly minimal definable $\mathcal{K}$-relic. Then $\dim(M)=1$, and $\mathcal{M}$ interprets a strongly minimal group.
In the next section, we provide a more detailed informal overview of the statement of this result and the general strategy of proof.
Theorem 2 relates to a recent work of Onshuus, Pinzon and the second author, [HOP24], where the problem of interpreting a field in definable ACVF-relics reduces to showing that the relic $\mathcal{M}$ interprets a group locally isomorphic to a local subgroup of either $(K,+)$ or $(K^{*},\cdot)$. As this can only happen if $M$ is (as a $\mathcal{K}$-definable set) 1-dimensional, the remaining problem splits in two: (1) showing that non-local modularity of $\mathcal{M}$ implies that $\dim(M)=1$, and then (2) showing that strongly minimal, non-locally modular, 1-dimensional relics interpret a group with the desired properties. We expect that (at least in characteristic 0) the axiomatic framework of Theorem 2 may suffice (with minor adjustments) to produce a field directly, without going through a group first. However, as this is not needed for our main results, and in order to not further lengthen the paper, we have not extended our axiomatization to cover this part of the argument.
The assumptions of Theorem 2 are, to a large extent, formulated to capture relics of analytic expansions of algebraically closed fields, and we do not expect them to apply in full in significantly different settings. However, as we will see below, our proof is built from the bottom up, strengthening the results as we specialize the axiomatization. This allows us to capture important steps of the strategy of proof in settings where only some of the axioms hold.
1.1. Strategy of proof
The proofs of our main theorems are split into three uneven parts. Sections 3 through 8 introduce the axiomatic framework and culminate with the proof of Theorem 2. In Sections 9 and 10, we verify that ACVF satisfies the assumptions of Theorem 2, and thus (in combination with the main result of [HOP24]) deduce Theorem 1. Section 12 is independent of the rest of the paper. It uses techniques introduced in [HHP22] to show that interpretable non-locally modular strongly minimal ACVF0,0-relics definably embed into a power of the valued field sort or the residue field sort (and thus are covered by Theorem 1 and Corollary Corollary).
Let us start by describing our axiomatic setting, and briefly explain the role of each set of axioms in the proof. The most general framework we work in, Hausdorff geometric structures, is introduced in Section 3. These are geometric structures (in the sense of [HP94, §2]) equipped with a Hausdorff topology. The axiomatization is aimed to capture the tameness of the topology (in the sense of e.g. [vdD98], [SW19] or [DG22]), without assuming its definability. Roughly, it assures a good interaction between the dimension theory of definable sets and the topology.
The main technical novelty in this section is the notion of enough open maps – a condition inspired by the most critical use of the complex analytic topology in [Cas23], as well as one of the critical uses of o-minimal geometry in the o-minimal variant [EHP21] of the restricted trichotomy. It aims to state, without assuming any differential structure, that the intersection of definable curves should be transverse, unless an obvious obstruction prevents it. It is designed to be a common generalization of this phenomenon in both the o-minimal and analytic settings, and as such it is rather technical; but the main point is that certain finite-to-one definable maps are stipulated to be open.
As the notion of enough open maps is rather technical and unpleasant to verify directly in applications, we give two simpler notions, each of which implies enough open maps. These are inspired by the examples above – namely, one (the open mapping property) is an abstraction of the complex algebraic setting, and the other (differentiability) is an abstraction of the o-minimal setting.
Section 4 begins the study of strongly minimal relics of Hausdorff geometric structures (with enough open maps). At this level of generality (covering such contexts as o-minimal strongly minimal relics, 1-h-minimal strongly minimal relics and more) we can define and prove a key topological property, that we call weak detection of closures. Detection of closures, as a key tool for the study of strongly minimal relics, goes back to [PS01], and plays a crucial role in several subsequent works on various restrictions of Zilber’s trichotomy ([HK08], [EHP21], and most importantly in [Cas23]). In Section 4 we define this notion (Definition 4.8), and in the subsequent section we prove one of the main results of this paper:
Theorem 3.
Let $(\mathcal{K},\tau)$ be a Hausdorff geometric structure with enough open maps. Let $\mathcal{M}$ be a non-locally modular strongly minimal definable $\mathcal{K}$-relic. Then, potentially after naming a small set of parameters, $\mathcal{M}$ weakly detects closures.
As was noted in [Cas23], at this level of generality, it seems somewhat over-optimistic to hope for the relic to detect all closure points of definable sets (in the sense, e.g., that if $X$ is $A$-definable in the relic and $x$ is in the frontier of $X$, then the Morley rank of $x$ over $A$ is less than that of $X$). The key idea, adapted from [Cas23] to our axiomatic setting, is to identify general enough situations when our over-optimistic hopes are in fact fulfilled. The proof of this result follows rather closely the analogous result from [Cas23], but has a few differences that we point out as we go.
In Section 6 we give applications of the above theorem to the Zilber trichotomy. We first show that from weak detection of closures one can detect certain double intersections of plane curves (see Definition 6.9 for details):
Theorem 4.
Let $(\mathcal{K},\tau)$ be a Hausdorff geometric structure. Let $\mathcal{M}$ be a non-locally modular strongly minimal definable $\mathcal{K}$-relic. If $\mathcal{M}$ weakly detects closures, then $\mathcal{M}$ detects multiple intersections.
We then conclude that, provided $(\mathcal{K},\tau)$ obeys a suitable form of the purity of the ramification locus, non-locally modular strongly minimal definable relics must have one-dimensional universes (see Definition 6.16 for the notion of ramification purity):
Theorem 5.
Let $(\mathcal{K},\tau)$ be a Hausdorff geometric structure. Let $\mathcal{M}=(M,...)$ be a non-locally modular strongly minimal definable $\mathcal{K}$-relic. Assume that $\mathcal{M}$ detects multiple intersections, and $(\mathcal{K},\tau)$ has ramification purity. Then $\dim(M)=1$ in the sense of $\mathcal{K}$.
We expect ramification purity to hold only in (expansions of) algebraically closed fields. In (o-minimal) expansions of real closed fields the first author proves a variant, sufficient to obtain an analogue of the above theorem, but we do not expect the same approach to apply in other settings. Of course, we do not expect non-locally modular strongly minimal relics to be definable in tame expansions of fields that are not real closed or algebraically closed – so one could hope for a uniform, dimension-independent, proof of such a result.
Section 7 axiomatizes the interpretation of a group when $\dim(M)=1$. We introduce a setting of definable slopes – requiring that curves in $K^{2}$ can be approximated near generic points by abstract ‘Taylor polynomials.’ We then introduce a technical notion called TIMI (tangent intersections are multiple intersections) – stipulating that if two curves share the same Taylor polynomial to unusually high order at a point, the point forms a ‘topological multiple intersection’ – which allows us to prove Theorem 2. This axiom has a distinct analytic flavour. It would be interesting to find a weaker axiom implying TIMI for curves definable in strongly minimal relics (this was one of the main technical results of [EHP21]).
In Section 9 we give examples of Hausdorff geometric structures with enough open maps. In particular, we note that certain visceral theories (in the sense of [DG22]) share these properties (covering 1-h-minimal valued fields and o-minimal expansions of fields) and that a certain class of éz topological fields (essentially coming from the notion of éz fields in [WY23]) also satisfies these axioms. This is our main example, as it is the only one covering ACVF in all characteristics.
Restricting further to algebraically closed éz fields, we next verify the axiom on the purity of the ramification locus. Then, for the rest of the paper, we work concretely in ACVF. First, in section 10, we show that ACVF has definable slopes as well as the remaining axiom TIMI. The main novelty in this section is the introduction of Taylor groupoids, allowing for greater flexibility in the study of slopes in definable families of curves, compared with earlier treatments of similar problems. Also, the definition of slope in positive characteristic is non-standard, absorbing powers of the Frobenius automorphism into the definition, allowing for a more uniform and less technical treatment of slopes.
The final section of the paper uses techniques from [HHP22] to extend the results of the main theorem to interpretable relics. This section is independent of the rest of the paper.
It seems that, at least in characteristic 00, large parts of the present argument could be extended to analytic expansions of ACVF (in the sense of Cluckers and Lipshitz [CL11]). Extending the results to arbitrary $V$-minimal theories may prove more challenging. More generally, in a recent preprint, [Joh24a], Johnson shows that $C$-minimal expansions of ACVF are geometric, implying that they are also Hausdorff geometric structures. As Johnson proves, moreover, that definable functions are generically strictly differentiable, it seems that the proof of enough open maps for 1-h-minimal fields could be extended to the $C$-minimal setting, but there are details to verify.
2. Preliminaries
We give a quick overview of some of the model-theoretic notions used extensively in the paper. Since modern detailed reviews of the relevant notions can be found in recent papers, we will be brief, directing interested readers to relevant references.
Throughout the paper, structures are denoted by calligraphic letters $\mathcal{K},\mathcal{M}$ etc. and their universes by the corresponding Roman letters $K,M$ etc. Roman letters $A,B$ will usually denote sets of parameters (i.e., subsets of the universe of the structure). Whenever the structure is assumed to be $\lambda$-saturated for some cardinal $\lambda$, all parameter sets are tacitly assumed to be of cardinality smaller than $\lambda$. Lowercase Latin letters $a,b,c$ usually denote finite tuples of elements in our structure. When no confusion can occur, we write $AB$ and $Aa$ etc. as a shorthand for $A\cup B$ and $A\cup\mathrm{dom}(a)$. Definable sets are usually denoted by $X,Y,Z$ etc. Unless specifically stated otherwise, the word definable allows parameters but does not allow imaginary sorts. If we wish to allow imaginary sorts, we either refer to interpretable sets, or we write explicitly definable in $T^{eq}$, definable in $\mathcal{K}^{eq}$ as the case may be.
We use standard model-theoretic notation and terminology. Readers are referred to any textbook in model theory such as [Maa15] or [TZ12] for further details.
Throughout the paper, we often work with finite correspondences between interpretable sets. There are competing versions of what this might mean – so before moving on, we clarify:
Definition 2.1.
Let $X$ and $Y$ be interpretable sets in some structure. A finite correspondence between $X$ and $Y$ is an interpretable set $Z\subset X\times Y$ whose projections to $X$ and $Y$ are finite-to-one and surjective. If there is a finite correspondence between $X$ and $Y$, we say that $X$ and $Y$ are in finite correspondence. If $X$ is in finite correspondence with a subset of $Y$, we say that $X$ is almost embeddable into $Y$.
2.1. Geometric Structures
Throughout this paper we will be working with geometric structures:
Definition 2.2.
We use $\operatorname{acl}$ to denote the model-theoretic algebraic closure. Let $T$ be a complete first-order theory. We say $T$ is geometric if:
  • $\operatorname{acl}$ satisfies exchange. More explicitly, for any $\mathcal{K}\models T$, $A\subseteq K$, and $b,c\in K$, $c\in\operatorname{acl}(Ab)\setminus\operatorname{acl}(A)$ implies that $b\in\operatorname{acl}(Ac)$.
  • $T$ eliminates $\exists^{\infty}$.
A structure $\mathcal{K}$ is geometric if its theory is geometric.
In the definition, by elimination of $\exists^{\infty}$ (also referred to as uniform finiteness in the literature) we mean that for any model $\mathcal{K}$ and any formula $\phi(x,\overline{y})$ (where $|x|=1$ and $|\overline{y}|=n$, say), the set $\{b\in K^{n}:|\phi(K,b)|<\infty\}$ is definable.
Working in a $|T|^{+}$-saturated geometric $\mathcal{K}$, there is a well-established theory of dimension and independence on $\mathcal{K}$, defined as follows. Let $A\subset K$ be small, let $a\in K^{m}$, and let $X$ be $A$-definable.
  • We $\dim(a/A)$ to be the length of a maximal $\operatorname{acl}_{A}$-independent subtuple of $a$. By the exchange property, this is independent of the choice of the maximal independent subtuple.
  • We define $\dim(X)=\max\{\dim(x/A):x\in X\}$. By a compactness argument, $\dim(X)$ does not depend on the parameter set $A$.
  • We say that $a$ is independent from $b$ over $A$ if $\dim(a/Ab)=\dim(a/A)$.
In this language, the exchange property can be equivalently written as the additivity formula:
$\dim(ab/A)=\dim(a/A)+\dim(b/Aa)$
for any $a\in K^{m}$, $b\in K^{n}$, and $A\subset\mathcal{K}$. This is, arguably, the single most useful property of dimension in geometric structures. Among other things, additivity gives the following:
  • Independence is symmetric: $a$ is independent from $b$ over $A$, if and only if $b$ is independent from $a$ over $A$, if and only if $\dim(ab/A)=\dim(a/A)+\dim(b/A)$ (thus, we freely use the ambiguous language ‘$a$ and $b$ are independent over $A$’).
  • Independence is transitive: $a$ is independent from $bc$ over $A$ if and only if $a$ is independent both from $b$ over $A$ and from $c$ over $Ab$.
  • Independence extends naturally to several tuples: $a_{1},...,a_{n}$ are independent over $A$ if
    $\dim(a_{1}...a_{n}/A)=\sum_{i}\dim(a_{i}/A),$
    if and only if each $a_{i}$ is independent from $\{a_{j}:j\neq i\}$ over $A$.
Using compactness, one can also characterize dimension as follows: $\dim(X)\geq n$ if and only if $\dim(\pi(X))=n$ for some coordinate projection $\pi:X\to K^{n}$, and $\dim(X)$ is the maximal such $n$; or, $\dim(X)\leq n$ if and only if there is a definable finite-to-one map $f:X\rightarrow K^{n}$, and $\dim(X)$ is the minimal such $n$. In fact, the finite-to-one map $f:X\rightarrow K^{\dim X}$ can be defined uniformly in families (by coordinate projections); combined with uniform finiteness, one concludes that dimension is definable in families: given a formula $\phi(x,y)$, the set $\{y:\dim(\phi(x,y)=d\}$ is definable for any integer $d$. We will use these facts without further mention.
Many other basic properties follow easily from the definition and saturation. For example:
  • $\dim(X\times Y)=\dim(X)+\dim(Y)$.
  • $\dim(X\cup Y)=\max\{\dim(X),\dim(Y)\}$.
  • If $X$ and $Y$ are in finite correspondence, then $\dim(X)=\dim(Y)$.
We will use all of these facts without further mention for the rest of the paper.
For an $A$-definable set $X$ and $B\supseteq A$, we say that $x\in X$ is a generic point of $X$ over $B$ if $\dim(x/B)=\dim X$. A definable set $Y\subseteq X$ is generic in $X$ if $\dim(X)=\dim(Y)$, if and only if $Y$ contains some $y$ generic in $X$ (over some parameter set over which $Y$ is defined). Two definable sets $X$ and $Y$ are almost equal if the symmetric difference $X\Delta Y$ is not generic in either of them (equivalently, by the above, if $X$ and $Y$ have the same generic points over some parameter set defining both of them). It is easy to see that almost equality is an equivalence relation on definable sets, and is definable in families (because dimension is).
Finally, we briefly note that dimension and independence can be extended to interpretable sets in a way that preserves additivity (but may lose the equivalence with acl-independence for interpretable sets that are not definable) – see section 3 of [Gag05]. Thus, the notation $\dim(a/A)$ is well-defined even for $a\in\mathcal{K}^{eq}$, and one still has all consequences of additivity. We will use this, but very rarely (in fact, only in Section 7, when computing dimensions of slopes and their coherent representatives).
For more details on the basic properties of dimension in geometric structures, we direct the reader to e.g. [AH24, §2].
2.2. Strongly Minimal Structures
In the paper, we study strongly minimal relics of certain geometric structures. Recall that a structure is strongly minimal if, in all of its elementary extensions, every definable set in one variable is finite or cofinite. Strongly minimal structures are themselves geometric structures. We now recall some facts and terminology specific to the strongly minimal case. First, the dimension associated with strongly minimal structures is known as Morley Rank, and in this paper will be called rank and denoted $\operatorname{rk}$.
The main advantage of strongly minimal structures (viewed within the larger class of geometric structures) is the notion of stationarity. In the context of geometric structures, a definable set is stationary if it cannot be partitioned into two generic definable subsets. A key feature of stationary sets is that their almost equality classes can be coded: that is, if $X$ is stationary, there is $c\in\mathcal{M}^{eq}$ such that for all $\sigma\in Aut(\mathcal{M})$, $\sigma(c)=c$ if and only if $\sigma(X)$ is almost equal to $X$. This tuple $c$ is unique up to interdefinability, and is called the canonical base of $X$ and denoted $\operatorname{Cb}(X)$. (See [Mar02, §8.2] for more on canonical bases and [Cas23, §2.3] for some relevant applications).
In many geometric structures, only singletons are stationary. Notably, strongly minimal structures are exactly those geometric structures $\mathcal{M}=(M,...)$ such that the universe $M$ is stationary. In fact, it follows that $M^{n}$ is also stationary for each $n$, and that every definable subset $X\subset M^{n}$ is a disjoint union of finitely many stationary definable sets $Y_{i}$ such that $\operatorname{rk}(Y_{i})=\operatorname{rk}(X)$. These stationary sets are unique up to almost equality, and are called the stationary components of $X$.
In the context of a geometric structure, a curve is a definable set of dimension $1$. For a strongly minimal structure $\mathcal{M}=(M,...)$, we frequently use plane curve to describe curves in $M^{2}$. Note that this is not the same as an algebraic curve (even if $M$ is a field). To avoid confusion, we will clarify our interpretation of ‘curve’ when it is not clear from the context.
Following [Cas23], we call a plane curve $C\subset M^{2}$ non-trivial if both projections $C\rightarrow M$ are finite-to-one (equivalently, no stationary component of $C$ is almost equal to a horizontal or vertical line). By uniform finiteness, non-triviality is definable in families. We will largely be able to assume that all plane curves considered are non-trivial – but we will clarify this as we go.
We will frequently use intersections of relic-definable families of curves (as well as higher dimensional objects) to study an ambient topology from within the relic. Since, a priori, these curves are not geometric objects, we have to identify some combinatorial properties of families of curves allowing us to study them geometrically. This is discussed extensively in [Cas23, §2], so we will be brief.
Suppose $\mathcal{M}$ is strongly minimal. A parametrized family $\mathcal{X}=\{X_{t}:t\in T\}$ of subsets of $M^{n}$ is a definable family if the parameter set $T$ and the graph $X=\{(x,t):x\in X_{t},t\in T\}\subset M^{n}\times T$ are definable. If $\mathcal{X}=\{X_{t}:t\in T\}$ is a definable family, we call $\mathcal{X}$ almost faithful if
  1. (1)
    each fiber $X_{t}$ has the same dimension $d$,
  2. (2)
    for each $t$, there are only finitely many $t^{\prime}$ with $\dim(X_{t}\cap X_{t^{\prime}})=d$.
If $\mathcal{X}$ is almost faithful, the rank of $\mathcal{X}$ is the rank of the parameter set, $\operatorname{rk}(T)$. It is not hard to check that every stationary definable set $S$ is, up to almost equality, a generic member of an almost faithful definable family of rank $\operatorname{rk}(\operatorname{Cb}(S))$.
For the purposes of the present work, it is convenient to use the following definition:
Definition 2.3.
A strongly minimal structure is not locally modular if it admits a definable rank 2 almost faithful family of plane curves.
This is well known to be equivalent to the standard definition (see, e.g., [Mar02, Theorem 8.2.11]). Clearly, algebraically closed fields are not locally modular (as witnessed, e.g., by the family of affine lines). Moreover, it is not hard to see that local modularity is preserved under interpretations between strongly minimal structures. Thus, non-local modularity is a necessary condition for a strongly minimal structure to interpret such a field.
Finally, we recall some special types of families of plane curves that were used extensively in [Cas23]. These are thought of as families that have been presented most conveniently and had various types of irregularities removed.
Definition 2.4.
Let $\mathcal{C}=\{C_{t}:t\in T\}$ be a definable almost faithful family of plane curves in a strongly minimal structure $\mathcal{M}=(M,...)$, with graph $C\subset M^{2}\times T$.
  1. (1)
    $\mathcal{C}$ is a standard family if $T$ is a generic subset $M^{n}$ for some $n\geq 1$, and for each $p\in M^{2}$, the set $\{t\in T:p\in C_{t}\}$ is non-generic in $T$.
  2. (2)
    $\mathcal{C}$ is excellent if $T$ is a generic subset of $M^{2}$, each $C_{t}$ is non-trivial, and for each $p\in M^{2}$, the set $\{t\in T:p\in C_{t}\}$ is either empty or a non-trivial plane curve. (In particular, note that excellent families are also standard).
3. Hausdorff Geometric Structures and Enough Open Maps
[…]
3.1. The Definition
The following will be the basic framework we assume throughout:
Definition 3.1.
Let $(\mathcal{K},\tau,\mathcal{L})$ be as above. We say that $(\mathcal{K},\tau)$ is a Hausdorff geometric structure if the following hold:
  1. (1)
    $\tau$ is Hausdorff.
  2. (2)
    $\mathcal{K}$ is geometric and $\aleph_{1}$-saturated (but $\mathcal{L}$ might be uncountable).
  3. (3)
    (Strong Frontier Inequality) If $X\subset K^{n}$ is definable over $A$ and $a\in\overline{\operatorname{Fr}(X)}$ then $\dim(a/A)<\dim(X)$.
  4. (4)
    (Baire Category Axiom) Let $X\subset K^{n}$ be definable over a countable set $A$, and let $a\in X$ be generic over $A$. Let $B\supset A$ be countable. Then every neighborhood of $a$ contains a generic of $X$ over $B$.
  5. (5)
    (Generic Local Homeomorphism Property) Suppose $X$, $Y$, and $Z\subset X\times Y$ are definable over $A$ of the same dimension, and each of $Z\rightarrow X$ and $Z\rightarrow Y$ is finite-to-one. Let $(x,y)$ be generic in $Z$ over $A$. Then there are open neighborhoods $U$ of $x$ in $X$, and $V$ of $y$ in $Y$, such that the restriction of $Z$ to $U\times V$ is the graph of a homeomorphism $U\rightarrow V$.
Warning 3.2.
We caution that Definition 3.1 does not assume the existence of a definable basis for $\tau$ – or even a basis whole members are definable. Thus $\tau$ is attached to the specific structure $\mathcal{K}$, and need not induce a similar topology on any elementary extension of $\mathcal{K}$.
Indeed, we will work throughout in a fixed Hausdorff geometric structure $(\mathcal{K},\tau)$, without considering any other models of its theory. This is why we assume $\mathcal{K}$ is $\aleph_{1}$-saturated in Definition 3.1 – so that we still have some of the tools of saturated models at our disposal (e.g. generic points will always exist over countable sets). In particular, with almost no exceptions, we will not use monster models of $\operatorname{Th}(\mathcal{K})$ (and if we do use one, we will make clear what we are doing when it happens).
Our reason for making this choice (i.e. to not require a definable basis) is given by two of the motivating examples. First, in the complex numbers with the analytic topology (considered in [Cas23]), the only definable open sets are the Zariski open sets, which do not form a basis for the analytic topology. Second, in many éz fields (see below for a discussion of these fields), one has only an ind-definable basis (i.e. there is a basis that is a countable union of definable families of open sets). Thus, Definition 3.1 is designed to accommodate these settings. Note, however, that in all other examples we know, the topology does have a definable basis, and this issue is irrelevant.
Remark 3.3.
It is easy to see that Hausdorff geometric structures are preserved under naming a countable set of constants (and indeed the same will hold of all additional properties we discuss). This will be a useful observation for the following reason: suppose $\mathcal{K}$ is a Hausdorff geometric structure, and $\mathcal{M}=(M,...)$ is a relic of $\mathcal{K}$ whose language is countable. Then we can add to $\mathcal{L}$ all parameters needed to define $\mathcal{M}$ (i.e. the universe $M$ and all basic relations of $\mathcal{M}$). Thus, we may assume that every $\emptyset$-definable set in $\mathcal{M}$ is $\emptyset$-definable in $\mathcal{K}$.
We will see detailed accounts of various Hausdorff geometric structures in Section 9. For now, we list them without proof for the reader’s intuition. Key examples include the complex field with the analytic (i.e. Euclidean) topology; any $\aleph_{1}$-saturated algebraically closed valued field with the valuation topology; any $\aleph_{1}$-saturated 1-h-minimal valued field with the valuation topology; any $\aleph_{1}$-saturated characteristic zero Henselian field with the valuation topology (in particular $\aleph_{1}$-saturated $p$-adically closed fields); and any $\aleph_{1}$-saturated o-minimal expansion of a real closed field with the order topology.
Generalizing characteristic zero Henselian fields, one can also take a large class of éz fields. Recall (see [JTWY23]) that the étale open topology on a large field $K$ is a system of topologies on the sets $V(K)$ for all varieties $V$ over $K$, where basic open sets are taken to be images of étale morphisms. A pure field is éz if every definable set is a finite union of definable étale open subsets of Zariski closed sets (see [WY23]). It is proved in [JTWY23] (Proposition 4.9) that the étale open topology of $K$ is induced by a field topology on $K$ if and only if it respects products of varieties in the obvious sense. Now suppose $K$ is an $\aleph_{1}$-saturated éz field whose étale open topology is induced by a field topology; we will see later that $K$ forms a Hausdorff geometric structure when equipped with the étale open topology on $K$.
Remark 3.4.
Note that it is still open whether many of our results can be proven for arbitrary éz fields (where the étale open topology might not respect products). This case could still be useful, as it could have implications for relics of pseudo algebraically closed fields (in particular pseudo-finite fields).
Assumption 3.5.
From now until the end of section 7, fix a Hausdorff geometric structure $(\mathcal{K},\tau)$. All sets $A,B,...$ of parameters are now assumed to be countable.
In a previous draft of this paper, we proved several basic properties of Hausdorff geometric structures that are not used in the sequel. Nevertheless, we list a few of them below, as they may give some intuition on the interaction between the topology and dimension theory. The interested reader may want to check these facts as an easy exercise.
Lemma 3.6.
Let $X\subset K^{n}$ be definable over $A$, and let $a\in X$ be generic over $A$.
  1. (1)
    $X$ is locally closed in a neighborhood of $a$ – that is, there is a neighborhood $U$ of $a$ in $K^{n}$ such that $U\cap X$ is relatively closed in $U$.
  2. (2)
    Any first-order property over $A$ which holds of $a$, holds for all points in some relative neighborhood of $a$ in $X$.
  3. (3)
    If $Y\subset X$ is definable and contains a neighborhood of $a$ in $X$, then $\dim(Y)=\dim(X)$.
  4. (4)
    In particular, no infinite definable subset of $K^{n}$ is discrete.
Before moving on, we make some further comments on Definition 3.1.
Remark 3.7.
The Baire category axiom should be viewed as a substitute for the more common statement ‘acl dimension equals topological dimension.’ In particular, this axiom is needed for Lemma 3.6(3)-(4). The reason for choosing our version of the axiom is that the latter becomes less useful when the topology is not definable.
Remark 3.8.
On a similar note to above, suppose $\tau$ has a definable basis. Then (modulo the generic local homeomorphism property and compactness) the Baire category axiom follows from the simpler statement ‘every definable open subset of $K^{n}$ has dimension $n$’. Moreover, the latter condition holds automatically, provided $K$ has no isolated points. Indeed, in a Hausdorff topology with no isolated points, all open sets are infinite, so – when definable in a geometric structure – one dimensional. Since the topology on $K^{n}$ is the product topology, and dimension in geometric structures is additive in Cartesian products, definable open sets contain $n$-dimensional boxes, which implies the desired conclusion.
On the other hand, in the complex field with the analytic topology, the Baire category axiom follows easily from Baire’s theorem – hence the name for the axiom.
Remark 3.9.
One typically thinks of the ‘frontier inequality’ as stipulating that $a\in\operatorname{Fr}(X)$ implies $\dim(a/A)<\dim(X)$. Our ‘strong’ version (condition (3) above) also applies to some elements of $X$; however, the two conditions are equivalent if the closure of a definable set is always definable. As written, the strong frontier inequality essentially amounts to ‘generic local closedness’ (see Lemma 3.6(2)).
3.2. Germs of Type-Definable Sets
In the next subsections, we will need to talk about the local behavior of definable sets and functions near a given point. An ideal way of doing this would be to use ‘infinitesimal neighborhoods’. If the topology $\tau$ is definable, one can generate an infinitesimal neighborhood of a point $x$ by intersecting infinitely many definable open neighborhoods of $x$. The result is a type-definable set. An example of particular interest is the set of realizations of $\operatorname{tp}(x/A)$ for a small set $A$ (indeed, it follows by the frontier inequality that every $\pi(x,a)\in\operatorname{tp}(x/A)$ holds on a neighborhood of $x$).
In our case, the topology might not be definable; however, certain type-definable sets – including complete types as above – still give some intuition of infinitesimal neighborhoods (essentially, because the frontier inequality still applies as above). Our goal now is to develop a general notion of germed type-definable sets which captures this idea. For instance, the frontier inequality will imply that all complete types over small sets are germed at every point. Importantly, the class of germed type-definable sets is richer than that of complete types, as new examples can be generated by closing under intersections and products (Lemma 3.15). This allows us to talk about germs of certain more intricate configurations when defining the key property of enough open maps (Definition 3.25) – which ultimately makes this notion much easier to state.
To start, let us clarify what we mean by type-definable sets in this paper:
Definition 3.10.
Let $X\subset K^{n}$, and let $A$ be a countable parameter set.
  1. (1)
    $X$ is type-definable over $A$ if $X$ is a (potentially infinite) intersection of $A$-definable sets.
  2. (2)
    In general, $X$ is type-definable if it is type-definable over some countable set $B$.
  3. (3)
    $X$ is a complete type over $A$ if $X$ is type-definable over $A$ and whenever $Y$ is definable over $A$, we either have $X\subset Y$ or $X\cap Y=\emptyset$.
  4. (4)
    If $X$ is type-definable, we call $X$ a complete type if it is a complete type over some countable set.
  5. (5)
    A type-definable function is a function $f:Y\rightarrow Z$ such that $Y$, $Z$, and the graph of $f$, are all type-definable (so we include $Y$ and $Z$ in the data of the function). Note that, by compactness, this is the same as the restriction of a definable function to a type-definable domain and target.
By compactness (and the fact that $(\mathcal{K},\tau)$ is $\aleph_{1}$-saturated), note that we have the following useful fact:
Fact 3.11.
Let $A$ be countable, and let $X=\bigcap\mathcal{Y}$ where $\mathcal{Y}$ is a collection of $A$-definable sets. Assume that $\mathcal{Y}$ is closed under finite intersections. Then $\mathcal{Y}$ is cofinal in the definable supersets of $X$. That is, if $Z\supset X$ is any definable set, then there is $Y\in\mathcal{Y}$ with $Z\supset Y$.
As is well-known in model theory, notions of dimension and genericity transfer naturally to type-definable sets:
Definition 3.12.
Let $X$ be type-definable over $A$. By $\dim(X)$, we mean any of the following values (which are all the same by compactness and Fact 3.11):
  • the smallest dimension of a definable set containing $X$
  • the smallest dimension of an $A$-definable set containing $X$
  • the largest value of $\dim(a/A)$ for $a\in X$.
If $a\in X$ with $\dim(a/A)=\dim(X)$, we say that $a$ is generic in $X$ over $A$.
Now let us discuss germs.
Definition 3.13.
Let $X$ be type-definable, and $a\in X$. We say that $X$ is germed at $a$ if there is a definable $Y\supset X$ such that, for all definable $Z$ with $X\subset Z\subset Y$, $Y$ and $Z$ agree in a neighborhood of $a$. If $f:X\rightarrow Y$ is a type-definable projection (i.e. a projection between type-definable sets), we say that $f$ is germed at $a$ if $X$ is germed at $a$ and $Y$ is germed at $f(a)$.
Example 3.14.
For intuition, we give an example of a type-definable set which is not germed. Suppose $(\mathcal{K},\tau)$ is $\mathbb{C}$ with the analytic topology. Now let $X\subset\mathbb{C}^{2}$ be the union of all lines $y=ax$ where $a\in\mathbb{C}-\mathbb{Q}$ – in other words, the complement of the non-zero parts of all lines through $(0,0)$ with rational slope. Clearly, $X$ is type-definable over $\emptyset$. But a definable approximation to $X$ will only remove finitely many lines with rational slope, and thus does not give the correct germ at $(0,0)$.
One now easily checks the following:
Lemma 3.15.
Let $X$, $Y$, and $Z$ be type-definable, $a\in X\cap Y$, and $b\in Z$.
  1. (1)
    If $X$ is definable, then $X$ is germed at $a$.
  2. (2)
    If $X$ is type-definable over $A$, and $a\in X$ is generic over $A$, then $X$ is germed at $a$.
  3. (3)
    If $X$ is a complete type, then $X$ is germed at $a$.
  4. (4)
    If $X$ is germed at $a$ and $Y$ is germed at $a$, then $X\cap Y$ is germed at $a$.
  5. (5)
    If $X$ is germed at $a$ and $Z$ is germed at $b$, then $X\times Z$ is germed at $(a,b)$.
Proof.
(1) is trivial. (2) uses the strong frontier inequality. (3) is an application of (2). (4) and (5) use Fact 3.11. ∎
Note that if $X$ is germed at $a$, then $X$ determines a unique germ of open neighborhoods of $a$. We call this the germ of $X$ at $a$ (here we treat germs as equivalence classes of sets, where two sets are equivalent if they agree on a neighborhood of $a$). We will use germs to talk about local properties of $X$. Let us be more precise:
Definition 3.16.
Let $f:X\rightarrow Y$ be a type-definable projection, $a\in X$, and suppose $f$ is germed at $a$.
  1. (1)
    A d-approximation of $X$ at $a$ is a definable set $X^{\prime}\supset X$ such that $\dim(X)=\dim(X^{\prime})$ and $X$ and $X^{\prime}$ have the same germ at $a$.
  2. (2)
    A d-approximation of $f$ at $a$ is a projection $f^{\prime}:X^{\prime}\rightarrow Y^{\prime}$, where $X^{\prime}$ is a d-approximation of $X$ at $a$ and $Y^{\prime}$ is a d-approximation of $Y$ at $f(a)$.
Definition 3.17.
Let $P$ be a property of a set with a distinguished point. We say that $P$ is d-local if whenever $X$ is a definable set, $a\in X$, and $X^{\prime}$ is a d-approximation of $X$ at $a$, then $P(X,a)$ holds if and only if $P(X^{\prime},a)$ holds.
We also make the analogous definition of $d$-local properties of projections (using Definition 3.16(2) instead of (1)).
The main point of d-local properties is that they extend naturally from definable sets to germed type-definable sets. Namely, the following is well-defined:
Definition 3.18.
Let $P$ by a d-local property of a set with a distinguished point. Let $X$ be a type-definable set, let $a\in X$, and assume that $X$ is germed at $a$. We say that $X$ satisfies $P$ near $a$ if some (equivalently any) d-approximation of $X$ satisfies $P$ with the distinguished point $a$.
As in Definition 3.17, we make the analogous definition for type-definable projections.
In particular, the following are clear:
Lemma 3.19.
For a projection $f:X\rightarrow Y$ and a point $a$, each of the following properties is d-local:
  1. (1)
    $f$ is finite-to-one on some neighborhood of $a$.
  2. (2)
    $f$ is locally open at $a$: there are neighborhoods $U$ of $a$ in $X$, and $V$ of $f(a)$ in $Y$, such that $f$ restricts to an open map $U\rightarrow V$.
Before moving on, we also show the following useful facts:
Lemma 3.20.
Suppose $X\subset K^{n}$ is definable over $A$, and $x\in X$. The following are equivalent:
  1. (1)
    $X$ is a d-approximation of $\operatorname{tp}(a/A)$ at $a$.
  2. (2)
    $a$ is generic in $X$ over $A$.
Proof.
Suppose (1) holds. Then by definition, $\dim(X)=\dim(a/A)$, so (2) holds.
Now suppose (2) holds. Then $\dim(X)=\dim(a/A)$. So to show that $X$ is a d-approximation of $\operatorname{tp}(a/A)$, it suffices to show that $X$ has the same germ as $\operatorname{tp}(a/A)$ at $a$. To do this, let $Y$ be any d-approximation of $\operatorname{tp}(a/A)$ at $a$. We want to show that $X$ and $Y$ agree in a neighborhood of $a$. Since $Y$ is a d-approximation of $\operatorname{tp}(a/A)$, some neighborhood of $a$ in $Y$ is contained in $X$. So we need to show the same statement with $X$ and $Y$ reversed.
Shrinking $Y$ if necessary, by Fact 3.11, we may assume $Y$ is definable over $A$. If no neighborhood of $a$ in $X$ is contained in $Y$, then $a$ is in the frontier of $Y-X$. But $Y-X$ is definable over $A$, so the strong frontier inequality gives
$\dim(a/A)<\dim(Y-X)\leq\dim(Y)=\dim(a/A),$
a contradiction. ∎
Lemma 3.21.
Suppose that $x\in K^{n}$ and $A\subset B$ are parameter sets. If $x$ is independent from $B$ over $A$, then $\operatorname{tp}(x/A)$ and $\operatorname{tp}(x/B)$ have the same germ at $x$.
Proof.
Let $X$ be a d-approximation of $\operatorname{tp}(x/A)$. By Lemma 3.20, $x$ is generic in $X$ over $A$. Since $x$ is independent from $B$ over $A$, $x$ is also generic in $X$ over $B$. So by Lemma 3.20 again, $X$ is a d-approximation of $\operatorname{tp}(x/B)$. Thus $\operatorname{tp}(x/A)$ and $\operatorname{tp}(x/B)$ have a common d-approximation at $x$, and so they have the same germ at $x$. ∎
3.3. Enough Open Maps
The main technical result of the paper concerns Hausdorff geometric structures satisfying an additional axiom about the openness of certain definable maps (called enough open maps). Because this axiom is rather technical to state, we give it a separate treatment. Let us start by giving an example to motivate the definition.
Later in the paper, we will show that non-locally modular definable strongly minimal $\mathcal{K}$-relics can reconstruct a fragment of the topology $\tau$, assuming only that $(\mathcal{K},\tau)$ has enough open maps. To sketch the setting, suppose $(\mathcal{K},\tau)$ is an algebraically closed valued field (so $\tau$ is the valuation topology), and $\mathcal{M}=(K,...)$ is a strongly minimal relic with universe $K$ (for example, $\mathcal{M}$ could be the pure field structure). Let $X\subset K^{2}$ be a plane curve with a frontier point at $(a,b)$. If $X$ happens to be definable in $\mathcal{M}$, say over a tuple $t$, we want to ‘recognize’ the frontier point $(a,b)$ using only the language of $\mathcal{M}$ (in an ideal world, this would mean showing $(a,b)\in\operatorname{acl}_{\mathcal{M}}(t)$). Our only tool for doing this is the existence of a Morley rank 2 family of plane curves in $\mathcal{M}$. For illustration, suppose this family is the family of all lines $y=\alpha x+\beta$ in $K^{2}$ (parametrized by $(\alpha,\beta)\in K^{2}$).
Now after potentially tweaking the setup (e.g. if addition is definable, we may want to translate $X$ by a generic), our strategy is as follows: we fix an independent generic slope $\alpha$, which determines a line $l$ of slope $\alpha$ through $(a,b)$. Suppose $l$ intersects $X$ in $m$ points. One can show, using the genericity of $\alpha$ (and the aforementioned ‘tweaks’), that $X$ and $l$ are not tangent at any of these $m$ points (that is, $X$ and $l$ are transverse at each point). It then follows that if we translate $l$ to a nearby line $l^{\prime}$ (still of slope $\alpha$), none of these $m$ points can ‘disappear’ – that is, $X\cap l^{\prime}$ will contain at least $m$ points, one near each point of $X\cap l$. Meanwhile, since $(a,b)\in\operatorname{Fr}(X)$, infinitely many such translations also produce an $(m+1)$-st point of $X\cap l^{\prime}$, this time near $(a,b)$. We conclude (by strong minimality) that cofinitely many of the lines of slope $\alpha$ intersect $X$ in at least $m+1$ points, and thus that $l\in\operatorname{acl}_{\mathcal{M}}(t\alpha)$. It is now an exercise in forking calculus (using the independence of $\alpha$ from $tab$) to show that this implies $(a,b)\in\operatorname{acl}_{\mathcal{M}}(t)$.
In the abstract topological setting, we need a replacement for the ‘transversality’ of $X$ and $l$ above: that is, we need an axiom saying that the $m$ given intersection points of $X\cap l$ will not disappear when we perturb $l$. This is what we wish to define. It will be most natural to work at the level of types and germs as in the previous subsection.
To start, we define:
Definition 3.22.
Let $x\in K^{m}$ and $y\in K^{n}$. Let $A$, $B\supset A$, and $C\supset A$ be parameter sets. We call $(x,y,A,B,C)$ a transversality configuration if the following hold:
  1. (1)
    $\dim(x/B)=\dim(y/C)$.
  2. (2)
    $x\in\operatorname{acl}(By)$ and $y\in\operatorname{acl}(Cx)$.
  3. (3)
    $x$ is independent from $C$ over each of $A$ and $Ay$.
If $(x,y,A,B,C)$ is a transversality condition, we further define the intersection family of $(x,y,A,B,C)$ to be the $(BC)$-type-definable set $W$ of all $(x^{\prime},y^{\prime})$ such that $(x^{\prime},y^{\prime})\models\operatorname{tp}(x,y/C)$ and $x^{\prime}\models\operatorname{tp}(x/B)$.
To illustrate Definition 3.22, we revisit the example given above. In the language of the example, one should think of the following analogs:
  • $x$ is analogous to a point of $X\cap l$.
  • $y$ is analogous to the parameter for the line $l$.
  • $A=\emptyset$, and $\operatorname{tp}(x/A)$ is analogous to $K^{2}$ (or rather, $K^{2}$ is analogous to a d-approximation of $\operatorname{tp}(x/A)$).
  • $B$ is a parameter defining $X$, and $\operatorname{tp}(x/B)$ is analogous to $X$.
  • $C=\alpha$.
  • $\operatorname{tp}(x/Ay)$ is analogous to $l$ (or rather, the germ of $l$ at $x$).
  • The assumption that $\dim(x/B)=\dim(y/C)$ replaces the strong minimality of the set of lines of slope $\alpha$.
  • The assumption that $x\in\operatorname{acl}(By)$ replaces the finiteness of $X\cap l$.
  • The assumption that $y\in\operatorname{acl}(Cx)$ replaces the fact that $l$ is determined by its slope and one point.
  • The independence of $x$ and $C$ over $A$ and $Ay$ replaces the fact that $\alpha$ is chosen independently of the other data. In particular, this means that $\operatorname{tp}(x/Cy)$ is also analogous to $l$.
  • If $W$ is the intersection family of $(x,y,A,B,C)$, then the fibers of the projection $W\rightarrow\operatorname{tp}(y/C)$ correspond to the intersections of $X$ with translates of $l$.
Now suppose $(x,y,A,B,C)$ is any transversality configuration, with intersection family $W$. The general idea is to think of $\operatorname{tp}(y/C)$ as parametrizing the family of conjugates of $\operatorname{tp}(x/Cy)$, and the projection $W\rightarrow\operatorname{tp}(y/C)$ as recording the intersection points of each such conjugate with $\operatorname{tp}(x/B)$. The desired conclusion (at least in the relevant cases) is that the point $(x,y)\in W$ ‘moves’ when $y$ does. Abstractly, we want the projection $W\rightarrow\operatorname{tp}(y/C)$ to be open at $(x,y)$.
Before giving a precise statement, we need to check that $W$ is germed at $(x,y)$:
Lemma 3.23.
Let $(x,y,A,B,C)$ be a transversality configuration, with family of intersections $W$.
  1. (1)
    $W$ is germed at $(x,y)$.
  2. (2)
    $\dim(W)=\dim(x/B)=\dim(y/C)$.
Proof.
  1. (1)
    By repeated applications of instances of Fact 3.15. Namely, $W$ is the intersection of the complete type $\operatorname{tp}(x,y/C)$ with the set of $(x^{\prime},y^{\prime})$ satisfying $x^{\prime}\models\operatorname{tp}(x/B)$; and the latter is the product of $\operatorname{tp}(x/B)$ with (the definable set) $K^{n}$ (where $n$ is the length of $y$).
  2. (2)
    First, we check that $\dim(W)\leq\dim(x/B)$. Indeed, let $(x^{\prime},y^{\prime})\in W$. Then $y^{\prime}\in\operatorname{acl}(Cx^{\prime})$ and $\dim(x^{\prime}/BC)\leq\dim(x^{\prime}/B)$, so $\dim(x^{\prime}y^{\prime}/BC)\leq\dim(x/B)$.
    Now we show that $\dim(W)\geq\dim(x/B)$. It suffices to show that every d-approximation of $W$ has dimension at least $\dim(x/B)$. So let $W^{\prime}$ be such a d-approximation. Using Fact 3.11, after shrinking if necessary, we can write $W^{\prime}$ as the set of $(x^{\prime},y^{\prime})$ satisfying $(x^{\prime},y^{\prime})\in Z$ and $x^{\prime}\in X_{1}$, where $X_{1}$ is a d-approximation of $\operatorname{tp}(x/B)$ and $Z$ is a d-approximation of $\operatorname{tp}(xy/C)$. Let us also fix a d-approximation $X$ of $\operatorname{tp}(x/A)$. By Fact 3.11, we can assume that $X$, $X_{1}$, and $Z$ are definable over $A$, $B$, and $C$, respectively.
    Now let $S$ be the set of $x^{\prime}$ such that $(x^{\prime},y^{\prime})\in Z$ for some $y^{\prime}$. Then $S$ is definable over $C$ and contains all realizations of $\operatorname{tp}(x/C)$. So the germ of $S$ at $x$ contains the germ of $\operatorname{tp}(x/C)$ at $x$ (this means that there are a neighborhood $V$ of $x$, and a $C$-definable set $Y$ containing $x$, such that $Y\cap V\subset S\cap V$). On the other hand, since $x$ and $C$ are independent over $A$, and by Lemma 3.21, the germ of $\operatorname{tp}(x/C)$ at $x$ is the same as the germ of $\operatorname{tp}(x/A)$ at $x$ – that is, the germ of $X_{1}$ at $x$. In other words, we have shown that the germ of $S$ at $x$ contains the germ of $X_{1}$ at $x$, and thus there is a neighborhood $U$ of $x$ such that $X_{1}\cap U\subset S\cap U$.
    Now by construction, $x$ is generic in $X_{1}$ over $B$. By the Baire category axiom, there is $x^{\prime}\in U$ which is generic in $X_{1}$ over $BC$. Thus $x^{\prime}\in S$, and so there is some $y^{\prime}$ with $(x^{\prime},y^{\prime})\in Z$. Thus $(x^{\prime},y^{\prime})\in W^{\prime}$, and moreover,
    $\dim(x^{\prime}y^{\prime}/BC)\geq\dim(x^{\prime}/BC)=\dim(X_{1})=\dim(x/B),$
    which completes the proof.
Now that we know $W$ is germed at $(x,y)$, we can define various versions of what it means for a transversality configuration to actually be transverse:
Definition 3.24.
Let $(x,y,A,B,C)$ be a transversality configuration, with intersection family $W$.
  1. (1)
    $(x,y,A,B,C)$ is weakly transverse if $W\rightarrow\operatorname{tp}(y/C)$ is finite-to-one near $(x,y)$.
  2. (2)
    $(x,y,A,B,C)$ is transverse if $W\rightarrow\operatorname{tp}(y/C)$ is locally open and finite-to-one near $(x,y)$.
  3. (3)
    $(x,y,A,B,C)$ is strongly transverse if $x$ is independent from $C$ over $B$ (equivalently, if $(x,y)$ is generic in $W$ over $BC$).
We often abbreviate ‘weakly transverse transversality configuration’ by ‘weakly transverse configuration’, and similarly for transverse and strongly transverse configurations.
One can check that strong transversality implies transversality, which implies weak transversality, though we will not need this.
We are finally ready to introduce the notion of a Hausdorff geometric structure having enough open maps. The name refers to the local openness of the map $W\rightarrow\operatorname{tp}(y/C)$ above (see Definition 3.24) for a transversality configuration $(x,y,A,B,C)$. We would like to say that ‘enough of the time’, this map is open at $(x,y)$ – or in the language of Definition 3.24, ‘enough’ transversality configurations are transverse.
Our easiest desired conclusion would be that every transversality configuration is transverse. However, this seems too much to ask for. Instead, we need a more complicated statement involving all three transversality notions:
Definition 3.25.
We say that $(\mathcal{K},\tau)$ has enough open maps if the following holds: let $(x,y,A,B,C)$ and $(x,y,A,B,C^{\prime})$ be transversality configurations. If $(x,y,A,B,C)$ is weakly transverse, and $(x,y,A,B,C^{\prime})$ is strongly transverse, then $(x,y,A,B,C)$ is transverse.
Remark 3.26.
A simpler statement would have been ‘every weakly transverse configuration is transverse’. However, the condition of weak transversality alone did not seem to be enough to imply transversality in examples. So one could think of Definition 3.25 as saying that, in nice enough situations (i.e. when strong transversality is consistent), transversality can only fail for a very good reason (being non-weakly-transverse).
3.4. The Open Mapping Property and Enough Open Maps
Due to the technical nature of the notion of $(\mathcal{K},\tau)$ having enough open maps, we will now provide two other notions, each of which implies enough open maps and is easier to verify in examples. The first notion below is to be thought of as an ‘algebraically closed’ condition, and generalizes the open mapping theorem from complex analysis. The second is a ‘characteristic zero’ condition, and essentially amounts to a version of Sard’s Theorem.
Before giving either of these notions, we need to develop ‘smoothness’ in Hausdorff geometric structures:
Definition 3.27.
By a notion of smoothness on $(\mathcal{K},\tau)$, we mean a map $X\mapsto X^{S}$ sending each definable set $X\subset K^{n}$ (for all $n$) to a (not necessarily definable) subset $X^{S}\subset X$, satisfying the following:
  1. (1)
    If $x$ is generic in $X$ over any set of parameters defining $X$, then $x\in X^{S}$.
  2. (2)
    If $x\in X^{S}$ and $y\in Y^{S}$ then $(x,y)\in X^{S}\times Y^{S}$.
  3. (3)
    If $x\in X^{S}$ and $\sigma$ is a coordinate-permutation, then $\sigma(x)\in(\sigma(X))^{S}$.
  4. (4)
    If $Z\subset X\times Y$ are definable and $W=\{(x,x,y):(x,y)\in Z\}$, then for all $x,y$ we have $(x,y)\in Z^{S}$ if and only if $(x,x,y)\in W^{S}$.
  5. (5)
    Suppose $f:X\rightarrow Y$ is an $A$-definable projection, $x\in X$ and $y=f(x)\in Y$ are both generic over $A$. Let $Z$ be definable over $B$ so that $y\in Z$ is generic over $B$ and $Z\cap U\subset Y\cap U$ for some neighborhood $U$ of $y$. Then $x\in(f^{-1}(Z\cap Y))^{S}$.
  6. (6)
    The assertion ‘$x\in X^{S}$’ is a d-local property of $x$ and $X$.
If $x\in X^{S}$, we say that $x$ is smooth in $X$ or $X$ is smooth at $x$.
Remark 3.28.
Definition 3.27(5) is our attempt at capturing the submersion theorem of differential geometry (asserting that every submersion of smooth manifolds is locally diffeomorphic to a coordinate projection of affine spaces) and its corollary on fibers (that preimages of smooth manifolds under submersions have a smooth structure). For technical reasons involving the Frobenius map, we could not find a more direct statement of this nature that holds in ACVF and is suitable for our needs. For the reader’s intuition, we note the informal translation between our statement and the usual statement. Given our $f:X\rightarrow Y$, the genericity of $x\in X$ and $f(x)\in Y$ lets us informally visualize $f$ as a submersion near $x$. The submersion theorem would then allow us to visualize $X$ as $W\times Y$ for some $W$, and thus $x=(w,y)$ for some $w$ and $y$. The restriction of $X$ over $Z$ then corresponds to $W\times Z$, which remains smooth at $(w,y)$ by looking at $W$ and $Z$ separately and applying Definition 3.27(2) (where the smoothness of $Z$ comes from the genericity of $y\in Z$ over $B$).
Thus, informally, a notion of smoothness is any d-local notion which holds generically and is closed under products, permutations, concatenations, and preimages under nice enough maps. In many natural examples of fields (including ACVF), there is a canonical such notion coming from the smooth locus of a variety (see Lemma 9.36): one takes $X^{S}$ to be the points that are locally open in the smooth locus of the Zariski closure of $X$. For o-minimal expansions of real closed fields, one can fix $n\geq 1$ and declare $x\in X^{S}$ if after restricting to a neighborhood of $x$, $X$ becomes a $C^{n}$-submanifold of the ambient space.
By d-locality, any notion of smoothness extends to germs of type-definable sets. Thus, if $X$ is type-definable and germed at $x$, the assertion ‘$x\in X^{S}$’ is well-defined. We note the translations of Definition 3.27 (1) to this context:
Lemma 3.29.
Let $X\mapsto X^{S}$ be a notion of smoothness on $(\mathcal{K},\tau)$. Let $x\in K^{n}$, and let $A$ be a set of parameters. Then $\operatorname{tp}(x/A)$ is smooth at $x$.
In particular, we conclude:
Lemma 3.30.
Fix a notion of smoothness $X\mapsto X^{S}$ on $(\mathcal{K},\tau)$. Let $(x,y,A,B,C)$ be a transversality configuration, with family of intersections $W$. Then $(x,y)\in W^{S}$.
Proof.
By Lemma 3.23, $W$ is germed at $(x,y)$, so the conclusion makes sense. Let $W^{\prime}$ be a d-approximation of $W$ at $(x,y)$. By Fact 3.11, we may assume $W^{\prime}=\{(x,y)\in X:x\in Z\}$, where $X$ is some d-approximation of $\operatorname{tp}(x,y/C)$ and $Z$ is some d-approximation of $\operatorname{tp}(x/B)$. We may assume that $X$ and $Z$ are definable over $C$ and $B$, respectively.
Let $Y$ be a d-approximation of $\operatorname{tp}(x/C)$, which we may assume is definable over $C$. Shrinking $X$ if necessary, we may assume the projection of $X$ is contained in $Y$.
We want to show that $(x,y)$ is smooth in $W^{\prime}$, by applying Definition 3.27(5). Almost all the hypotheses are already met. It remains only to show that the germ of $Z$ at $x$ is contained in the germ of $Y$ at $x$ – that is, that the germ of $\operatorname{tp}(x/B)$ at $x$ is contained in the germ of $\operatorname{tp}(x/C)$ at $x$. But this follows from Lemma 3.21. Indeed, since $x$ and $C$ are independent over $A$, $\operatorname{tp}(x/A)$ and $\operatorname{tp}(x/C)$ have the same germ at $x$. So it suffices to show that the germ of $\operatorname{tp}(x/B)$ at $x$ is contained in the germ of $\operatorname{tp}(x/A)$ at $x$. And this is clear, since $B\supset A$. ∎
We now give our first simplified condition:
Definition 3.31.
Let $(\mathcal{K},\tau)$ be a Hausdorff geometric structure. We say that $(\mathcal{K},\tau)$ has the open mapping property if there is a notion of smoothness $X\mapsto X^{S}$ satisfying the following: suppose $X$ and $Y$ are definable of the same dimension, $f:X\rightarrow Y$ is a projection, $x\in X^{S}$, and $f(x)\in Y^{S}$. If $f$ is finite-to-one in a neighborhood of $x$, then $f$ is open in a neighborhood of $x$.
As with Lemma 3.29, we immediately get the following for germed sets:
Lemma 3.32.
Assume $(\mathcal{K},\tau)$ has the open mapping property, witnessed by the notion of smoothness $X\mapsto X^{S}$. Let $X$ and $Y$ be type-definable of the same dimension, and $f:X\rightarrow Y$ a projection which is germed at $x\in X$. If $f$ is finite-to-one near $x$, then $f$ is open near $x$.
The argument in [Cas23] crucially used that the complex field has the open mapping property, which follows from the usual open mapping theorem. Moreover, we will see later (Corollary 9.60) that ACVF has the open mapping property in all characteristics. Now we show:
Proposition 3.33.
Assume that $(\mathcal{K},\tau)$ has the open mapping property. Then every weakly transverse configuration is transverse. Thus, $(\mathcal{K},\tau)$ has enough open maps.
Proof.
Let $(x,y,A,B,C)$ be a weakly transverse configuration, with family of intersections $W$. By Lemma 3.23, $W$ is germed and has the same dimension as $\operatorname{tp}(y/C)$. By assumption, $W\rightarrow\operatorname{tp}(y/C)$ is finite-to-one near $(x,y)$. Thus, by Definition 3.31, $W\rightarrow\operatorname{tp}(y/C)$ is open near $(x,y)$. Thus $(x,y,A,B,C)$ is transverse. ∎
3.5. Differentiability and Enough Open Maps
We now introduce differentiability, which complements the open mapping property as another simplified condition implying enough open maps. Roughly speaking, we want to say that $(\mathcal{K},\tau)$ can be equipped with a tangent space functor satisfying versions of the Inverse Function Theorem and Sard’s Theorem.
Definition 3.34.
We say that $(\mathcal{K},\tau)$ is differentiable if there are a notion of smoothness $X\mapsto X^{S}$ on $(\mathcal{K},\tau)$, and a covariant functor $\mathcal{F}:\mathcal{C}\rightarrow\mathcal{D}$, such that the following hold:
  1. (1)
    $\mathcal{D}$ is the category of $F$-vector spaces for some field $F$.
  2. (2)
    The objects of $\mathcal{C}$ are smooth-pointed definable sets in $\mathcal{K}$: that is, pairs $(X,x)$ where $X\subset K^{n}$ is definable and $x\in X^{S}$.
  3. (3)
    For each object $(X,x)$ in $\mathcal{C}$, the dimension of $\mathcal{F}(X,x)$ as an $F$-vector space is precisely $\dim X$.
  4. (4)
    The morphisms of $\mathcal{C}$ are compositions of inclusions and projections.
  5. (5)
    $\mathcal{F}$ sends inclusions in $\mathcal{C}$ to inclusions in $\mathcal{D}$.
  6. (6)
    If $x\in X^{S}$ and $y\in Y^{S}$, then $\mathcal{F}(X\times Y,(x,y))$ is the direct product of $\mathcal{F}(X,x)$ and $\mathcal{F}(Y,y)$. Moreover, $\mathcal{F}$ sends projections in $\mathcal{C}$ to the corresponding projections in $\mathcal{D}$.
  7. (7)
    $\mathcal{F}$ sends constant functions in $\mathcal{C}$ to the zero map in $\mathcal{D}$.
  8. (8)
    (Weak Inverse Function Theorem) If $f:(X,x)\rightarrow(Y,y)$ is a morphism in $\mathcal{C}$, and $\mathcal{F}(f)$ is an isomorphism, then $f$ is an open map in a neighborhood of $x$.
  9. (9)
    (Sard’s Theorem) If a morphism $f:(X,x)\rightarrow(Y,y)$ in $\mathcal{C}$ is definable over $A$, $x$ is generic in $X$ over $A$, and $y$ is generic in $Y$ over $A$, then $\mathcal{F}(f)$ is surjective.
We will see later (Theorem 9.4 and Lemma 9.40) that o-minimal expansions of fields, as well as pure Henselian fields of characteristic 00, are differentiable. We now aim toward showing that differentiability implies enough open maps.
For the rest of this subsection, we assume that $(\mathcal{K},\tau)$ is differentiable, and fix witnessing data $X\mapsto X^{S}$, $\mathcal{C}$, $\mathcal{D}$, and $\mathcal{F}$.
Remark 3.35.
Following the intuition of tangent spaces, we will denote the space $\mathcal{F}(X,x)$ by $T_{x}(X)$, and call it the tangent space to $X$ at $x$.
First we check:
Lemma 3.36.
Tangent spaces are d-local. That is, suppose $x\in X\subset Y$, where $X$ is definable, $x\in X^{S}$, and $Y$ is a d-approximation of $X$ at $x$. Then $T_{x}(X)=T_{x}(Y)$.
Proof.
Since smoothness is d-local (Definition 3.27(6)), we have $x\in Y^{S}$, so $T_{x}(Y)$ is well-defined. Now by Definition 3.34(5), the inclusion $X\xhookrightarrow{}Y$ induces an inclusion $T_{x}(X)\xhookrightarrow{}T_{x}(Y)$. On the other hand, since $Y$ is a d-approximation of $X$, we have $\dim(X)=\dim(Y)$, and thus $\dim(T_{x}(X))=\dim(T_{x}(Y))$ (by Definition 3.34(3)). So $T_{x}(X)=T_{x}(Y)$. ∎
By Lemma 3.36, we can extend tangent spaces to germed type-definable sets at smooth points. That is, if $X$ is type-definable and germed at $x$, and $x\in X^{S}$, then $T_{x}(X)$ is well-defined.
Notation 3.37.
If $\operatorname{tp}(x/A)$ is a complete type, we abbreviate $T_{x}(\operatorname{tp}(x/A))$ by $T(x/A)$.
In the language of types, we conclude:
Lemma 3.38.
Let $x$ and $y$ be tuples, and $A\subset B$ parameter sets. Let $0_{y}$ denote the identity of $T(y/A)$.
  1. (1)
    $\dim(T(x/A))=\dim(x/A)$.
  2. (2)
    If $x$ and $B$ are independent over $A$, then $T(x/B)=T(x/A)$.
  3. (3)
    $T(xy/A)\rightarrow T(y/A)$ is surjective, with kernel $T(x/Ay)\times\{0_{y}\}$.
Proof.
(1) follows from Definition 3.34(3). (2) follows from Lemma 3.21. For (3), the surjectivity follows first from Sard’s Theorem. Now let $V_{1}=T(x/Ay)\times\{0_{y}\}$, and let $V_{2}$ be the kernel of $T(xy/A)\rightarrow T(y/A)$. We want to show that $V_{1}=V_{2}$. Notice that there is an inclusion $\operatorname{tp}(x/Ay)\times\{y\}\xhookrightarrow{}\operatorname{tp}(xy/A)$. By Definition 3.34(5), there is an induced inclusion $V_{1}\xhookrightarrow{}T(xy/A)$, whose image is contained in $V_{2}$ (by composing with $T(xy/A)\rightarrow T(y/A)$ and applying Definition 3.34(7)). Thus $V_{1}\subset V_{2}$, and it suffices to show that $\dim(V_{1})=\dim(V_{2})$. For this, since $T(xy/A)\rightarrow T(y/A)$ is surjective, and using (1), we have
$\dim(V_{2})=\dim(T(xy/A))-\dim(T(y/A))$
$=\dim(xy/A)-\dim(y/A)=\dim(x/A)=\dim(T(x/A))=\dim(V_{1}).$
The key observation in showing enough open maps is:
Lemma 3.39.
Let $(x,y,A,B,C)$ be a transversality configuration, with family of intersections $W$. Then the kernel of $T_{(x,y)}(W)\rightarrow T(y/C)$ is precisely $(T(x/Ay)\cap T(x/B))\times\{0_{y}\}$. In particular, it does not depend on $C$.
Proof.
By Lemma 3.38(3), $T(xy/C)\rightarrow T(x/C)$ is surjective. But $y\in\operatorname{acl}(Cx)$ by assumption, so $\dim(xy/C)=\dim(x/C)$, thus $\dim(T(xy/C))=\dim(T(x/C))$, and thus $T(xy/C)\rightarrow T(x/C)$ is an isomorphism. Next, by Lemma 3.38(2), we have $T(x/C)=T(x/A)$. Now there is an inclusion $\operatorname{tp}(x/B)\xhookrightarrow{}\operatorname{tp}(x/A)$, so there is an induced inclusion $T(x/B)\xhookrightarrow{}T(x/A)=T(x/C)$. In particular, we have $T(x/B)\subset T(x/C)$.
Let $V$ be the preimage of $T(x/B)$ under the isomorphism $T(xy/C)\rightarrow T(x/C)$. Note that there is an inclusion of $T_{(x,y)}(W)$ into $V$ (induced by the embedding $W\xhookrightarrow{}\operatorname{tp}(xy/C)$). On the other hand, by Lemma 3.23, we have
$\dim(T_{(x,y)}(W))=\dim(W)=\dim(x/B)=\dim(T(x/B))=\dim(V).$
So in fact $T_{(x,y)}(W)=V$ (that is, $T_{(x,y)}(W)$ is formed by restricting $T(xy/C)$ to those points whose $T_{x}(C)$ coordinate lies in $T_{x}(B)$).
In particular, the kernel of $T_{(x,y)}(W)\rightarrow T(y/C)$ is now the restriction of the kernel of $T(xy/C)\rightarrow T(y/C)$ to those elements whose $T_{x}(C)$ coordinate lies in $T_{x}(B)$. By Lemma 3.38(3), this is the same as $(T(x/Cy)\cap T(x/B))\times\{0_{y}\}$. Finally, since $x$ and $C$ are independent over $Ay$, Lemma 3.38(2) gives that $T(x/Cy)=T(x/Ay)$, and we are done. ∎
Finally, we conclude:
Proposition 3.40.
Under our assumption that $(\mathcal{K},\tau)$ is differentiable, $(\mathcal{K},\tau)$ has enough open maps.
Proof.
Let $(x,y,A,B,C)$ and $(x,y,A,B,C^{\prime})$ be transversality configurations, with $(x,y,A,B,C)$ weakly transverse and $(x,y,A,B,C^{\prime})$ strongly transverse. Let $W$ and $W^{\prime}$ be their respective families of intersections. Let $f$ and $f^{\prime}$ denote the maps $T_{(x,y)}(W)\rightarrow T(y/C)$ and $T_{(x,y)}(W^{\prime})\rightarrow\operatorname{tp}(y/C^{\prime})$, respectively. By Lemma 3.23, each of these maps goes between two vector spaces of the same dimension.
Since $(x,y,A,B,C^{\prime})$ is strongly transverse, $(x,y)$ is generic in $W^{\prime}$ over $BC^{\prime}$ and $y$ is generic in $\operatorname{tp}(y/C)$ over $BC^{\prime}$. Thus, by Sard’s Theorem, $f^{\prime}$ is surjective. By dimension considerations, $f^{\prime}$ thus has trivial kernel. Then, by Lemma 3.39, $f$ also has trivial kernel. Finally, by dimension considerations again, $f$ is surjective, and thus an isomorphism. It now follows by the weak inverse function theorem that $W\rightarrow\operatorname{tp}(y/C)$ is open near $(x,y)$, and thus $(x,y,A,B,C)$ is transverse. ∎
4. Strongly Minimal Relics
We now study the behavior of strongly minimal structures definable in $\mathcal{K}$. Our goal is to make progress toward the Zilber trichotomy for such structures, by showing that (assuming non-local modularity) they are able to partially reconstruct the topology $\tau$. Thus, we now fix:
Assumption 4.1.
From now until the end of section 7, we fix a non-locally modular $\emptyset$-definable strongly minimal $\mathcal{K}$-relic $\mathcal{M}=(M,...)$. In other words, $\mathcal{M}$ is a non-locally modular strongly minimal structure; the set $M$ is a subset of some $K^{d}$; and every $\emptyset$-definable set in $\mathcal{M}$ is also $\emptyset$-definable in $\mathcal{K}$. We also assume the language of $\mathcal{M}$ is countable; that $\operatorname{acl}_{\mathcal{M}}(\emptyset)$ is infinite; and that there is a $\emptyset$-definable excellent family of plane curves in $\mathcal{M}$ (see Definition 2.4).
Remark 4.2.
Similarly to Remark 3.3, one might be concerned about the assumptions above regarding $\emptyset$-definability and $\operatorname{acl}_{\mathcal{M}}(\emptyset)$. We point out that all of these conditions can be arranged harmlessly by adding a countable set of constants to the languages of $\mathcal{M}$ and $\mathcal{K}$.
Remark 4.3.
Since $\mathcal{K}$ is $\aleph_{1}$-saturated and the language of $\mathcal{M}$ is countable, one checks easily that $\mathcal{M}$ is also $\aleph_{1}$-saturated. We use this throughout.
4.1. Notational Conventions
Several confusions can arise from working simultaneously with two geometric structures. Before proceeding, we clarify a few issues:
Convention 4.4.
Following [Cas23], we make the following notational conventions to distinguish between $\mathcal{K}$ and $\mathcal{M}$:
  1. (1)
    Unless otherwise stated, all tuples are assumed to be taken in $\mathcal{M}^{\textrm{eq}}$, and all parameter sets are assumed to be countable sets of tuples from $\mathcal{M}^{\textrm{eq}}$.
  2. (2)
    The term rank, and the notation rk, will always refer to the notion of dimension in $\mathcal{M}$.
  3. (3)
    We use the notation $\mathcal{M}(A)$-definable to refer to sets definable in $\mathcal{M}$ over $A$. We similarly use $\mathcal{K}(A)$-definable, $\mathcal{M}(A)$-interpretable, and $\mathcal{K}(A)$-interpretable.
  4. (4)
    The term plane curve, and all properties of plane curves and families of plane curves, are interpreted in the sense of $\mathcal{M}$. Similarly, the terms stationary, stationary component, and canonical base always refer to $\mathcal{M}$-definable objects, and are interpreted in the sense of $\mathcal{M}$.
  5. (5)
    When referring to the dimension functions of $\mathcal{K}$, we will use the term dimension, and the notation dim.
  6. (6)
    The terms generic, independent, and algebraic, and the notation acl(A), are always interpreted in the sense of $\mathcal{K}$. We will refer to the corresponding terms in $\mathcal{M}$ with the prefix $\mathcal{M}$ (e.g. $\mathcal{M}$-generic), and algebraic closure in $\mathcal{M}^{\textrm{eq}}$ will be denoted $\operatorname{acl}_{\mathcal{M}}(A)$.
4.2. Stating the Main Result
Let us begin our work with $\mathcal{M}$ by stating our first main goal. As stated above, the idea is to show that $\mathcal{M}$ detects $\tau$ in a precise sense. Following [Cas23], we make the following definitions:
Definition 4.5.
Let $x=(x_{1},...,x_{n})\in M^{n}$ be a tuple. We say that $x$ is coordinate-wise generic if each $x_{i}$ is generic in $M$ over $\emptyset$ (that is, in the sense of $\mathcal{K}$).
Definition 4.6.
Let $X\subset M^{n}$ be $\mathcal{M}(A)$-definable, and let $\pi_{i},\pi_{j}:M^{n}\rightarrow M$ be the $i$th and $j$th projections. We say that $\pi_{i}$ and $\pi_{j}$ are independent on $X$ if for all $x=(x_{1},...,x_{n})\in X$ generic over $A$, the elements $x_{i}$ and $x_{j}$ are independent over $A$.
Remark 4.7.
It is easy to see that Definition 4.6 is unchanged if ‘generic’ and ‘independent’ are replaced by their counterparts in $\mathcal{M}$, and that moreover the definition as a whole does not depend on the particular parameter set $A$.
Definition 4.8.
We say that $\mathcal{M}$ weakly detects closures if the following holds: Suppose $X\subset M^{n}$ is $\mathcal{M}(A)$-definable of rank $r\geq 0$, and let $x=(x_{1},...,x_{n})\in\overline{X}$ be a coordinate-wise generic point. Then $\operatorname{rk}(x/A)\leq r$. Moreover, one of the following holds:
  1. (1)
    $\operatorname{rk}(x/A)<r$.
  2. (2)
    For all $i\neq j$ such that the projections $\pi_{i},\pi_{j}:M^{n}\rightarrow M$ are independent on $X$, the elements $x_{i}$ and $x_{j}$ are $\mathcal{M}$-independent over $A$.
Example 4.9.
To aid in understanding condition (2) above, we point out that it always holds for non-trivial plane curves. Indeed, let $X$, $A$, and $x$ be as in Definition 4.8, such that, moreover, $X$ is a non-trivial plane curve. Then for generic $(y_{1},y_{2})\in X$, the $y_{i}$ are interalgebraic, and thus dependent, over $A$. So the two projections $X\rightarrow M$ are dependent, and thus (2) is vacuous in this case.
Now our first main result is:
Theorem 4.10.
If $(\mathcal{K},\tau)$ has enough open maps, then $\mathcal{M}$ weakly detects closures.
We postpone the proof to the next section.
4.3. Coherence
Before moving toward the proof of Theorem 4.10, we briefly study the notion of coherence, which allows us to move between the dim and rk functions smoothly, and in many instances to forgo computations with dim in favor of simpler computations with rk. The material in this subsection is analogous to [Cas23, §3.3].
We begin by noting the following basic facts, which are all easy and can be proven word for word as in [Cas23]:
Lemma 4.11.
Let $X$ be $\mathcal{M}(A)$-interpretable, and let $a\in X$.
  1. (1)
    $\dim X=\operatorname{rk}X\cdot\dim M$.
  2. (2)
    $\dim(a/A)\leq\operatorname{rk}(a/A)\cdot\dim M$.
  3. (3)
    If $a$ is generic in $X$ over $A$, then $a$ is $\mathcal{M}$-generic in $X$ over $A$.
In light of Lemma 4.11(2), the following makes sense:
Definition 4.12.
Let $a$ be a tuple and $A$ a set. We say that $a$ is coherent over $A$ if $\dim(a/A)=\operatorname{rk}(a/A)\cdot\dim M$. We say that $a$ is coherent if $a$ is coherent over $\emptyset$.
We will aim to work with coherent tuples as much as possible, because computations with them tend to be significantly easier. The following facts are very helpful:
Lemma 4.13.
Let $X$ be $\mathcal{M}(A)$-interpretable, let $a\in X$, and let $b$ be any tuple.
  1. (1)
    $a$ is generic in $X$ over $A$ if and only if it is both $\mathcal{M}$-generic in $X$ over $A$ and coherent over $A$.
  2. (2)
    $ab$ is coherent over $A$ if and only if $a$ is coherent over $A$ and $b$ is coherent over $Aa$.
  3. (3)
    If $a$ is coherent over $A$ and $b\in\operatorname{acl}_{\mathcal{M}}(Aa)$, then $b$ is coherent over $A$ and $a$ is coherent over $Ab$.
  4. (4)
    If $a$ is coherent over $A$ and $a$ is independent from $B$ over $A$, then $a$ is coherent over $AB$, and $a$ is $\mathcal{M}$-independent from $B$ over $A$.
  5. (5)
    If $a$ is coherent over $A$ and $B$ is any set, then there is $b$ realizing $\operatorname{tp}_{\mathcal{K}}(a/A)$ with $b$ coherent over $AB$.
  6. (6)
    There is a realization $a^{\prime}$ of $\operatorname{tp}_{\mathcal{M}}(a/A)$ which is coherent over $A$.
Proof.
(1), (2), and (3) are exactly as in [Cas23] (Lemmas 3.16 and 3.17). For (5), let $b$ be any relation of $\operatorname{tp}_{\mathcal{K}}(a/A)$ which is independent from $B$ over $A$, and apply (4).
For (6), choose an $\mathcal{M}(A)$-interpretable set $Y$ of minimal Morley rank and degree containing $a$, so that any $\mathcal{M}$-generic element of $Y$ over $A$ has the same $\mathcal{M}$-type as $a$ over $A$. Now let $a^{\prime}$ be a generic element of $Y$ over $A$, and apply (1).
It remains only to show (4). So assume $a$ is coherent over $A$ and independent from $B$ over $A$. Then $\dim(a/A)=\dim(a/AB)=\operatorname{rk}(a/A)\cdot\dim(M)=k$, say. The statement of (4) would follow if we can show $k=\operatorname{rk}(a/AB)\cdot\dim(M)$. But $\operatorname{rk}(a/AB)\cdot\dim(M)\leq k$ follows since $\operatorname{rk}(a/AB)\leq\operatorname{rk}(a/A)$, and $\operatorname{rk}(a/AB)\cdot\dim(M)\geq k$ follows by Lemma 4.11(2). ∎
5. The Proof of Theorem 4.10
Let us now begin the proof of Theorem 4.10. Almost all of the proof is independent of the topology on $\mathcal{K}$, and uses only dimension computations in $\mathcal{K}$ and $\mathcal{M}$. In particular, most of the proof can be copied word-for-word from [Cas23]. The key difference occurs in an important geometric step in the proof, where the assumption of enough open maps (Definition 3.25) replaces some application of complex geometry in the original proof.
Rather than duplicating twenty pages of material, we will largely give an outline of the proof, pointing out which parts can be copied and which ones involve the topological axioms. In the former case, we will point the reader to the identical steps in [Cas23] for details. We will then focus almost exclusively on the latter case.
As in Definition 4.8, we are given an $\mathcal{M}(A)$-definable set with coordinate-wise generic closure point $x$. Our goal is to show that $\operatorname{rk}(x/A)\leq r$, and that one of the two options in the conclusion of Definition 4.8 holds. The overarching structure of the proof will be an induction on the value $r=\operatorname{rk}(X)$, with almost the whole proof contained in the inductive step. Thus, we make the following convention (as we will see later, the case $r=0$ is trivial):
Assumption 5.1.
Until otherwise stated, we assume that $r\geq 1$ is fixed and the statement of Theorem 4.10 holds for all $r^{\prime}<r$.
5.1. The Main Argument
As in [Cas23], we deduce the inductive step of Theorem 4.10 from a different, more technical statement (Proposition 5.4 below), whose proof contains the main geometric idea of the argument as a whole. This proposition will apply to generic members of almost faithful families of non-trivial hypersurfaces. Let us first recall what this means:
Definition 5.2.
An $r$-hypersurface is an $\mathcal{M}$-definable subset of $M^{r+1}$ of rank $r$. An $r$-hypersurface $H$ is non-trivial if for $(x_{0},...,x_{r})\in H$ generic over the parameters defining $H$ in $\mathcal{M}$, any $r$ of $x_{0},...,x_{r}$ have rank $r$ over the same parameters.
Remark 5.3.
Note that a plane curve is the same as a 1-hypersurface; in particular, one easily checks that Definition 5.2 generalizes non-triviality for plane curves (introduced in Subsection 2.2).
As discussed in Section 2.2, we say that an $\mathcal{M}$-definable family $\{H_{t}:t\in T\}$ of $r$-hypersurfaces is almost faithful if for each $t\in T$, there are only finitely many $t^{\prime}\in T$ with $\operatorname{rk}(H_{t}\cap H_{t^{\prime}})=r$. We will (later on) use the existence of such families. Namely, let $H$ be a stationary non-trivial hypersurface. Then, as also discussed in Section 2.2, there are an $\mathcal{M}(\emptyset)$-definable almost faithful definable family of (without loss of generality non-trivial) hypersurfaces $\{H_{t}:t\in T\}$, and an $\mathcal{M}$-generic $\hat{t}\in T$, such that $H$ is almost equal to $H_{\hat{t}}$. See [Cas23, Lemma 2.25] as well as the discussion following Definition 8.11 loc. cit. for details. Moreover, we may assume $T\subseteq M^{n}$ For some $n$, and if $H$ is defined over a coherent parameter (over $\emptyset$), then $\hat{t}$ will be generic in $T$ (i.e. in the sense of $\mathcal{K}$).
Now, the main technical step of Theorem 4.10 (analogous to [Cas23, Proposition 8.15]) is:
Proposition 5.4.
Let =$\mathcal{H}=\{H_{t}:t\in T\}$ be an almost faithful family of non-trivial $r$-hypersurfaces of rank $k>(r+1)\cdot\dim M$, and let $\mathcal{C}=\{C_{s}:s\in S\}$ be an excellent family of plane curves. Assume that each of $\mathcal{H}$ and $\mathcal{C}$ is $\mathcal{M}(A)$-definable for some countable set $A$. Let $\hat{t}\in T$ and $\hat{x}=(\hat{x}_{0},...\hat{x}_{r})\in M^{r+1}$ each be generic over $A$, and assume that $\hat{x}\in\overline{H_{\hat{t}}}$. Then at least one of the following happens:
  1. (1)
    $\operatorname{rk}(\hat{x}/A\hat{t})<r$.
  2. (2)
    There is some $x\in H_{\hat{t}}$ such that for each $i=1,...,r$, there are infinitely many $s\in S$ with $(\hat{x}_{0},\hat{x}_{i}),(x_{0},x_{i})\in C_{s}$.
  3. (3)
    There is some $i\geq 2$ such that $\hat{x}_{i}\in\operatorname{acl}_{\mathcal{M}}(A\hat{t}\hat{x}_{0}...\hat{x}_{i-1})$.
Proof.
For ease of notation, we will assume $A=\emptyset$.
Arguing by contradiction, assume that each of (1), (2), and (3) fails. As in [Cas23], we will proceed by studying the pairwise intersections of $H_{\hat{t}}$ with a certain family of curves in $M^{r+1}$. We begin by defining the various objects we will need for the argument.
Notation 5.5.
For the rest of the proof of Proposition 5.4, we fix the following:
  • Let $H\subset M^{r+1}\times T$ be the graph of $\mathcal{H}$.
  • Let $U=S^{r}$.
  • For $u=(s_{1},...,s_{r})\in U$, set
    $D_{u}=\{(x_{0},...,x_{r})\in M^{r+1}:(x_{0},x_{i})\in C_{s_{i}}\textrm{ for each }i=1,...,r\}.$
  • Let $\mathcal{D}$ be the family of $D_{u}$ as $u$ varies in $U$, and let $D\subset M^{r+1}\times U$ denote its graph.
  • Let $\hat{p}=(\hat{p}_{1},...,\hat{p}_{r})$ be a fixed element of $M^{r}$ generic over $\hat{t}\hat{x}$.
  • Let $U_{\hat{p}}$ be the set of all $u=((p_{1},q_{1}),...,(p_{r},q_{r}))\in U$ such that $p_{i}=\hat{p}_{i}$ for each $i=1,...,r$.
  • Let $\mathcal{Y}$ be the subfamily of $\mathcal{D}$ indexed by $U_{\hat{p}}$, with graph $Y\subset M^{r+1}\times U_{\hat{p}}$.
  • Let $I$ be the set of $(x,u)\in Y$ with $x\in H_{\hat{t}}$.
  • For $x\in M^{r+1}$, we let ${{}_{x}D}:=\{u\in U:x\in D_{u}\}$ and ${{}_{x}Y}:={{}_{x}D}\cap U_{\hat{p}}$.
Remark 5.6.
A minor point of departure from [Cas23] is the definition of $I$. In [Cas23], $I$ also had a $T$-coordinate, so one took triples $(x,t,u)$ with $x\in H_{t}\cap D_{u}$. Our definition essentially prohibits $t$ from varying. This will make the current version of the argument easier, though the role played by $I$ will be the same.
The following facts are all now easy to see. Note that (5) is Lemma 8.21 of [Cas23].
Fact 5.7.
For the various objects specified above, we have:
  1. (1)
    $U$ is stationary of rank $2r$, and $U_{\hat{p}}$ is stationary of rank $r$.
  2. (2)
    $\mathcal{D}$ and $\mathcal{Y}$ are families of rank one subsets of $M^{r+1}$.
  3. (3)
    For generic $x\in M^{r+1}$, we have $\operatorname{rk}({{}_{x}D})=r$ and $\operatorname{rk}({{}_{x}Y})=0$.
  4. (4)
    For independent generic $x,x^{\prime}\in M^{r+1}$, we have $\operatorname{rk}({{}_{x}D}\cap{{}_{x^{\prime}}D})=0$.
  5. (5)
    The projection $Y\rightarrow M^{r+1}$ is finite-to-one and almost surjective.
  6. (6)
    $\hat{x}$ is generic in $M^{r+1}$ over $\hat{p}$.
By Fact 5.7 (5), (6) and the genericity of $\hat{x}$, the set ${{}_{\hat{x}}Y}$ is non-empty and finite.
Notation 5.8.
For the rest of the proof of Proposition 5.4, fix $\hat{u}\in{{}_{\hat{x}}Y}$.
The following facts about $\hat{u}$ are also straightforward (see [Cas23, Lemma 8.23, Lemma 8.24] for the proofs):
Fact 5.9.
$\hat{u}$ is generic in each of the following senses:
  1. (1)
    in $U$ over $\emptyset$; in particular, $\hat{u}$ is coordinate-wise generic.
  2. (2)
    in $U_{\hat{p}}$ over $\hat{p}$.
  3. (3)
    in ${{}_{\hat{x}}D}$ over $\hat{t}\hat{x}$.
Later, on multiple occasions, we will apply the inductive hypothesis of Theorem 4.10 with $\hat{u}$ as the distinguished closure point of a certain set. Fact 5.9(1) says that these applications will be valid.
Now that we have the main pieces in place, let us outline the proof. The argument can be organized into the following four main steps:
Step I: $\hat{u}$ is $\mathcal{M}$-generic in $U_{\hat{p}}$ over $\hat{t}\hat{p}$.
Step II: If $x\in H_{\hat{t}}\cap D_{\hat{u}}$, then $x$ is generic in $H_{\hat{t}}$ over $\hat{t}\hat{x}$, in $M^{r+1}$ over $\hat{x}$, and in $D_{\hat{u}}$ over $\hat{u}\hat{x}$.
Step III: If $x\in H_{\hat{t}}\cap D_{\hat{u}}$, then the projection $I\rightarrow U_{\hat{p}}$ is open in a neighborhood of $(x,\hat{t},\hat{u})$.
Step IV: $\hat{u}$ is not $\mathcal{M}$-generic in $U_{\hat{p}}$ over $\hat{t}\hat{p}$.
Note that Steps I and IV are contradictory, so this will conclude the proof. Step I is done in isolation from the other three. Meanwhile, II is used to prove III, and III is used to prove IV.
Steps I and II have nothing to do with the topology, and are proven entirely by rank and dimension computations. They can be copied word-for-word from [Cas23]. We will not reproduce the details here.
Step III (or rather an analogous statement) is done in [Cas23] using complex geometry. In that setting, II allows one to reduce III to a problem about smooth complex spaces and holomorphic maps, and apply an open mapping theorem from complex analysis. In the abstract setting, we will prove III using the assumption that $(\mathcal{K},\tau)$ has enough open maps. Indeed, most of the required hypotheses are straightforward to verify, with the most difficult ones being given exactly by II.
Step IV is the key topological argument. Roughly, we show using III that $\hat{u}$ belongs to the closure of an $\mathcal{M}(\hat{t}\hat{p})$-definable set of small rank, and conclude using the inductive hypothesis. This argument is largely unchanged from [Cas23], but we still give the details because it is the most crucial step in the proof of Theorem 4.10.
Step I
Fact 5.10.
$\hat{u}$ is $\mathcal{M}$-generic in $U_{\hat{p}}$ over $\hat{t}\hat{p}$.
Proof.
Exactly the same as Lemma 8.25 of [Cas23]. ∎
Step II
Fact 5.11.
If $x\in H_{\hat{t}}\cap D_{\hat{u}}$, then $x$ is generic in each of the following senses:
  1. (1)
    in $H_{\hat{t}}$ over $\hat{t}\hat{x}$.
  2. (2)
    in $M^{r+1}$ over $\hat{x}$.
  3. (3)
    in $D_{\hat{u}}$ over $\hat{u}\hat{x}$.
Proof.
Copy Lemmas 8.29, 8.34, 8.39, 8.40, 8.41, and 8.44 from [Cas23]. ∎
Step III
First, we need the following, which is also used in Step IV:
Lemma 5.12.
The intersection $H_{\hat{t}}\cap D_{u}$ is finite for all $u\in U_{\hat{p}}$ in some neighborhood of $\hat{u}$.
Proof.
Let $J$ be the set of $u\in U_{\hat{p}}$ such that $H_{\hat{t}}\cap D_{u}$ is infinite. So $J$ is $\mathcal{M}$-definable over $\hat{t}\hat{p}$, and if the lemma fails, then $\hat{u}\in\overline{J}$. We want to apply the inductive hypothesis to $J$. Thus we check:
Claim 5.13.
$\operatorname{rk}(J)<r$.
Proof.
If not, there is $u\in J$ with $\operatorname{rk}(u/\hat{t}\hat{p})=r$. Since $u\in J$, there is $x\in H_{\hat{t}}\cap D_{u}$ with $\operatorname{rk}(x/\hat{t}\hat{p}u)\geq 1$. So $\operatorname{rk}(xu/\hat{t}\hat{p})\geq r+1$. On the other hand, $\operatorname{rk}(x/\hat{t})\leq r$ since $x\in H_{\hat{t}}$, and $\operatorname{rk}(u/\hat{p}x)=0$ since $u\in{{}_{x}Y}$. Thus $\operatorname{rk}(xu/\hat{t}\hat{p})\leq r$, a contradiction. ∎
Now suppose the lemma fails, so $\hat{u}\in\overline{J}$. Then by the inductive hypothesis we get $\operatorname{rk}(\hat{u}/\hat{t}\hat{p})\leq\operatorname{rk}(J)<r$, contradicting Lemma 5.10. ∎
Recall that our aim in this step is to show that if $x\in H_{\hat{t}}\cap D_{\hat{u}}$, then the projection $I\rightarrow U_{\hat{p}}$ is open in a neighborhood of $(x,\hat{u})$. So for the rest of Step III, let us fix $x\in H_{\hat{t}}\cap D_{\hat{u}}$. We will deduce the desired openness of $I\rightarrow U_{\hat{p}}$ using an application of the assumption on enough open maps for $(\mathcal{K},\tau)$. Namely, in the terminology of transversality configurations from section 3, we show:
  • $(x,\hat{u},\emptyset,\hat{t},\hat{p})$ and $(x,\hat{u},\emptyset,\hat{u},\hat{x})$ are transversality configurations.
  • $(x,\hat{u},\emptyset,\hat{t},\hat{p})$ is weakly transverse.
  • $(x,\hat{u},\emptyset,\hat{t},\hat{x})$ is strongly transverse.
It will follow from enough open maps that $(x,\hat{u},\emptyset,\hat{t},\hat{p})$ is transverse. Thus $W_{\hat{p}}\rightarrow\operatorname{tp}(\hat{u}/\hat{p})$ is open near $(x,\hat{u})$, where $W_{\hat{p}}$ is the family of intersections of $(x,\hat{u},\emptyset,\hat{t},\hat{p})$. Finally, we note that $I\rightarrow U_{\hat{p}}$ realizes the same germ as $W_{\hat{p}}\rightarrow\operatorname{tp}(\hat{u}/\hat{p})$ at $(x,\hat{u})$, so is also open at $(x,\hat{u})$.
Let us proceed. We begin by showing:
Lemma 5.14.
$(x,\hat{u},\emptyset,\hat{t},\hat{p})$ and $(x,\hat{u},\emptyset,\hat{t},\hat{x})$ are transversality configurations.
Proof.
The three properties in Definition 3.22 translate into the following three claims:
Claim 5.15.
$\dim(x/\hat{t})=\dim(\hat{u}/\hat{p})=\dim(\hat{u}/\hat{x})$.
Proof.
By Fact 5.11(1), $\dim(x/\hat{t})=\dim(H_{\hat{t}})$. By Fact 5.9(2), $\dim(\hat{u}/\hat{p})=\dim(U_{\hat{p}})$. By Fact 5.9(3), $\dim(\hat{u}/\hat{x})=\dim(_{\hat{x}}D)$. So it suffices to show that $\dim(H_{\hat{t}})=\dim(U_{\hat{p}})=\dim(_{\hat{x}}D)$. By Lemma 4.11, it moreover suffices to show that $\operatorname{rk}(H_{\hat{t}})=\operatorname{rk}(U_{\hat{p}})=\operatorname{rk}(_{\hat{x}}D)$. But $\operatorname{rk}(H_{\hat{t}})=r$ by construction, $\operatorname{rk}(U_{\hat{p}})=r$ by Fact 5.7(1), and $\operatorname{rk}(_{\hat{x}}D)=r$ by Fact 5.7(3). ∎
Claim 5.16.
$x\in\operatorname{acl}(\hat{t}\hat{u})$, $\hat{u}\in\operatorname{acl}(\hat{p}x)$, and $\hat{u}\in\operatorname{acl}(\hat{x}x)$.
Proof.
Lemma 5.12 gives $x\in\operatorname{acl}(\hat{t}\hat{u})$ immediately. That $\hat{u}\in\operatorname{acl}(\hat{p}x)$ follows from Fact 5.7(5). That $\hat{u}\in\operatorname{acl}(\hat{x}x)$ follows from Facts 5.11(2) and 5.7(4). ∎
Claim 5.17.
$x$ is independent from each of $\hat{p}$ and $\hat{x}$ over each of $\emptyset$ and $\hat{u}$.
Proof.
That $x$ is independent from $\hat{x}$ over $\emptyset$ follows from Fact 5.11(2). That $x$ is independent from $\hat{x}$ over $\hat{u}$ follows from Fact 5.11(3). That $x$ is independent from $\hat{p}$ over $\hat{u}$ is automatic since $\hat{p}\in\operatorname{acl}(\hat{u})$ (indeed $\hat{p}$ contains $\hat{u}$ among its coordinates).
Finally, we check that $x$ is independent from $\hat{p}$ over $\emptyset$. First, it follows from Facts 5.9(2) and 5.11(3) that $(x,\hat{u})$ is generic in $Y$ over $\hat{p}$. But by Fact 5.7, this implies that $x$ is generic in $M^{r+1}$ over $\hat{p}$, which is enough. ∎
As $(x,\hat{u},\emptyset,\hat{t},\hat{p})$ and $(x,\hat{u},\emptyset,\hat{t},\hat{x})$ are transversality configurations, we can form the associated families of intersections, $W_{\hat{p}}$ and $W_{\hat{x}}$. Note that if $(x^{\prime},u^{\prime})\in W_{\hat{p}}$, then $\operatorname{tp}(x^{\prime}/\hat{t})=\operatorname{tp}(x/\hat{t})$ and $\operatorname{tp}(x^{\prime}u^{\prime}/\hat{p})=\operatorname{tp}(x\hat{u}/\hat{p})$. In particular, $x^{\prime}\in H_{\hat{t}}\cap D_{u^{\prime}}$. The following is then immediate:
Lemma 5.18.
$(x,\hat{u},\emptyset,\hat{t},\hat{p})$ is weakly transverse.
Proof.
By Lemma 5.12. ∎
To finish setting up the application of enough open maps, we check:
Lemma 5.19.
$(x,\hat{u},\emptyset,\hat{t},\hat{x})$ is strongly transverse.
Proof.
An equivalent statement is that $x$ and $\hat{x}$ are independent over $\hat{t}$. This follows from Fact 5.11(1). ∎
We may now conclude:
Lemma 5.20.
$(x,\hat{u},\emptyset,\hat{t},\hat{p})$ is transverse. That is, the projection $W_{\hat{p}}\rightarrow\operatorname{tp}(\hat{u}/\hat{p})$ is open at $(x,\hat{u})$.
Proof.
By the definition of enough open maps. ∎
To finish step III, it remains to relate Lemma 5.20 back to the projection $I\rightarrow U_{\hat{p}}$. To that end, we show:
Lemma 5.21.
The projections $I\rightarrow U_{\hat{p}}$ and $W_{\hat{p}}\rightarrow\operatorname{tp}(\hat{u}/\hat{p})$ have the same germ at $(x,\hat{u})$. Thus, $I\rightarrow U_{\hat{p}}$ is open at $(x,\hat{u})$.
Proof.
By their definitions, we can write
$I=Y\cap(H_{\hat{t}}\times U_{\hat{p}}),\;\;\;\;W_{\hat{p}}=\operatorname{tp}(x\hat{u}/\hat{p})\cap(\operatorname{tp}(x/\hat{t})\times\operatorname{tp}(\hat{u}/\hat{p})).$
So to show equality of germs, it is enough to show that each set in the intersection on the left hand side has the same germ as the corresponding set in the intersection on the right hand side. That is, it suffices to show the following:
Claim 5.22.
$(x,\hat{u})$ is generic in $Y$ over $\hat{p}$, and thus $Y$ realizes the germ of $\operatorname{tp}(x\hat{u}/\hat{p})$.
Proof.
The genericity statement follows from Facts 5.9(2) and 5.11(3). Now apply Lemma 3.20. ∎
Since the germ of a Cartesian product is the Cartesian product of the germs, to show the equality of the germs of $H_{\hat{t}}\times U_{\hat{p}}$ and $\operatorname{tp}(x/\hat{t})\times\operatorname{tp}(\hat{u}/\hat{p})$ we show:
Claim 5.23.
$x$ is generic in $H_{\hat{t}}$ over $\hat{t}$, and thus $H_{\hat{t}}$ realizes the germ of $\operatorname{tp}(x/\hat{t})$.
Proof.
The genericity statement follows from Fact 5.11(1). Now apply Lemma 3.20. ∎
Claim 5.24.
$\hat{u}$ is generic in $U_{\hat{p}}$, and thus $U_{\hat{p}}$ realizes the germ of $\operatorname{tp}(\hat{u}/\hat{p})$.
Proof.
The genericity statement follows from Fact 5.9(2). Now apply Lemma 3.20. ∎
Concluding the proof of the lemma. ∎
This finishes the proof of Step III.
Step IV
To obtain a contradiction to Step I it remains to show:
Lemma 5.25.
$\hat{u}$ is not $\mathcal{M}$-generic in $U_{\hat{p}}$ over $\hat{t}\hat{p}$.
Proof.
Let $w_{1},...,w_{l}$ be the distinct intersection points of $H_{\hat{t}}$ and $D_{\hat{u}}$ ($l$ is finite by Lemma 5.12). If $\hat{x}$ is among $w_{1},...,w_{l}$, then (2) in the statement of Proposition 5.4 holds with $x=\hat{x}$. Thus, we assume the $l+1$ points $\hat{x},w_{1},...,w_{l}$ are distinct.
Now let $Z$ be the set of all $u\in U_{\hat{p}}$ such that $H_{\hat{t}}\cap D_{u}$ is not of size $l$. So $Z$ is $\mathcal{M}$-definable over $\hat{t}\hat{p}$. Then the main point is the following:
Claim 5.26.
$\hat{u}\in\overline{Z}$.
Proof.
Let $V$ be any neighborhood of $\hat{u}$ in $U_{\hat{p}}$. We will find some $u\in V\cap Z$.
By Lemma 5.21, the projection $I\rightarrow U_{\hat{p}}$ is open in a neighborhood of each $(w_{i},\hat{u})$. Moreover, the projection $Y\rightarrow M^{r+1}$ is open in a neighborhood of $(\hat{x},\hat{u})$, by the generic local homeomorphism property of $(\mathcal{K},\tau)$ (see Fact 5.7 (5) and (6)). Then using each of these instances of local openness, and after applying various shrinkings, we can choose sets $W_{0},...,W_{l}$ such that:
  1. (1)
    $W_{0}$ is a neighborhood of $\hat{x}$ in $M^{r+1}$, and for each $i\geq 1$, $W_{i}$ is a neighborhood of $w_{i}$ in $M^{r+1}$.
  2. (2)
    The sets $W_{0},...,W_{l}$ are pairwise disjoint (this can be arranged since the topology $\tau$ is Hausdorff).
  3. (3)
    For each $x\in W_{0}$ there is some $u\in V$ such that $(x,u)\in Y$.
  4. (4)
    For each $u\in V$ and $i\geq l$ there is some $x\in W_{i}$ with $(x,u)\in I$.
Now using (1) and the fact that $\hat{x}\in\overline{H_{\hat{t}}}$, we can choose $x\in W_{0}\cap H_{\hat{t}}$, and then choose $u$ as in (3). Then, by (4), the intersection $H_{\hat{t}}\cap D_{u}$ contains $x$, in addition to one point in each $W_{i}$. By (2) all of these points are distinct, so $|H_{\hat{t}}\cap D_{u}|\geq l+1$, and thus $u\in V\cap Z$. ∎
We now finish the proof of Lemma 5.25 in two cases:
  • First, suppose that $Z$ is generic in $U_{\hat{p}}$. Then since $U_{\hat{p}}$ is stationary (see Fact 5.7(1)), $U_{\hat{p}}-Z$ is non-generic in $U_{\hat{p}}$. Since $\hat{u}\in U_{\hat{p}}-Z$, we are done.
  • Now suppose that $Z$ is non-generic in $U_{\hat{p}}$. So $\operatorname{rk}(Z)<r$. By Claim 5.26 and the inductive hypothesis, we get
    $\operatorname{rk}(\hat{u}/\hat{t}\hat{p})\leq\operatorname{rk}(Z)<r,$
    so we are again done.
Finally, by Steps I and IV (i.e. Fact 5.10 and Lemma 5.25), the proof of Proposition 5.4 is complete. ∎
5.2. Generic Non-trivial $r$-Hypersurfaces
The next three subsections complete the proof of Theorem 4.10, using Proposition 5.4 as the main tool. They are analogous to subsections 8.4-8.6 of [Cas23]. These remaining cases are essentially word-for-word copies of their analogs in [Cas23], so we will be brief. As in [Cas23], we use the following:
Definition 5.27.
Let $x=(x_{0},...,x_{r})\in M^{r+1}$ be a tuple, and $A$ a set of parameters. We say that $x$ satisfies * over $A$, also denoted $*(x,A)$, if the following hold:
  1. (1)
    $\operatorname{rk}(x/A)\leq r$.
  2. (2)
    If $\operatorname{rk}(x/A)=r\geq 2$ then for all $i\neq j\in\{0,...,r\}$ we have $\operatorname{rk}(x_{i}x_{j}/A)=2$.
As in [Cas23], we have the following properties:
Lemma 5.28.
Let $x$ and $y$ be tuples in $M^{r+1}$, and $A$ and $B$ sets of parameters.
  1. (1)
    If $x$ is $\mathcal{M}$-independent from $B$ over $A$, then $*(x,A)$ and $*(x,B)$ are equivalent.
  2. (2)
    If $x$ and $y$ are coordinate-wise $\mathcal{M}$-interalgebraic over $A$, then $*(x,A)$ and $*(y,A)$ are equivalent.
  3. (3)
    $*(x,A)$ is equivalent to $*(x,\operatorname{acl}_{\mathcal{M}}(A))$.
  4. (4)
    If $*(x,A)$ and $B\supset A$ then $*(x,B)$.
Let us now proceed. First we show that in the context of Theorem 5.4, the full statement of Theorem 4.10 follows from several applications of Proposition 5.4:
Proposition 5.29.
Let =$\mathcal{H}=\{H_{t}:t\in T\}$ be an almost faithful family of non-trivial $r$-hypersurfaces of rank $k>(r+1)\cdot\dim M$, and assume $\mathcal{H}$ is $\mathcal{M}$-definable over a set $A$. Let $\hat{t}\in T$ and $\hat{x}=(\hat{x}_{0},...,\hat{x}_{r})\in M^{r+1}$ each be generic over $A$, and assume that $\hat{x}\in\overline{H_{\hat{t}}}$. Then $*(\hat{x},A\hat{t})$ holds.
Proof.
Exactly as in [Cas23, Proposition 8.50]. ∎
5.3. Stationary Non-trivial $r$-Hypersurfaces
We next prove Theorem 4.10 for all stationary non-trivial hypersurfaces and all coordinate-wise generic points, without the added genericity assumptions of Proposition 5.29. As in [Cas23, §8.5], the trick is to use a sequence of independent plane curves to ‘translate’ the initial setup into a more generic one. That this is possible is provided by the local homeomorphism property of Definition 3.1.
Proposition 5.30.
Let $X$ be a stationary non-trivial $r$-hypersurface which is $\mathcal{M}$-definable over a set $A$. Let $\hat{x}=(\hat{x}_{0},...,\hat{x}_{r})$ be coordinate-wise generic, and assume that $\hat{x}\in\overline{X}$. Then $*(\hat{x},A)$ holds.
Proof.
As described above, the first step is to ‘translate’ $X$ and $\hat{x}$ along a sequence of ‘independent plane curves’. The following is straightforward and independent of the topology, and is proven in [Cas23] using pure rank and dimension computations:
Fact 5.31.
There are plane curves $C_{0},...,C_{r}$, tuples $c_{0},...,c_{r}$ from $\mathcal{M}^{\textrm{eq}}$, a positive integer $k>(r+1)\cdot\dim M$, and an element $\hat{y}=(\hat{y}_{0},...,\hat{y}_{r})\in M^{r+1}$, which satisfy the following:
  1. (1)
    Each $C_{i}$ is stationary and non-trivial.
  2. (2)
    Each $C_{i}$ is $\mathcal{M}$-definable over $c_{i}$, and each $c_{i}=\operatorname{Cb}(C_{i})$.
  3. (3)
    The $c_{i}$ are independent over $\emptyset$, and each individual $c_{i}$ is coherent of rank $k$ over $\emptyset$.
  4. (4)
    The tuple $(c_{0},...,c_{r})$ is independent from $A\hat{x}$ over $\emptyset$.
  5. (5)
    Each $(\hat{x}_{i},\hat{y}_{i})$ is a generic element of $C_{i}$ over $c_{i}$. In particular, $\hat{x}$ and $\hat{y}$ are coordinate-wise $\mathcal{M}$-interalgebraic over $Ac_{0}...c_{r}$.
  6. (6)
    $\hat{y}$ is generic in $M^{r+1}$ over $A\hat{x}$.
Proof.
See Lemmas 8.56 and 8.59 of [Cas23]. In particular, (5) is by the proof of Lemma 8.59. ∎
Fix $C_{0},...,C_{r},c_{0},...,c_{r},k,\hat{y}$ as in Fact 5.31. Let $c=(c_{0},...,c_{r})$. The tuple $\hat{y}$ is our ‘translated’ $\hat{x}$. We now ‘translate’ $X$ compatibly:
Notation 5.32.
We set
$Y=\{(y_{0},...,y_{r}):\textrm{ for some }(x_{0},...,x_{r})\in X\textrm{ we have that each }(x_{i},y_{i})\in C_{i}\}.$
So $Y$ is $\mathcal{M}$-definable over $Ac$. It is easy to verify that $Y$ is also a non-trivial $r$-hypersurface. Moreover, the assumption that $\hat{x}\in\overline{X}$ transfers to $\hat{y}$ and $Y$:
Lemma 5.33.
$\hat{y}\in\overline{Y}$.
Proof.
Let $V=V_{0}\times...\times V_{r}$ be any neighborhood of $\hat{y}$ in $M^{r+1}$. We will find an element $y\in V\cap Y$. Since each $(\hat{x}_{i},\hat{y}_{i})$ is generic in $C_{i}$ over $c_{i}$, the generic local homeomorphism property (see Definition 3.1) implies that the projection $\pi:C_{i}\rightarrow M$, $(x,y)\mapsto x$, is a homeomorphism near $(\hat{x}_{i},\hat{y}_{i})$. After applying various shrinkings, it follows that we can find a neighborhood $U=U_{0}\times...\times U_{r}$ of $\hat{x}$ in $M^{r+1}$, so that for each $i$ and each $u\in U_{i}$ there is some $v\in V_{i}$ with $(u,v)\in C_{i}$. Since $\hat{x}\in\overline{X}$, there is some $x=(x_{0},...,x_{r})\in U\cap X$. We thus obtain $y=(y_{0},...,y_{r})\in V$ with each $(x_{i},y_{i})\in C_{i}$, which implies that $y\in Y$, as desired. ∎
The rest of the proof of Proposition 5.30 is identical to the analogous proof in[Cas23], so we just give a brief outline.
Since $\hat{y}\in\overline{Y}$, it belongs to the closure of one of the stationary components of $Y$. So let $Z\subset Y$ be a stationary component, $\mathcal{M}$-definable over $\operatorname{acl}_{\mathcal{M}}(Ac)$, with $\hat{y}\in\overline{Z}$. Let $z=\operatorname{Cb}(Z)$. The main point is:
Fact 5.34.
$z$ is coherent of rank at least $k$ over $A$.
Proof.
Exactly as in Lemmas 8.61, 8.62 of [Cas23]. ∎
By Fact 5.34, we can realize $Z$ up to almost equality as a generic member of a large family. That is, let =$\mathcal{H}=\{H_{t}:t\in T\}$ be an almost faithful $\mathcal{M}(A)$-definable family of non-trivial $r$-hypersurfaces, and let $t\in T$ be generic over $A$, so that $Z$ is almost equal to $H_{t}$.
Lemma 5.35.
Keeping the above notation, $*(\hat{y},Azct)$ holds.
Proof.
If $\hat{y}\in\overline{H_{\hat{t}}}$, then Proposition 5.29 gives $*(\hat{y},At)$, and we conclude by Lemma 5.28(4). If $\hat{y}\notin\overline{H_{\hat{t}}}$, then $\hat{y}\in\overline{Z-H_{\hat{t}}}$. By assumption $\operatorname{rk}(Z-H_{\hat{t}})<r$, so by the inductive hypothesis, $\operatorname{rk}(\hat{y}/Azt)<r$, and we again conclude by Lemma 5.28. (To use the inductive hypothesis, one most note that $\hat{y}$ is coordinate-wise generic; this is by Fact 5.31(6)). ∎
To end the proof, we repeatedly use Lemma 5.28. First, by the definition of $Z$ and the almost faithfulness of $\mathcal{H}$, we have $z,t\in\operatorname{acl}_{\mathcal{M}}(Ac)$, so we get $*(\hat{y},Ac)$. Since $\hat{x}$ and $\hat{y}$ are coordinate-wise $\mathcal{M}$-interalgebraic over $Ac$, this implies $*(\hat{x},Ac)$. Finally, since $c$ is independent from $A\hat{x}$, this implies $*(x,A)$.
5.4. The Remaining Cases
The rest of the proof of Theorem 4.10 is quite straightforward, and does not use any properties of the topology $\tau$ other than those true in all topological spaces. Thus, we do not repeat the whole proof here.
Proposition 5.36.
Theorem 4.10 holds whenever $X\subset M^{n}$ has rank $r$ and $n\geq r+1$.
Proof.
Exactly as in Propositions 8.68, 8.70, and 8.75 of [Cas23]. ∎
We now drop the inductive assumption, i.e. Assumption 5.1, and summarize the completed proof of the theorem:
Proof of Theorem 4.10.
Let $X$, $n$, $A$, and $x=(x_{1},...,x_{n})$ be as in Definition 4.8. It is easy to see that the statement of the theorem holds if $x\in X$.
We now induct on $r=\operatorname{rk}(X)$. If $r=0$ then $x\in X$, so we are done. Now assume $r\geq 1$ and the statement holds for all $r^{\prime}<r$. Clearly $n\geq r$. If $n=r$ then, since $M^{n}$ is stationary, we either have $x\in X$ or $\operatorname{rk}(x/A)<r$. In either case, we are done. So assume $n\geq r+1$. Then by we done by Proposition 5.36. ∎
6. Consequences of Weak Detection of Closures
We now give applications of weak detection of closures to trichotomy problems. The main point is that weak detection of closures allows for the detection of ramification points of certain maps, and that this can be used in some cases to detect tangency of plane curves. As a corollary, assuming an appropriate ‘purity of ramification’ statement, we show that $\dim(M)=1$.
6.1. Weakly Generic Intersections
In sections 6 and 7, we will frequently consider geometric properties of pairwise intersections between definable families of sets. Many natural properties become more difficult to state in the abstract setup, because notions such as ‘variety’ are not accessible. In order to make the presentation smoother, we begin by establishing some terminology. The following will be used extensively:
Definition 6.1.
Let $\mathcal{Y}=\{Y_{t}:t\in T\}$ and $\mathcal{Z}=\{Z_{u}:u\in U\}$ be $\mathcal{K}(A)$-definable families of subsets of a $\mathcal{K}(A)$-definable set $X$.
  1. (1)
    By a weakly generic $(\mathcal{Y},\mathcal{Z})$-intersection over $A$, we mean a tuple $(x,t,u)$ such that $x\in X$, $t\in T$, $u\in U$, $(x,t)\in Y$, and $(x,u)\in Z$, are all generic over $A$.
  2. (2)
    Let $(x,t,u)$ be a weakly generic $(\mathcal{Y},\mathcal{Z})$-intersection over $A$. We call $(x,t,u)$ a strongly generic $(\mathcal{Y},\mathcal{Z})$-intersection over $A$ if $t$ and $u$ are independent over both $A$ and $Ax$.
  3. (3)
    Let $(x,t,u)$ be a weakly generic $(\mathcal{Y},\mathcal{Z})$-intersection over $A$. We call $(x,t,u)$ strongly approximable over $A$ if every neighborhood of $(x,t,u)$ contains a strongly generic $(\mathcal{Y},\mathcal{Z})$-intersection over $A$.
Remark 6.2.
This is a slight abuse of notation, as Definition 6.1 depends on $X$, $T$, and $U$. In our applications, these sets will be clear from context, so we just write $(\mathcal{Y},\mathcal{Z})$.
The idea is that a weakly generic $(\mathcal{Y},\mathcal{Z})$-intersection is an intersection $x\in Y_{t}\cap Z_{u}$ which is ‘as generic as possible’ when looking at each family individually; while a strongly generic $(\mathcal{Y},\mathcal{Z})$-intersection is as generic as possible when looking at the two families simultaneously.
Note that for $\mathcal{Y}$ and $\mathcal{Z}$, weakly generic $(\mathcal{Y},\mathcal{Z})$-intersections need not exist (e.g. if one of $Y$ and $Z$ concentrates only on a small portion of $X$). Moreover, there might be weakly generic $(\mathcal{Y},\mathcal{Z})$-intersections but no strongly generic $(\mathcal{Y},\mathcal{Z})$-intersections (e.g. if $\mathcal{Y}=\mathcal{Z}$ and the sets in $\mathcal{Y}$ are pairwise disjoint).
On the other hand, for families of plane curves in $\mathcal{M}$, one can always find both types of intersections:
Lemma 6.3.
Let $\mathcal{C}=\{C_{t}:t\in T\}$ and $\mathcal{D}=\{D_{u}:u\in U\}$ be almost faithful $\mathcal{M}(A)$-definable families of plane curves in $\mathcal{M}$, each of rank at least 2.
  1. (1)
    Suppose $t\in T$ and $u\in U$ are generic over $A$, and $x$ is generic in both $C_{t}$ over $At$, and $D_{u}$ over $Au$. Then $(x,t,u)$ is a weakly generic $(\mathcal{C},\mathcal{D})$-intersection.
  2. (2)
    There is a tuple $(x,t,u)$ satisfying the assumptions of (1) above. In particular, there is a weakly generic $(\mathcal{C},\mathcal{D})$-intersection.
  3. (3)
    Every weakly generic $(\mathcal{C},\mathcal{D})$-intersection is strongly approximable. In particular, there is a strongly generic $(\mathcal{C},\mathcal{D})$-intersection.
Proof.
To simplify notation, throughout, we assume $A=\emptyset$.
  1. (1)
    Everything is clear except the genericity of $x\in M^{2}$. But by Lemma 4.13, $x$ is coherent (because $t$ is coherent and $x$ is coherent over $t$). So it suffices to show that $x$ is $\mathcal{M}$-generic in $M^{2}$. This is well known easy (see [Cas23, Lemma 2.40]).
  2. (2)
    Let $I$ be the set of $(x,t,u)$ with $x\in C_{t}\cap D_{u}$. Fix $(x,t,u)\in I$ generic. Then $(x,t,u)$ is such a tuple (this also follows from [Cas23, Lemma 2.40]).
  3. (3)
    Let $(x,t,u)$ be a weakly generic $(\mathcal{C},\mathcal{D})$-intersection. Let $V$ be any neighborhood of $u$. It will suffice to find a strongly generic $(\mathcal{C},\mathcal{D})$-intersection of the form $(x,t,u^{\prime})$ with $u^{\prime}\in V$.
    Let ${{}_{x}}D$ be the set of $u^{\prime}$ with $x\in D_{u^{\prime}}$. Then $u$ is generic in ${{}_{x}}D$ over $x$. By the Baire category axiom, there is some $u^{\prime}\in V$ which is generic in ${{}_{x}}D$ over $tx$. It follows easily that $(x,t,u^{\prime})$ is a strongly generic $(\mathcal{C},\mathcal{D})$-intersection (again, using [Cas23, Lemma 2.40]).
6.2. Topological Ramification and Multiple Intersections
We will also extensively study the related notions of ‘ramification’ and ‘multiple intersections’ in sections 6 and 7. Let us now make these precise.
Definition 6.4.
Let $f:X\rightarrow Y$ be any function, where $X\subset K^{m}$ and $Y\subset K^{n}$. We say that $x_{0}\in X$ is a topological ramification point of $f$ if $(x_{0},x_{0})$ belongs to the closure of the set of $(x,x^{\prime})\in X^{2}$ with $x\neq x^{\prime}$ and $f(x)=f(x^{\prime})$.
Definition 6.5.
Let $\mathcal{Y}=\{Y_{t}:t\in T\}$ and $\mathcal{Z}=\{Z_{t}:t\in T\}$ be $\mathcal{K}(A)$-definable families of subsets of a $\mathcal{K}(A)$-definable set $X$. Let $(x,t,u)$ be a weakly generic $(\mathcal{Y},\mathcal{Z})$-intersection over $A$. We call $(x,t,u)$ a generic multiple $(\mathcal{Y},\mathcal{Z})$-intersection over $A$ if $(x,t,u)$ is a topological ramification point of the projection $I\rightarrow T\times U$, where $I$ is the set of $(x^{\prime},t^{\prime},u^{\prime})$ with $(x^{\prime},t^{\prime})\in Y$ and $(x^{\prime},u^{\prime})\in Z$.
6.3. Detection of Generic Multiple Intersections
The first main goal of section 6 is to show that, if $\mathcal{M}$ weakly detects closures (Definition 4.8), then $\mathcal{M}$ can detect multiple intersections of plane curves in a precise sense. It will be convenient to only consider curves coming from standard families. Recall (Definition 2.4):
Definition 6.6.
A standard family of plane curves over $A$ is an almost faithful $\mathcal{M}(A)$-definable family of plane curves $\mathcal{C}=\{C_{t}:t\in T\}$ such that:
  1. (1)
    $\operatorname{rk}(T)\geq 1$.
  2. (2)
    $T$ is a generic subset of $M^{\operatorname{rk}(T)}$.
  3. (3)
    For each $x\in M^{2}$, the set ${{}_{x}C}=\{t:x\in C_{t}\}$ has rank at most $\operatorname{rk}(T)-1$.
Remark 6.7.
Condition (3) does not matter to us. We include it for consistency with [Cas23].
Standard families always exist in the following sense:
Lemma 6.8.
Let $C$ be a strongly minimal plane curve in $\mathcal{M}$, with canonical base $c$. Assume that $c$ is coherent of rank at least 1 over $A$. Then there are a standard family of plane curves $\mathcal{C}=\{C_{t}:t\in T\}$ over $A$, and a generic $t\in T$ over $A$, with $C$ almost contained in $C_{t}$.
Proof.
See Lemmas 2.42 and 3.21 of [Cas23]. ∎
We can now state what it means for $\mathcal{M}$ to detect multiple intersections:
Definition 6.9.
We say that $\mathcal{M}$ detects multiple intersections if whenever $\mathcal{C}=\{C_{t}:t\in T\}$ and $\mathcal{D}=\{D_{u}:u\in U\}$ are standard families of plane curves over $A$, and $(x,t,u)$ is a generic multiple $(\mathcal{C},\mathcal{D})$-intersection over $A$, the parameters $t$ and $u$ are $\mathcal{M}$-dependent over $A$.
We now prove, in analogy to [Cas23, Theorem 9.1]:
Theorem 6.10.
If $\mathcal{M}$ weakly detects closures, then $\mathcal{M}$ detects multiple intersections.
Proof.
Let $\mathcal{C}=\{C_{t}:t\in T\}$ and $\mathcal{D}=\{D_{u}:u\in U\}$ be standard families of plane curves over $A$. For ease of notation, we assume $A=\emptyset$. Let $\hat{w}=(\hat{x},\hat{t},\hat{u})$ be a generic multiple $(\mathcal{C},\mathcal{D})$-intersection over $\emptyset$. Note that $\hat{x}$, $\hat{t}$, and $\hat{u}$ are generic in their respective powers of $M$, and thus $\hat{w}$ is coordinate-wise generic; this will guarantee that applications of weak detection of closures below are valid.
Let $I$ be the set of $(x,t,u)$ with $x\in C_{t}\cap D_{u}$ – so $\hat{w}$ is a topological ramification point of the projection $I\rightarrow T\times U$. Also let $Z\subset T\times U$ be the set of $(t,u)$ such that $C_{t}\cap D_{u}$ is infinite, and note that by almost faithfulness the projection $Z\rightarrow T$ is finite-to-one. Finally, let $r_{\mathcal{C}}=\operatorname{rk}(T)$ and $r_{\mathcal{D}}=\operatorname{rk}(U)$. So $\operatorname{rk}(\hat{t})=r_{\mathcal{C}}$ and $\operatorname{rk}(\hat{u})=r_{\mathcal{D}}$, and our goal is to show that $\operatorname{rk}(\hat{t}\hat{u})<r_{\mathcal{C}}+r_{\mathcal{D}}$.
Lemma 6.11.
We may assume that $(\hat{t},\hat{u})\notin\overline{Z}$.
Proof.
Assume that $(\hat{t},\hat{u})\in\overline{Z}$. Then by weak detection of closures and the fact that $Z\rightarrow T$ is finite-to-one, we have
$\operatorname{rk}(\hat{t}\hat{u})\leq\operatorname{rk}(Z)\leq\operatorname{rk}(T)=r_{\mathcal{C}}<r_{\mathcal{C}}+r_{\mathcal{D}},$
which proves the theorem in this case. ∎
So, assume that $(\hat{t},\hat{u})\notin\overline{Z}$. Let $I^{\prime}$ be the set of $(x,t,u)\in I$ such that $(t,u)\notin Z$. Then by assumption, it follows that $\hat{w}$ is still a topological ramification point of $I^{\prime}\rightarrow T\times U$. Let $P$ be the set of $(x,x^{\prime},t,u)$ such that $x\neq x^{\prime}$ and $(x,t,u),(x^{\prime},t,u)\in I^{\prime}$ – so the tuple $\hat{z}=(\hat{x},\hat{x},\hat{t},\hat{u})$ belongs to the frontier $\textrm{Fr}(P)$. Moreover, note by definition of $P$, that the projection $P\rightarrow T\times U$ is finite-to-one, which shows that $\operatorname{rk}(P)\leq r_{\mathcal{C}}+r_{\mathcal{D}}$.
Now we would like to apply weak detection of closures to $\hat{z}\in\operatorname{Fr}(P)$, but we need to ‘preen’ $P$ first to get the relevant projections to be independent on $P$. Let $\pi_{1},\pi_{2}:M^{2}\rightarrow M$ denote the two projections. Then we define the sets
$P_{1}=\{(x,x^{\prime},t,u)\in P:\pi_{1}(x)\neq\pi_{1}(x^{\prime})\},$
$P_{2}=\{(x,x^{\prime},t,u)\in P:\pi_{2}(x)\neq\pi_{2}(x^{\prime})\}.$
It is evident from the definition of $P$ that $P=P_{1}\cup P_{2}$. Thus, we get $\hat{z}\in\operatorname{Fr}(P_{i})$ for some $i=1,2$. Without loss of generality, we assume that $\hat{z}\in\operatorname{Fr}(P_{1})$. We then define:
Definition 6.12.
Let $x_{1}\neq x_{1}^{\prime}\in M$. Then $x_{1}$ and $x_{1}^{\prime}$ are extendable if the preimage of $(x_{1},x_{1}^{\prime})$ in $P_{1}$ under the projection to the first and third coordinates has rank at least $r_{\mathcal{C}}+r_{\mathcal{D}}-1$.
Let $E$ be the set of extendable pairs. It is evident from the definition, and the fact that $\operatorname{rk}(P)\leq r_{\mathcal{C}}+r_{\mathcal{D}}$, that $\operatorname{rk}(E)\leq 1$. Then, denoting $\hat{x}$ as the pair $(\hat{x}_{1},\hat{x}_{2})$, we conclude the following:
Lemma 6.13.
$(\hat{x}_{1},\hat{x}_{1})\notin\overline{E}$.
Proof.
If not, then $(\hat{x}_{1},\hat{x}_{1})\in\operatorname{Fr}(E)$, since the two coordinates of $(\hat{x}_{1},\hat{x}_{1})$ are equal. But since $\operatorname{rk}(E)\leq 1$, we have $\dim(E)\leq\dim M$; so if $(\hat{x}_{1},\hat{x}_{1})\in\operatorname{Fr}(E)$ then by the strong frontier inequality (see Definition 3.1) we get $\dim(\hat{x}_{1},\hat{x}_{1})<\dim M$, which contradicts that $\hat{x}$ is generic in $M^{2}$. ∎
Finally, let $P^{\prime}$ be the set of all $(x,x^{\prime},t,u)\in P_{1}$ such that $\pi_{1}(x)$ and $\pi_{1}(x^{\prime})$ are not extendable. Note that $P^{\prime}$ is $\mathcal{M}$-definable over $\emptyset$, because $P$ is. Moreover, by Lemma 6.13, it follows that $\hat{z}\in\operatorname{Fr}(P^{\prime})$. Then we note:
Lemma 6.14.
If $\operatorname{rk}(P^{\prime})=r_{\mathcal{C}}+r_{\mathcal{D}}$, then the projections to the first and third $M$-coordinates are independent on $P^{\prime}$.
Proof.
Assume that $\operatorname{rk}(P^{\prime})=r_{\mathcal{C}}+r_{\mathcal{D}}$, and let $(x,x^{\prime},t,u)\in P^{\prime}$ be generic. Let $x_{1}$ and $x_{1}^{\prime}$ be the first and third coordinates of $(x,x^{\prime},t,u)$ – that is, the first coordinate of each of $x$ and $x^{\prime}$. Then by definition of $P^{\prime}$, $x_{1}$ and $x_{1}^{\prime}$ are not extendable. By definition of extendability, we conclude that $\operatorname{rk}(xx^{\prime}tu/x_{1}x_{1}^{\prime})\leq r_{\mathcal{C}}+r_{\mathcal{D}}-2$. But by the choice of $(x,x^{\prime},t,u)$ we have $\operatorname{rk}(xx^{\prime}tu)=r_{\mathcal{C}}+r_{\mathcal{D}}$. So by additivity, we obtain $\operatorname{rk}(x_{1}x_{1}^{\prime})\geq 2$. Thus $x_{1}$ and $x_{1}^{\prime}$ are $\mathcal{M}$-independent $\mathcal{M}$-generics in $M$ over $\emptyset$, which is enough to prove the lemma. ∎
We now apply weak detection of closures to $P^{\prime}$. Since $P^{\prime}\subset P$ and $\hat{z}\in\operatorname{Fr}(P^{\prime})$, we first get
$\operatorname{rk}(\hat{z})\leq\operatorname{rk}(P^{\prime})\leq\operatorname{rk}(P)\leq r_{\mathcal{C}}+r_{\mathcal{D}}.$
We want this to be a strict inequality between $\hat{z}$ and $r_{\mathcal{C}}+r_{\mathcal{D}}$. But if equality holds then all terms above must be equal, so in particular $\operatorname{rk}(\hat{z})=\operatorname{rk}(P^{\prime})=r_{\mathcal{C}}+r_{\mathcal{D}}$. So by Lemma 6.14, the projections of $P^{\prime}$ to the first and third coordinates are independent. By weak detection of closures, this forces the first and third coordinates of $\hat{z}$ – i.e. $\hat{x}_{1}$ and $\hat{x}_{1}$ – to be $\mathcal{M}$-independent over $\emptyset$. But this is clearly not true, since $\hat{x}_{1}$ and $\hat{x}_{1}$ are equal generics.
So the inequality must be strict, i.e. $\operatorname{rk}(\hat{z})<r_{\mathcal{C}}+r_{\mathcal{D}}$. But $(\hat{t},\hat{u})$ is a subtuple of $\hat{z}$ – so we also get that $\operatorname{rk}(\hat{t}\hat{u})<r_{\mathcal{C}}+r_{\mathcal{D}}$, which proves the theorem. ∎
6.4. Purity of Ramification and the Dimension of $M$
Our next goal is to show that in contexts where purity of ramification applies, Theorem 6.10 is enough to show that $\dim(M)=1$. This is analogous to section 6 of [Cas23]. In the abstract setting, the notion of purity of ramification is more difficult to state – the issue being that we need to identify an appropriate class of maps which could reasonably be expected to have the relevant property (see Remark 6.17). To remedy this, we will use the terminology of weakly generic intersections developed in the previous subsection, with one addition:
Definition 6.15.
Let $\mathcal{Y}=\{Y_{t}:t\in T\}$ and $\mathcal{Z}=\{Z_{u}:u\in U\}$ be $\mathcal{K}(A)$-definable families of subsets of a $\mathcal{K}(A)$-definable set $X$. Let $(x,t,u)$ be a weakly generic $(\mathcal{Y},\mathcal{Z})$-intersection over $A$. The codimension of $(x,t,u)$ over $A$ is
$\operatorname{codim}_{A}(x,t,u)=\dim(t/Ax)+\dim(u/Ax)-\dim(tu/Ax).$
One should think of codimension as a measure of how far a weakly generic intersection is from being strongly generic. With this intuition, we now define:
Definition 6.16.
$(\mathcal{K},\tau)$ has ramification purity if the following holds: Let $\mathcal{Y}=\{Y_{t}:t\in T\}$ and $\mathcal{Z}=\{Z_{u}:u\in U\}$ be $\mathcal{K}(A)$-definable families of subsets of a $\mathcal{K}(A)$-definable set $X$. Suppose there is a generic multiple $(\mathcal{Y},\mathcal{Z})$-intersection over $A$ which is strongly approximable over $A$. Then there is a generic multiple $(\mathcal{Y},\mathcal{Z})$-intersection over $A$ which has codimension at most 1 over $A$.
Remark 6.17.
Our originally intended statement of Definition 6.16 was much simpler and closer to the usual statement: namely, that there is a notion of smoothness $X\mapsto X^{S}$ on $(\mathcal{K},\tau)$ such that, if a generically finite-to-one projection $f:X\rightarrow Y$ (where $\dim(X)=\dim(Y)$) topologically ramifies at a point $x$ with $x\in X^{S}$ and $y\in Y^{S}$, then there are codimension 1 topological ramification points of $f$ in any neighborhood of $x$. However, even in the specific example of ACVF, we were unable to prove that the ramification point we wish to consider is actually smooth in its domain. In a sense, one needs to replace ‘smooth varieties’ with ‘varieties universally homeomorphic to smooth varieties’ – and this has no obvious analog in the abstract setting. This is the reason we needed to use the more complicated definition above.
We now show:
Theorem 6.18.
Assume that $\mathcal{M}$ detects multiple intersections, and $(\mathcal{K},\tau)$ has ramification purity. Then $\dim(M)=1$.
Proof.
Let $\mathcal{C}=\{C_{t}:t\in T\}$ be an $\mathcal{M}(\emptyset)$-definable excellent family of plane curves. Let $\hat{t}\in T$ be generic, and let $\hat{x}\in C_{\hat{t}}$ be generic over $\hat{t}$. By Lemma 6.3, $(\hat{x},\hat{t},\hat{t})$ is a weakly generic $(\mathcal{C},\mathcal{C})$-intersection which is strongly approximable over $\emptyset$. On the other hand, we also have:
Claim 6.19.
$(\hat{x},\hat{t},\hat{t})$ is a generic multiple $(\mathcal{C},\mathcal{C})$-intersection over $\emptyset$.
Proof.
We need to show that $(\hat{x},\hat{t},\hat{t})$ is a topological ramification point of $I\rightarrow T^{2}$, where $I$ is the set of $(x,t,u)$ with $x\in C_{t}\cap C_{u}$. But $\hat{x}$ is generic in $C_{\hat{t}}$ over $\hat{t}$. By the Baire category axiom, it follows that every neighborhood of $\hat{x}$ contains infinitely many points of $C_{\hat{t}}$. In particular, every neighborhood of $(\hat{x},\hat{t},\hat{t})$ contains infinitely many points of $I$ which map to $(\hat{t},\hat{t})\in T^{2}$. This is enough to prove the claim. ∎
So $(\hat{x},\hat{t},\hat{t})$ is a generic multiple $(\mathcal{C},\mathcal{C})$-intersection which is strongly approximable. Moreover, $\mathcal{C}$ is a standard family over $\emptyset$ (since this is weaker than being excellent). So, since $(\mathcal{K},\tau)$ has ramification purity, there is a generic multiple $(\mathcal{C},\mathcal{C})$-intersection over $\emptyset$ which has codimension at most 1 over $\emptyset$. Call this intersection $(x,t,u)$. Note that, since $\mathcal{M}$ detects multiple intersections, the parameters $t$ and $u$ are $\mathcal{M}$-dependent over $\emptyset$.
We now have two bounds on $\dim(xtu)$ – a lower bound coming from the fact that $\operatorname{codim}_{\emptyset}(x,t,u)\leq 1$, and an upper bound coming from the $\mathcal{M}$-dependence of $t$ and $u$. The next two claims make these explicit:
Claim 6.20.
$\dim(xtu)\geq 4\cdot\dim(M)-1$.
Proof.
Using that $(x,t)$ and $(x,u)$ are generic in $C$, it is straightforward to see that $\dim(x)=2\cdot\dim(M)$, and $\dim(t/x)=\dim(u/x)=\dim(M)$. Since $\operatorname{codim}_{\emptyset}(x,t,u)=1$, this implies $\dim(tu/x)\geq 2\cdot\dim(M)-1$. Thus, by additivity,
$\dim(xtu)=\dim(x)+\dim(tu/x)\geq 4\cdot\dim(M)-1.$
Claim 6.21.
$\dim(xtu)\leq 3\cdot\dim(M)$.
Proof.
It is enough to show that $\operatorname{rk}(xtu)\leq 3$. Now since $t$ and $u$ are $\mathcal{M}$-dependent over $\emptyset$, we have $\operatorname{rk}(tu)\leq 3$. So if $\operatorname{rk}(x/tu)=0$, we are done. Otherwise, we get that $C_{t}\cap C_{u}$ is infinite. By almost faithfulness, $u$ is thus $\mathcal{M}$-algebraic over $t$, which gives that $\operatorname{rk}(xtu)=\operatorname{rk}(xt)=3$. ∎
The theorem follows by combining the two bounds above. Namely, by the two previous claims, we have $4\cdot\dim(M)-1\leq 3\cdot\dim(M)$. Rearranging gives $\dim(M)\leq 1$, and since $M$ is infinite, we get $\dim(M)=1$. ∎
7. Definable Slopes and Interpreting a Group
Assumption 7.1.
Throughout Section 7, in addition to Assumption 4.1, we assume that $\dim(M)=1$.
Our goal in this section is to give conditions under which $\mathcal{M}$ interprets a strongly minimal group. Essentially, the conditions say that the germ of a curve at a generic point is determined by a sequence of definable approximations, each one of which adds one parameter over the previous one – and that, moreover, these approximations can be ‘detected’ by $\mathcal{M}$ in a certain sense. This setting is an abstraction of power series expansions, where the definable approximations are Taylor polynomials.
Once we have our definable ‘Taylor expansions’, we then build a group in a similar way to previous trichotomy papers (e.g. [HS24] and [Cas23]) – by considering a collection of ‘slopes’ equipped with a composition operation. As in [Cas23], we find it most convenient to work with groupoids rather than groups, because this allows us to work solely with generic points of curves. In the end, this is still sufficient to recover a group, by Hrushovski’s group configuration theorem (see Fact 7.3).
7.1. Group Configurations
Our main tool for interpreting a group is Hrushovski’s group configuration theorem, which is well-known in model theory. Let us now make explicit the version of this theorem we will use:
Definition 7.2.
Let $A$ be a parameter set, and let
$\bar{s}=(s_{12},s_{13},s_{14},s_{23},s_{24},s_{34})$
be tuples in $\mathcal{M}^{\textrm{eq}}$. We say that $\bar{s}$ is a rank 1 group configuration in $\mathcal{M}$ over $A$ if the following hold:
  1. (1)
    Each of the six points in $\bar{s}$ has rank 1 over $A$.
  2. (2)
    Any two distinct points in $\bar{s}$ have rank 2 over $A$.
  3. (3)
    $\operatorname{rk}(\bar{s}/A)=3$.
  4. (4)
    If $i,j,k\in\{1,2,3,4\}$ with $i<j<k$ then $\operatorname{rk}(s_{ij}s_{ik}s_{jk}/A)=2$.
The canonical example is as follows: suppose $\mathcal{M}$ is a strongly minimal expansion of a group. Let $a,b,c\in M$ be independent generics, and let $s_{12}=a$, $s_{23}=b$, $s_{34}=c$, $s_{13}=ba$, $s_{24}=cb$, and $s_{14}=cba$. Then $\bar{s}$ is a rank 1 group configuration in $\mathcal{M}$ over $\emptyset$.
The main fact we need is:
Fact 7.3 (Hrushovski (see [Bou89])).
Suppose there is a rank 1 group configuration $\bar{s}$ in $\mathcal{M}$ over $A$ as above. Then there are
  • $B\supset A$,
  • an $\mathcal{M}(B)$-interpretable group $G$ which is strongly minimal as an interpretable set in $\mathcal{M}$, and
  • For each $i,j\in\{1,2,3\}$ with $i<j$, an $\mathcal{M}$-generic element $g_{ij}$ over $B$,
such that:
  1. (1)
    $g_{12}$ and $g_{23}$ are $\mathcal{M}$-independent over $B$,
  2. (2)
    $g_{13}=g_{23}g_{12}$, and
  3. (3)
    For each $i,j\in\{1,2,3\}$ with $i<j$, $g_{ij}$ and $s_{ij}$ are $\mathcal{M}$-interalgebraic over $B$.
7.2. Definable Slopes
We now define our abstraction of Taylor expansions, as described above: namely, we introduce the notion of $(\mathcal{K},\tau)$ having ‘definable slopes’. This will involve several intermediate notions.
Definition 7.4.
Let $\mathcal{LHG}$ be the category of local homeomorphisms at generic points in $(\mathcal{K},\tau)$, defined as follows:
  1. (1)
    An object in $\mathcal{LHG}$ is a generic point of $K$ (over $\emptyset$).
  2. (2)
    A morphism $x\rightarrow y$ is a germ of a local homeomorphism from $K$ to $K$ sending $x$ to $y$: that is, we consider homeomorphisms between neighorhoods of $x$ and $y$, modulo agreement in a neighborhood of $x$.
Definition 7.5.
Let $f:x\rightarrow y$ be a morphism in $\mathcal{LHG}$. We call $f$ a basic invertible arc if there are a set $X\subset K^{2}$ and a parameter set $A$ such that:
  1. (1)
    $X$ is $\mathcal{K}(A)$-definable.
  2. (2)
    $\dim(X)=1$, and each projection $X\rightarrow K$ is finite-to-one.
  3. (3)
    $(x,y)$ is generic in $X$ over $A$.
  4. (4)
    There are neighborhoods $U$ and $V$ of $x$ and $y$, respectively, so that $X\cap(U\times V)$ is the graph of a homeomorphism of germ $f$.
Definition 7.6.
Let $f:x\rightarrow y$ be a morphism in $\mathcal{LHG}$. We call $f$ an invertible arc if it is a composition of finitely many basic invertible arcs.
Suppose $X\subset K^{2}$ is $\mathcal{K}(A)$-definable and of dimension 1, with each projection $X\rightarrow K$ finite-to-one, and let $(x,y)\in X$ be generic over $A$. It follows from the generic local homeomorphism property that $X$ determines a basic invertible arc $x\rightarrow y$.
Notation 7.7.
Let $x$, $y$, $X$, and $A$ be as above. We call the basic invertible arc $x\rightarrow y$ determined by $X$ the arc of $X$ at $(x,y)$, denoted $\alpha_{xy}(X)$.
It is easy to see that for each generic $x\in K$, the identity morphism $\operatorname{id}:x\rightarrow x$ from $\mathcal{LHG}$ is a basic invertible arc (witnessed by the diagonal in $K^{2}$). Moreover, by switching the roles of the domain and target, one sees that basic invertible arcs are closed under inverses. It follows that the invertible arcs form the morphisms of a groupoid (in fact a subgroupoid of $\mathcal{LHG}$), with the same objects as $\mathcal{LHG}$.
Definition 7.8.
We let $\mathcal{IA}_{\infty}$ be the groupoid of invertible arcs.
Remark 7.9.
In contrast, the basic invertible arcs typically do not form a groupoid (that is, not every invertible arc is a basic invertible arc). For example, for generic $x\in K$, the identity is the only basic invertible arc $x\rightarrow x$ (as is easily checked from the definition); while other invertible arcs $x\rightarrow x$ can typically be constructed as compositions of basic invertible arcs $x\rightarrow y\rightarrow x$ where $y$ is generic over $x$. In a sense, then, the reason we need compositions is to recover the full collection of morphisms at points $(x,y)$ where $x$ and $y$ are dependent.
Note that $\mathcal{IA}_{\infty}$ is typically not a definable object in any reasonable sense. In contrast, the main point of our setting of ‘Taylor approximations’ is to require that $\mathcal{IA}_{\infty}$ can be approximated by definable objects:
Definition 7.10.
We say that $(\mathcal{K},\tau)$ has definable slopes if there are groupoids $\{\mathcal{I}A_{n}\}_{n\in\mathbb{Z}^{+}}$, and covariant functors $T_{m,n}:\mathcal{IA}_{m}\rightarrow\mathcal{IA}_{n}$ for all $n<m\in\mathbb{Z}^{+}\cup\{\infty\}$, such that:
  1. (1)
    The objects of each $\mathcal{IA}_{n}$ are the same as in $\mathcal{IA}_{\infty}$, and all $T_{m,n}$ are the identity on objects.
  2. (2)
    The $T_{m,n}$ are compatible, and $\mathcal{IA}_{\infty}$ is the inverse limit of the $\mathcal{IA}_{n}$ for $n<\infty$. More precisely, if $m>n>k$ then $T_{mk}=T_{nk}\circ T_{nm}$, and for any two objects $x$ and $y$, the $T_{m,n}$ turn the set of morphisms $x\rightarrow y$ in $\mathcal{IA}_{\infty}$ into the inverse limit of the sets of morphisms $x\rightarrow y$ in the $\mathcal{IA}_{n}$ for $n<\infty$.
  3. (3)
    Let $f:x\rightarrow y$ be a morphism in $\mathcal{IA}_{n}$ for some $n<\infty$. Then $f$ is a tuple in $\mathcal{K}^{eq}$, and $(x,y)$ is $\mathcal{K}(f)$-definable.
  4. (4)
    (Uniform Definability of Slopes) Let $T$ and $X\subset K^{2}\times T$ be $\mathcal{K}(A)$-definable. Assume that for each $t\in T$, the fiber $X_{t}\subset K^{2}$ has dimension 1 and projects finite-to-one to both copies of $K$. Then for each $n$, the set of $(x,y,t,\alpha)$ such that $(x,y)$ is generic in $X_{t}$ over $At$ and $T_{\infty,n}(\alpha_{xy}(X_{t}))=\alpha$, is type-definable in $\mathcal{K}$ over $A$.
  5. (5)
    (Definability of Composition and Inverse) If $f:x\rightarrow y$ and $g:y\rightarrow z$ are morphisms in $\mathcal{IA}_{n}$ for some $n<\infty$, then $g\circ f$ is $\mathcal{K}(f,g)$-definable, and $f^{-1}$ is $\mathcal{K}(f)$-definable.
  6. (6)
    (Coordinatization of Slopes) Suppose $f:x\rightarrow y$ is a morphism in $\mathcal{IA}_{n+1}$ for some $n<\infty$, and let $g=T_{n+1,n}(f)$. Then $g$ is $\mathcal{K}(f)$-definable, and $\dim(f/g)\leq 1$.
Remark 7.11.
Definition 7.10(2) (particularly the clause about inverse limits) is a strong assumption: it says that the germ of a curve at a generic point is determined by its sequence of $n$th approximations (at that point) for every $n$. For instance, this property holds in ACVF and in polynomially bounded o-minimal structures, but we don’t know if it holds in arbitrary o-minimal structures (because of the existence of ‘flat functions’ – non-constant functions with all derivatives vanishing at some point).
On the other hand, we will only use the inverse limit property once (in the proof of Lemma 7.22), to guarantee the existence of an infinite family of $n$-slopes for some $n$. We note, then, that the statement of Lemma 7.22 can be proven in many other settings (including all o-minimal fields) using a version of Sard’s Theorem. For this to work, one just needs a differentiable Hausdorff geometric structure whose differential structures is definable in an appropriate sense.
Finally, we note that one could amend all of Section 7 accordingly in this case – taking Lemma 7.22 as an axiom instead of the inverse limit clause of (2) – and all the arguments would still work. The reason we don’t do this is that Lemma 7.22 rather awkwardly involves the relic $\mathcal{M}$, and we prefer Definition 7.10 to be purely a property of $(\mathcal{K},\tau)$.
Assumption 7.12.
For the rest of Section 7, we assume that $(\mathcal{K},\tau)$ has definable slopes, witnessed by the fixed groupoids and functors $\mathcal{IA}_{n}$ and $T_{m,n}$.
Remark 7.13.
We think of the functors $T_{m,n}$ as truncation maps. Given a morphism $f$ in $\mathcal{IA}_{m}$ for some $m$, and given some $n<m$, we often call $T_{m,n}(f)$ the $n$th truncation of $f$.
Definition 7.14.
Assume that $(\mathcal{K},\tau)$ has definable slopes, witnessed by $\{\mathcal{IA}_{n}\}$ and $\{T_{m,n}\}$. Let $X\subset K^{2}$ be $\mathcal{K}(A)$-definable of dimension 1, with both projections $X\rightarrow K$ finite-to-one, and let $(x,y)\in X$ be generic over $A$. For $n<\infty$, we call $T_{\infty,n}(\alpha_{xy}(X))$ the $n$-slope of $X$ at $(x,y)$, denoted $\alpha_{xy}^{n}(X)$.
In this language, by an easy compactness argument, the uniform definability of slopes (Definition 7.10(4)) has the following simpler consequence:
Lemma 7.15.
Let $X\subset K^{2}$ be $\mathcal{K}(A)$-definable of dimension 1, with both projections finite-to-one. Let $(x,y)\in X$ be generic over $A$. Then for each $n<\infty$, $\alpha_{xy}^{n}(X)\in\operatorname{dcl}_{\mathcal{K}}(Axy)$.
Remark 7.16.
Definition 7.10(4) is stronger than Lemma 7.15, because Lemma 7.15 does not capture the uniformity of the way slopes are defined. This will matter in a couple of places, but usually Lemma 7.15 will suffice.
7.3. Coherent Slopes and Codes
In this subsection, we give axioms for an encoding system of slopes into $\mathcal{M}$. The idea is to assign each slope to an interalgebraic tuple in $\mathcal{M}$, so that the model-theoretic properties of $\mathcal{K}$ appearing in Definition 7.10 can be replaced with the analogous properties in $\mathcal{M}$. For this to have a chance of working, we first need to assign slopes to plane curves in $M$ (rather than sets in $K^{2}$). We then need to identify a class of coherent slopes – roughly, those appearing on coherently defined plane curves.
Let us begin by ‘lifting’ slopes to $M$. Since $\dim(M)=1$, and $M$ is $\mathcal{K}(\emptyset)$-definable, it follows that there is a finite-to-one $\mathcal{K}(\emptyset)$-definable function $\rho:M\rightarrow K$ (one can construct such a function piecewise by projections). For the rest of Section 7, we fix such a function $\rho$. Abusing notation, we also write $\rho:M^{n}\rightarrow K^{n}$ for all $n$, applying coordinate-wise.
Suppose $C\subset M^{2}$ is $\mathcal{K}(A)$-definable of dimension 1 with both projections $C\rightarrow M$ finite-to-one. Let $(x,y)\in C$ be generic over $A$. Then $\rho(C)\subset K^{2}$ is $A$-definable and projects finite-to-one in both directions, and $(\rho(x),\rho(y))$ is generic in $\rho(C)$ over $A$. By the generic local homeomorphism property, $\rho$ induces a homeomorphism between open subsets of $M^{2}$ and $K^{2}$ near $(x,y)$ and $(\rho(x),\rho(y))$, and also between $C$ and $\rho(C)$ near $(x,y)$ and $(\rho(x),\rho(y))$. This observation allows us, in a neighborhood of $(x,y)$, to think of $C$ as a curve in $K^{2}$, and subsequently to extract slopes at $(x,y)$:
Definition 7.17.
Let $C\subset M^{2}$, $A$, and $(x,y)$ be as in the above paragraph. For each $n$, we define the $n$-slope of $C$ at $(x,y)$, denoted $\alpha_{xy}^{n}(X)$, to be the $n$-slope of $\rho(C)$ at $(\rho(x),\rho(y))$.
We now proceed to discuss coherent slopes in $M$. Recall (see Remark 5.3) that a plane curve $C\subset M^{2}$ is non-trivial if each projection $C\rightarrow M$ is finite-to-one. Now we define:
Definition 7.18.
Let $n$ be a positive integer. By a coherent $n$-slope configuration, we mean a 5-tuple $(x,y,C,c,\alpha)$ such that:
  1. (1)
    $C\subset M^{2}$ is a stationary non-trivial plane curve in the sense of $\mathcal{M}$ (so in particular $C$ is $\mathcal{M}$-definable), and $c=\operatorname{Cb}(C)$.
  2. (2)
    $c$ is coherent, and $(x,y)$ is generic in $C$ over $c$ (so $cxy$ is coherent).
  3. (3)
    $\alpha_{xy}^{n}(C)=\alpha$.
Definition 7.19.
Let $x,y\in M$, and let $\alpha:\rho(x)\rightarrow\rho(y)$ be a morphism in $\mathcal{IA}_{n}$ for some $n<\infty$.
  1. (1)
    We say that $\alpha$ is a coherent $n$-slope at $(x,y)$ if there is a coherent $n$-slope configuration $(x,y,C,c,\alpha)$.
  2. (2)
    If $(x,y,C,c,\alpha)$ is a coherent $n$-slope configuration, we say that $c$ is a coherent representative of $\alpha$.
  3. (3)
    Suppose $\alpha$ is a coherent $n$-slope at $(x,y)$. If $\alpha\in\operatorname{acl}(xy)$, we call $\alpha$ an algebraic coherent $n$-slope at $(x,y)$. Otherwise, we call $\alpha$ a non-algebraic coherent $n$-slope at $(x,y)$.
We now give various properties of coherent slopes. Lemma 7.20 is essentially a restatement of Definition 7.10 (3) and (4), but is quite useful. Roughly, it allows us to perform dimension computations with coherent slopes, while (in the presence of any coherent representative) treating the slopes as elements of $\mathcal{M}^{\textrm{eq}}$.
Lemma 7.20.
Let $\alpha$ be a coherent $n$-slope at $(x,y)$, and let $c$ be any coherent representative of $\alpha$.
  1. (1)
    $(x,y)\in\operatorname{acl}(\alpha)$.
  2. (2)
    $\alpha$ is definable over $cxy$ (that is, $\alpha\in\operatorname{dcl}_{\mathcal{K}}(cxy)$).
  3. (3)
    $\alpha$ and $xy$ are interalgebraic over $c$.
Proof.
Clause (2) follows from Lemma 7.15 and the fact that $\rho$ is $\mathcal{K}(\emptyset)$-definable. Clause (3) follows from (1) and (2). For (1), since $\alpha$ is a morphism from $\rho(x)$ to $\rho(y)$, Definition 7.10(3) gives that $\rho(x)$ and $\rho(y)$ are $\mathcal{K}(\alpha)$-definable. Meanwhile, since $\rho$ is finite-to-one, $x$ and $y$ are algebraic over $(\rho(x),\rho(y))$. ∎
It will also be convenient to know that, when working with non-algebraic slopes, we can upgrade Definition 7.19 to include the genericity of $(x,y)$ in $M^{2}$:
Lemma 7.21.
Let $\alpha$ be a c coherent $n$-slope at $(x,y)$. Then:
  1. (1)
    If $\alpha\notin\operatorname{acl}(x,y)$, then $(x,y)$ is generic in $M^{2}$.
  2. (2)
    If $(x,y)$ is generic in $M^{2}$, and $c$ is any coherent representative of $\alpha$, then $\dim(c/xy)=\dim(c)-1$.
Proof.
For (1), suppose toward a contradiction that $(x,y)$ is not generic in $M^{2}$. Then by coherence, $\operatorname{rk}(xy)\leq 1$. But $\operatorname{rk}(xy/c)=1$ by assumption, so $\operatorname{rk}(xy/c)=\operatorname{rk}(xy)=1$. So $p=\operatorname{tp}_{\mathcal{M}}(xy/c)$ does not fork over $\emptyset$. This implies that $c=\operatorname{Cb}(p)\in\operatorname{acl}_{\mathcal{M}}(\emptyset)$. Then by Lemma 7.20, $\alpha\in\operatorname{acl}(cxy)\subset\operatorname{acl}(xy)$, so $\alpha$ is algebraic.
For (2), let $c$ be any coherent representative. Then $\dim(xy/c)=1$, so $\dim(cxy)=\dim(c)+1$. On the other hand, $\dim(xy)=2$ by assumption. So
$\dim(c/xy)=(\dim(c)+1)-2=\dim(c)=1.$
Later, under further assumptions, we will interpret a strongly minimal group in $\mathcal{M}$ by finding codes for a one-dimensional family of coherent slopes. Lemmas 7.22 and 7.25 guarantee that such one-dimensional families exist.
Lemma 7.22.
There is a non-algebraic coherent $n$-slope for some $n$.
Proof.
Fix any stationary non-trivial plane curve $C$ with coherent canonical base $c$, and fix a generic element $(x,y)\in C$ over $c$. So for each $n<\infty$, $\alpha_{xy}^{n}(C)$ is a coherent $n$-slope at $(x,y)$.
Since the category $\mathcal{IA}_{\infty}$ is the inverse limit of the $\mathcal{IA}_{n}$ for $n<\infty$ (see Definition 7.10(2)), the basic invertible arc $\alpha_{\rho(x)\rho(y)}(\rho(C))$ (i.e. the germ of $\rho(C)$ at $(\rho(x),\rho(y))$) is determined by the truncations $\alpha_{xy}^{n}(C)$ for $n<\infty$. Since $\rho$ induces a local homeomorphism $C\rightarrow\rho(C)$ near $(x,y)$, it follows that these truncations, together with $(x,y)$, determine the germ of $C$ at $(x,y)$.
Let $A$ be the tuple of all $\alpha_{xy}^{n}(C)$ for $n<\infty$. It now follows that $c=\operatorname{Cb}(c)$ is determined by $A$ – or more precisely, $c\in\operatorname{dcl}_{\mathcal{K}}(A)$. Let us explain. By $\aleph_{1}$-saturation, it is enough to show that for any $c^{\prime}\models\operatorname{tp}_{\mathcal{K}}(c/A)$, we have $c^{\prime}=c$. So fix such $c^{\prime}$. Then $c^{\prime}=\operatorname{Cb}(C^{\prime})$ for some stationary non-trivial plane curve $C^{\prime}$, also containing $(x,y)$ as a generic point. Now by uniform definability of slopes, and since $\operatorname{tp}_{\mathcal{K}}(c^{\prime}/A)=\operatorname{tp}_{\mathcal{K}}(c/A)$, it follows that $\alpha_{xy}^{n}(C^{\prime})=\alpha_{xy}^{n}(C)$ for all $n<\infty$. Then by the discussion from the above paragraph, $C^{\prime}$ agrees with $C$ in a neighborhood of $(x,y)$. But by the Baire category axiom (see Remark 3.7), every neighborhood of $(x,y)$ in $C$ is infinite; so in fact $C^{\prime}$ and $C$ have infinite intersection, and thus by strong minimality they are almost equal. Thus, they have the same canonical base, i.e. $c^{\prime}=c$.
Now assume that all coherent slopes are algebraic. Then, in the notation above, we get $A\subset\operatorname{acl}(xy)$, and thus (by the previous paragraph) $c\in\operatorname{acl}(xy)$. In particular, $\dim(c)\leq 2$. But, by non-local modularity, $\dim(c)$ can be arbitrarily large, which is enough to prove the lemma. ∎
Notation 7.23.
For the rest of section 7, we let $n_{0}$ be the smallest $n$ such that there is a non-algebraic coherent $n$-slope.
Remark 7.24.
One could also define ‘non-algebraic coherent $n$-slopes’ relative to a parameter set $A$ (so coherence is interpreted over $A$, and non-algebraicity at $(x,y)$ means the slope is not algebraic over $Axy$). This would, a priori, change the way that $n_{0}$ is defined. We note, however, that the value of $n_{0}$ does not ultimately depend on parameters (which the reader can check if desired). Thus, we find it simplest to stick to the parameter-free version presented above.
Lemma 7.25.
Suppose $\alpha$ is a non-algebraic coherent $n_{0}$-slope at $(x,y)$.
  1. (1)
    $\dim(\alpha/xy)=1$.
  2. (2)
    If $c$ is any coherent representative of $\alpha$, then $\dim(c/\alpha)=\dim(c/xy)-1$.
Proof.
Let $\beta$ be the $(n_{0}-1)$-th truncation of $\alpha$. By the minimality of $n_{0}$, $\beta\in\operatorname{acl}(xy)$. By Definition 7.10(6), $\dim(\alpha/\beta)\leq 1$. Thus $\dim(\alpha/xy)\leq 1$, and since $\alpha$ is non-algebraic, $\dim(\alpha/xy)=1$, proving (1).
For (2), let $c$ be any coherent representative of $\alpha$. Repeatedly using Lemma 7.20 and (1), we compute:
$\dim(c/\alpha)=\dim(c/\alpha xy)=\dim(c\alpha/xy)-\dim(\alpha/xy)=\dim(c/xy)-1.$
Finally, we now end this subsection with our definition of codes of slopes in $\mathcal{M}$. This is the key notion allowing us to recover composition of slopes in $\mathcal{M}$.
Definition 7.26.
Let $\alpha$ be a coherent $n$-slope at $(x,y)$. A tuple $s$ from $\mathcal{M}^{\textrm{eq}}$ is a code of $\alpha$ if the following hold:
  1. (1)
    $(x,y)\in\operatorname{acl}_{\mathcal{M}}(s)$.
  2. (2)
    For each coherent representative $c$ of $\alpha$, $s\in\operatorname{acl}_{\mathcal{M}}(cxy)$.
  3. (3)
    $\alpha$ and $s$ are $\mathcal{K}$-interalgebraic.
Note that codes might not exist – that is, if $\alpha$ is a coherent $n$-slope at $(x,y)$, there is no reason in general to expect some $s\in\mathcal{M}^{eq}$ to satisfy (1)-(3) in Definition 7.26. Moving forward, one of our main jobs will be to find suitable topological conditions guaranteeing that many slopes do have codes.
7.4. Detecting Composition when Codes Exist
Definition 7.10(5) says that slope composition is definable in an appropriate sense. Our next goal is to transfer this property into $\mathcal{M}$, for ‘independent’ composition of coherent slopes that have codes. The idea is that the composition of slopes is respected by the $\mathcal{M}$-definable composition of plane curves.
The following notion will be useful in the next two sections:
Definition 7.27.
Let $a_{1},...,a_{n},b_{1},...,b_{n}\in\mathcal{K}^{\textrm{eq}}$. We say that $a_{1},...,a_{n}$ are maximally independent over $b_{1},...,b_{n}$ if
$\dim(a_{1}...a_{n}/b_{1}...b_{n})=\dim(a_{1}/b_{1})+...+\dim(a_{n}/b_{n}).$
Note that if $b_{1}=...=b_{n}=b$, this is the same as saying that $a_{1},...,a_{n}$ are independent over $b$.
The following transitivity property follows easily from additivity of dimension:
Lemma 7.28.
$a_{1}b_{1},...,a_{n}b_{n}$ are maximally independent over $c_{1},...,c_{n}$ if and only if $a_{1},...,a_{n}$ are maximally independent over $c_{1},...,c_{n}$ and $b_{1},...,b_{n}$ are maximally independent over $a_{1}c_{1}...a_{n}c_{n}$.
Proposition 7.29.
Let $\alpha_{1}$ be a coherent $n$-slope at $(x_{0},y_{0})$, and let $\alpha_{2}$ be a coherent $n$-slope at $(y_{0},z_{0})$. Assume that $x_{0}$ and $z_{0}$ are independent over $y_{0}$, and $\alpha_{1},\alpha_{2}$ are maximally independent over $(x_{0},y_{0}),(y_{0},z_{0})$. Then $\alpha_{3}=\alpha_{2}\circ\alpha_{1}$ is a coherent $n$-slope at $(x_{0},z_{0})$. Moreover, if $s_{i}$ is a code of $\alpha_{i}$ for $i=1,2,3$, then $s_{3}\in\operatorname{acl}_{\mathcal{M}}(s_{1}s_{2})$.
Proof.
Let $c_{1}$ and $c_{2}$ be coherent representatives of $\alpha_{1}$ and $\alpha_{2}$, respectively. We may assume that $c_{1},c_{2}$ are maximally independent over $\alpha_{1},\alpha_{2}$. Note also that $(x_{0},y_{0}),(y_{0},z_{0})$ are maximally independent over $y_{0},y_{0}$ by assumption. By repeated applications of Lemma 7.28, it then follows that $c_{1},c_{2}$ are maximally independent over $y_{0},y_{0}$. Thus, $c_{1}$ and $c_{2}$ are independent over $y_{0}$. But each $c_{i}$ is independent from $y_{0}$ by assumption. It then follows from additivity of dimension that $c_{1}$, $c_{2}$, and $y_{0}$ are independent over $\emptyset$. In particular, $c_{1}c_{2}y_{0}$ is coherent, and $\dim(y_{0}/c_{1}c_{2})=1$. Clearly $x_{0}$, $y_{0}$, and $z_{0}$ are pairwise $\mathcal{M}$-interalgebraic over $c_{1}c_{2}$. Thus, we have:
Claim 7.30.
$c_{1}c_{2}x_{0}y_{0}z_{0}$ is coherent, and $\dim(x_{0}z_{0}/c_{1}c_{2})=1$.
Now let $C_{1}$ and $C_{2}$ be such that $(x_{0},y_{0},C_{1},c_{1},\alpha_{1})$ and $(y_{0},z_{0},C_{2},c_{2},\alpha_{2})$ are coherent $n$-slope configurations. We may assume that each $C_{i}$ is $\mathcal{M}$-definable over $c_{i}$. Let $C_{3}=C_{2}\circ C_{1}$ – that is, the set of $(x,z)$ such that for some $y$ we have $(x,y)\in C_{1}$ and $(y,z)\in C_{2}$. Then $C$ is a non-trivial plane curve which is $\mathcal{M}(c_{1}c_{2})$-definable and (by Claim 7.30) contains $(x_{0},z_{0})$ as a generic element over $c_{1}c_{2}$.
Claim 7.31.
$\alpha_{x_{0}z_{0}}^{n}(C_{3})=\alpha_{3}$.
Proof.
Let $q_{1}$ stand for $(x_{0},y_{0})$, $q_{2}$ stand for $(y_{0},z_{0})$, and $q_{3}$ stand for $(x_{0},z_{0})$. By the generic local homeomorphism property, for each $i=1,2,3$, the points of $C_{i}$ in a neighborhood of $q_{i}$ form the graph of a local homeomorphism $f_{i}$. Shrinking these neighborhoods if necessary, it is clear that $f_{3}=f_{2}\circ f_{1}$.
Now, for each $i$, let $\rho(f_{i})$ be the image of $f_{i}$ under $\rho$, viewed as a local function on $K$ at $\rho(q_{i})$. It follows by definition that each $\rho(f_{i})$ is an invertible arc, with $n$th truncation $\alpha_{q_{i}}^{n}(C_{i})$. But since $\rho$ is a local homeomorphism $M\rightarrow K$ near each of $x_{0}$, $y_{0}$, and $z_{0}$, it also follows that $\rho(f_{3})=\rho(f_{2})\circ\rho(f_{1})$. So since the truncation maps are functors, we get $\alpha_{q_{3}}^{n}(C_{3})=\alpha_{q_{2}}^{n}(C_{2})\circ\alpha_{q_{1}}^{n}(C_{1})$. Equivalently, $\alpha_{x_{0}z_{0}}^{n}(C_{3})=\alpha_{2}\circ\alpha_{1}=\alpha_{3}$. ∎
Now since $(x_{0},z_{0})$ is generic in $C_{3}$, there is a strongly minimal component $C$ of $C_{3}$, with canonical base $c_{3}$, such that $(x_{0},z_{0})$ is generic in $C$ over $c_{3}$. Then $c_{3}\in\operatorname{acl}_{\mathcal{M}}(c_{1}c_{2})$, so $c_{3}$ is coherent. Moreover, by the strong frontier inequality, $C$ and $C_{3}$ agree in a neighborhood of $(x_{0},z_{0})$, which shows that $\alpha_{x_{0},z_{0}}^{n}(C)=\alpha_{x_{0},z_{0}}^{n}(C_{3})=\alpha_{3}$. It follows that $(x_{0},z_{0},C,c_{3},\alpha_{3})$ is a coherent $n$-slope configuration. So $\alpha_{3}$ is coherent, and $c_{3}$ is a coherent representative of $\alpha_{3}$.
Now assume that $s_{i}$ is a code of $\alpha_{i}$ for each $i$. By the definition of codes, we have $s_{i}\in\operatorname{acl}_{\mathcal{M}}(c_{i}x_{0}y_{0}z_{0})$ for each $i$. Since $c_{3}\in\operatorname{acl}_{\mathcal{M}}(c_{1}c_{2})$, this implies $s_{1}s_{2}s_{3}\in\operatorname{acl}_{\mathcal{M}}(c_{1}c_{2}x_{0}y_{0}z_{0})$. By Claim 7.30, $c_{1}c_{2}x_{0}y_{0}z_{0}$ is coherent, thus so is $s_{1}s_{2}s_{3}$. But by the definability of composition (Definition 7.10(5)), $\alpha_{3}\in\operatorname{acl}(\alpha_{1}\alpha_{2})$. So by interalgebraicity with the $s_{i}$ (Definition 7.26)(3), $s_{3}\in\operatorname{acl}(s_{1}s_{2})$, and thus by coherence, $s_{3}\in\operatorname{acl}_{\mathcal{M}}(s_{1}s_{2})$. ∎
Remark 7.32.
Note that the last clause of Proposition 7.29 requires that all three $\alpha_{i}$ have codes. That is, we do not claim that if $\alpha_{1}$ and $\alpha_{2}$ have codes, so does their composition.
7.5. Interpreting a Group when Codes Exist
Recall that $n_{0}$ is the smallest $n$ such that there is a non-algebraic coherent $n$-slope. We now use Proposition 7.29 to interpret a group in $\mathcal{M}$ assuming enough coherent $n_{0}$-slopes have codes. The statement of Proposition 7.33 below is somewhat complicated. The reason is that later on, in ACVF, we will need to analyze the specific structure of the interpreted group – a task which will use the extra data in the proposition. However, we stress the main point: if all coherent $n_{0}$-slopes have codes, then $\mathcal{M}$ interprets a strongly minimal group.
Proposition 7.33.
Suppose every coherent $n_{0}$-slope has a code. Then there are
  • A parameter set $B$,
  • An $\mathcal{M}(B)$-interpretable group $(G,\cdot)$ which is strongly minimal as in interpretable set in $\mathcal{M}$,
  • A generic element $y\in K$ over $\emptyset$,
  • For each $i,j\in\{1,2,3\}$ with $i<j$, an element $g_{ij}\in G$, and
  • For each $i,j\in\{1,2,3\}$ with $i<j$, a morphism $f_{ij}:y\rightarrow y$ in $\mathcal{IA}_{n_{0}}$,
such that:
  1. (1)
    $y$ is $\mathcal{K}(B)$-definable (that is, $y\in\operatorname{dcl}_{\mathcal{K}}(B)$).
  2. (2)
    The $(n_{0}-1)$-th truncation of each $f_{ij}$ is the identity at $y$.
  3. (3)
    $g_{12}$ and $g_{23}$ are independent generics in $G$ over $B$.
  4. (4)
    $g_{13}=g_{23}\cdot g_{12}$ and $f_{13}=f_{23}\circ f_{12}$.
  5. (5)
    For each $i,j\in\{1,2,3\}$ with $i<j$, $g_{ij}$ and $f_{ij}$ are interalgebraic over $B$.
Proof.
Our strategy is as follows. First, we use the definability of composition and inverse of slopes to find a ‘group configuration’ in the sense of dimension in $\mathcal{K}$, whose elements are non-algebraic $n_{0}$-slopes. Then we use Proposition 7.29 to check that the slopes in the configuration are coherent, and subsequently extract codes. Finally, we check that the tuple of codes in the entire configuration is coherent, and conclude that the codes form a group configuration in the sense of rank in $\mathcal{M}$.
Let us proceed. We first want to find an appropriate collection of points and slopes to form a group configuration. The first two lemmas below accomplish this.
Lemma 7.34.
Suppose there is a non-algebraic coherent $n_{0}$-slope at $(x_{0},y_{0})$ for some $(x_{0},y_{0})\in M^{2}$. Then there is a non-algebraic coherent $n_{0}$-slope at $(y_{0},x_{0})$.
Proof.
Suppose that $(x_{0},y_{0},C,c,\alpha)$ is a non-algebraic coherent $n_{0}$-slope configuration. It is then easy to check that $(y_{0},x_{0},C^{-1},c,\alpha^{-1})$ is a non-algebraic coherent $n_{0}$-slope configuration, where $C^{-1}$ is the set of $(y,x)$ with $(x,y)\in C$. ∎
Lemma 7.35.
There are independent generic elements $x_{1},x_{2},x_{3},x_{4}\in M$ such that each of $(x_{1},x_{2})$, $(x_{2},x_{3})$, and $(x_{3},x_{4})$ admits a non-algebraic coherent $n_{0}$-slope.
Proof.
Let $(x_{1},x_{2})$ be such that there is a non-algebraic coherent $n_{0}$-slope at $(x_{1},x_{2})$. By Lemma 7.21, $(x_{1},x_{2})$ is generic in $M^{2}$. Let $x_{3}$ be a realization of $\operatorname{tp}_{\mathcal{K}}(x_{1}/x_{2})$ independent over $x_{1}$, and let $x_{4}$ be a realization of $\operatorname{tp}_{\mathcal{K}}(x_{2}/x_{3})$ independent over $x_{1}x_{2}$. So $x_{1},x_{2},x_{3},x_{4}$ are independent generics in $M$, and all of $(x_{1},x_{2})$, $(x_{3},x_{2})$, and $(x_{3},x_{4})$ realize the same type in $\mathcal{K}$. Thus, each of $(x_{1},x_{2})$, $(x_{3},x_{2})$, and $(x_{3},x_{4})$ admits a non-algebraic coherent $n_{0}$-slope, and by Lemma 7.34, so does $(x_{2},x_{3})$. ∎
Notation 7.36.
For the rest of the proof of Proposition 7.33, we adopt the following:
  • We fix $x_{1},x_{2},x_{3},x_{4}$ as provided by Lemma 7.35.
  • We fix non-algebraic coherent $n_{0}$-slopes $\alpha_{12}$ at $(x_{1},x_{2})$, $\alpha_{23}$ at $(x_{2},x_{3})$, and $\alpha_{34}$ at $(x_{3},x_{4})$, chosen so that $\alpha_{12},\alpha_{23},\alpha_{34}$ are maximally independent over $x_{1}x_{2},x_{2}x_{3},x_{3}x_{4}$.
  • Finally, we let $\alpha_{13}=\alpha_{23}\circ\alpha_{12}$, $\alpha_{24}=\alpha_{34}\circ\alpha_{23}$, and $\alpha_{14}=\alpha_{34}\circ\alpha_{23}\circ\alpha_{12}$, and we denote the tuple $(\alpha_{12},\alpha_{23},\alpha_{34},\alpha_{13},\alpha_{24},\alpha_{14})$ by $\overline{\alpha}$.
Lemma 7.37 follows easily from Lemmas 7.25 and the definability of composition and inverse (Definition 7.10(5)). We leave the details to the reader.
Lemma 7.37.
$\overline{\alpha}$ forms a ‘one-dimensional group configuration in $\mathcal{K}$ over $x_{1}x_{2}x_{3}x_{4}$’. That is:
  1. (1)
    Each point in $\overline{\alpha}$ has dimension 1 over $x_{1}x_{2}x_{3}x_{4}$.
  2. (2)
    Any two distinct points in $\overline{\alpha}$ have dimension 2 over $x_{1}x_{2}x_{3}x_{4}$.
  3. (3)
    $\dim(\overline{\alpha}/x_{1}x_{2}x_{3}x_{4})=3$.
  4. (4)
    If $i,j,k\in\{1,2,3,4\}$ with $i<j<k$, then $\dim(\alpha_{ij}\alpha_{jk}\alpha_{ik}/x_{1}x_{2}x_{3}x_{4})=2$.
By Lemmas 7.25 and 7.37, any two elements $\alpha_{ij},\alpha_{kl}\in\overline{\alpha}$ are maximally independent over $x_{i}x_{j},x_{k}x_{l}$. This allows us to apply Proposition 7.29 to the elements of $\overline{\alpha}$. We first conclude:
Lemma 7.38.
Each of the six elements of $\overline{\alpha}$ is a coherent $n_{0}$-slope.
Proof.
$\alpha_{12}$, $\alpha_{23}$, and $\alpha_{14}$ are coherent by the way they are chosen. Now repeatedly apply Proposition 7.29. We get first that each of $\alpha_{13}$ and $\alpha_{24}$ is a composition of two maximally independent coherent slopes, so is coherent. Finally, $\alpha_{14}$ is now coherent by the same reasoning, by writing it as $\alpha_{34}\circ\alpha_{13}$. ∎
By Lemma 7.38 and the assumption of the proposition, each $\alpha_{ij}$ has a code.
Notation 7.39.
We fix a code $s_{ij}$ for each $\alpha_{ij}$, and denote the tuple of $s_{ij}$ by $\overline{s}$.
Arguing similarly to Lemma 7.38, we now conclude:
Lemma 7.40.
$\overline{s}$ is coherent.
Proof.
By definition of codes, each $s_{ij}$ is interalgebraic with $\alpha_{ij}$. It follows that $s_{12},s_{23},s_{34}$ are maximally independent over $x_{1}x_{2},x_{2}x_{3},x_{3}x_{4}$. It then follows easily that $s_{12}s_{23}s_{34}$ is coherent over $x_{1}x_{2}x_{3}x_{4}$, and since $x_{1}x_{2}x_{3}x_{4}$ is coherent by construction, we get that $s_{12}s_{23}s_{34}$ is coherent. But by repeated instances of Proposition 7.29 (exactly as in Lemma 7.38), we have $\overline{s}\in\operatorname{acl}_{\mathcal{M}}(s_{12}s_{23}s_{34})$. So $\overline{s}$ is coherent. ∎
Now, since each $s_{ij}$ is interalgebraic with $\alpha_{ij}$, Lemma 7.37 holds with each $\overline{\alpha}$ replaced with $\overline{s}$. Since $\overline{s}$ is coherent, all of the statements in Lemma 7.37 hold with rank instead of dimension. Thus, $\overline{s}$ is a rank 1 group configuration in $\mathcal{M}$ over $x_{1}x_{2}x_{3}x_{4}$, and thus, by Fact 7.3, $\mathcal{M}$ interprets a strongly minimal group.
If we only wanted to find a group, we could stop here. However, let us now apply Fact 7.3 more carefully in order to verify the extra conditions of Proposition 7.33. First, we define:
  • Let $\overline{\beta}\overline{t}$ be an independent realization of $\operatorname{tp}_{\mathcal{K}}(\overline{\alpha}\overline{s}/\operatorname{acl}(x_{1}x_{2}x_{3}x_{4}))$ over $\overline{\alpha}\overline{s}$ (by this we mean that for $i<j$ we have distinguished tuples $\beta_{ij}$ and $t_{ij}$).
  • Let $A=\overline{t}x_{1}x_{2}x_{3}x_{4}$.
  • Let $y=\rho(x_{4})$.
  • For $i,j\in\{1,2,3\}$ with $i<j$, let $f_{ij}=\beta_{j4}\circ\alpha_{ij}\circ\beta_{i4}^{-1}$.
We now check:
Lemma 7.41.
  1. (1)
    $y$ is generic in $K$ over $\emptyset$.
  2. (2)
    $y$ is $\mathcal{K}(A)$-definable.
  3. (3)
    Each $\alpha_{ij}$ has the same $(n_{0}-1)$-th truncation as $\beta_{ij}$
  4. (4)
    Each $f_{ij}$ is a morphism from $y$ to $y$ in $\mathcal{IA}_{n_{0}}$.
  5. (5)
    The $(n_{0}-1)$-th truncation of each $f_{ij}$ is the identity at $y$.
  6. (6)
    Each $f_{ij}$ is interalgebraic with $s_{ij}$ over $A$.
  7. (7)
    $f_{13}=f_{23}\circ f_{12}$.
  8. (8)
    $\overline{s}$ is a rank 1 group configuration in $\mathcal{M}$ over $A$.
Proof.
  1. (1)
    Clear, because $y$ and $x_{4}$ are interalgebraic.
  2. (2)
    Clear because $\rho$-is $\emptyset$-definable.
  3. (3)
    By the minimality of $n_{0}$, the $(n_{0}-1)$-th truncation of $\alpha_{ij}$ is contained in $\operatorname{acl}(x_{1}x_{2}x_{3}x_{4})$, and $\alpha_{ij}$ and $\beta_{ij}$ realize the same type over $\operatorname{acl}(x_{1}x_{2}x_{3}x_{4})$.
  4. (4)
    Clear, because $f_{ij}$ is the composition of $\beta_{i4}^{-1}:\rho(x_{4})\rightarrow\rho(x_{i})$, $\alpha_{ij}:\rho(x_{i})\rightarrow\rho(x_{j})$, and $\beta_{j4}:\rho(x_{j})\rightarrow\rho(x_{4})$.
  5. (5)
    By (3), $\beta_{i4}$ and $\beta_{j4}$ have the same $(n_{0}-1)$-th truncations as $\alpha_{i4}$ and $\alpha_{j4}$, respectively. So, the $(n_{0}-1)$-th truncation of $f_{ij}=\beta_{j4}\circ\alpha_{ij}\circ\beta_{i4}^{-1}$ is the same as that of
    $\alpha_{j4}\circ\alpha_{ij}\circ\alpha_{i4}^{-1}=\alpha_{j4}\circ\alpha_{ij}\circ(\alpha_{j4}\circ\alpha_{ij})^{-1}$
    $=\alpha_{j4}\circ\alpha_{ij}\circ\alpha_{ij}^{-1}\alpha_{j4}^{-1}=\mathrm{id}:y\rightarrow y.$
  6. (6)
    Clearly, $f_{ij}$ is interdefinable with $\alpha_{ij}$ over $\overline{\beta}$. Since $\overline{\beta}\overline{t}$ and $\overline{\alpha}\overline{s}$ realize the same $\mathcal{K}$-type over $x_{1}x_{2}x_{3}x_{4}$ (and by definition of codes), it follows that $\overline{\beta}$ is interalgebraic with $\overline{t}x_{1}x_{2}x_{3}x_{4}=A$. Thus $f_{ij}$ is interalgebraic with $\alpha_{ij}$ over $A$. Finally, again by definition of codes, $\alpha_{ij}$ is interalgebraic with $s_{ij}$ over $A$. Thus, $f_{ij}$ is interalgebraic with $s_{ij}$ over $A$.
  7. (7)
    We have
    $f_{23}\circ f_{12}=(\beta_{34}\circ\alpha_{23}\circ\beta_{24}^{-1})\circ(\beta_{24}\circ\alpha_{12}\circ\beta_{14}^{-1})$
    $=\beta_{34}\circ(\alpha_{23}\circ\alpha_{12})\circ\beta_{14}^{-1}=\beta_{34}\circ\alpha_{13}\circ\beta_{14}^{-1}=f_{13}.$
  8. (8)
    By construction (see Lemma 4.13(4)), $A\overline{s}$ is coherent, and $A$ is $\mathcal{M}$-independent from $\overline{s}$ over $x_{1}x_{2}x_{3}x_{4}$. So since $\overline{s}$ is a rank 1 group configuration in $\mathcal{M}$ over $x_{1}x_{2}x_{3}x_{4}$, the same holds over $A$.
Now using Lemma 7.41(8), we can apply Fact 7.3 to $\overline{s}$ and $A$. We obtain:
Lemma 7.42.
There are
  • A parameter set $B\supset A$,
  • An $\mathcal{M}(B)$-interpretable group $G$ which is strongly minimal as an interpretable set in $\mathcal{M}$, and
  • For each $i,j\in\{1,2,3\}$ with $i<j$, an $\mathcal{M}$-generic element $g_{ij}\in G$ over $B$,
such that:
  1. (1)
    $g_{12}$ and $g_{23}$ are $\mathcal{M}$-independent over $B$,
  2. (2)
    $g_{13}=g_{23}\cdot g_{12}$, and
  3. (3)
    For each $i,j\in\{1,2,3\}$ with $i<j$, $g_{ij}$ and $s_{ij}$ are $\mathcal{M}$-interalgebraic over $B$.
We can thus add to our above list of conclusions:
Lemma 7.43.
Each $g_{ij}$ is interalgebraic with $f_{ij}$ over $B$.
Proof.
Note that $s_{ij}$ is interalgebraic with each of them over $B$ (by Lemma 7.42(3) and Lemma 7.41(6)). ∎
At this point, Lemmas 7.41 and 7.43 cover all of the requirements of Proposition 7.33, with one exception: we need $g_{12}$ and $g_{23}$ to be independent generics in $G$ over $B$, and we only know they are $\mathcal{M}$-independent $\mathcal{M}$-generics. To remedy this, note that all of our initial assumptions on $B$ and the $g_{ij}$ (i.e. those in Lemma 7.42) are expressed solely in terms of the structure $\mathcal{M}$. So Lemmas 7.42 and 7.43 will still hold after replacing $Bg_{12}g_{23}g_{13}$ with any tuple realizing the same $\mathcal{M}$-type over $A$. Now by Lemma 4.13(6), there is one such tuple which is coherent over $A$ (in case $B$ is infinite, we interpret coherence as saying that all finite subsets are coherent; in this case, enumerate $B$ and apply Lemma 4.13(6) inductively). Then, by coherence, we can transfer $\mathcal{M}$-genericity to full genericity (i.e. Lemma 4.13(1)). In particular, $g_{12}$ and $g_{23}$ are now independent generics in $G$ over $B$, as desired. ∎
7.6. Detecting Tangency and Finding Codes
We now turn to the task of coding $n_{0}$-slopes. The following is an analog of the notion of ‘detecting generic non-transversalities’ from [Cas23]:
Definition 7.44.
$\mathcal{M}$ detects tangency if whenever $(x,y)\in M^{2}$ is generic, and $\alpha$ is a non-algebraic coherent $n$-slope at $(x,y)$ for some $n$, then any two coherent representatives of $\alpha$ are $\mathcal{M}$-dependent over $(x,y)$.
We will show that if $\mathcal{M}$ detects tangency, every coherent $n_{0}$-slope has a code, and thus by Proposition 7.33, $\mathcal{M}$-interprets a strongly minimal group. In [Cas23], one has a clear construction of codes: indeed, in ACF0, a rank 2 family of plane curves realizes every non-algebraic coherent $n_{0}$-slope finitely many times, and one can code a slope using its finite set of occurrences in any such family. However, in the current setting, there is no reason that a non-algebraic coherent slope must occur in a rank 2 family. So we might have an infinite set of occurrences in a larger family, any two of which are $\mathcal{M}$-dependent by assumption, and we want to ‘code’ this infinite set. The proper tool for doing such a thing is canonical bases. As it turns out, we will code each slope using the canonical base of the dependence between two independent occurrences of that slope in a family.
Proposition 7.45.
If $\mathcal{M}$ detects tangency, then every coherent $n_{0}$-slope has a code.
Proof.
The idea is to use the detection of tangency to $\mathcal{M}$-definably organize the curves in a family by slope, and then take the ‘equivalence classes’ (i.e. canonical bases) to be the codes. For this to work, we need to know that the $\mathcal{M}$-dependence between coherent representatives in Definition 7.44 is not ‘coarser’ than the tangency relation, so that the canonical bases we take are genuinely capturing curves up to tangency. This will work because we are only considering one-dimensional families of slopes (i.e. by the minimality of $n_{0}$). The precise statements we need are given in Lemmas 7.46 and 7.47 below:
Lemma 7.46.
Let $\alpha$ be a coherent $n_{0}$-slope at $(x,y)$. Let $c_{1}$ and $c_{2}$ be coherent representatives of $\alpha$, and assume that $c_{1}$ and $c_{2}$ are independent over $\alpha$. Then $c_{2}$ is coherent over $c_{1}xy$.
Proof.
If $\alpha\in\operatorname{acl}(xy)$, then $c_{1}$ and $c_{2}$ are independent over $xy$. By assumption, each $c_{i}$ is coherent over $xy$, so it follows that $c_{1}c_{2}$ is coherent over $xy$, making the lemma obvious. Thus, we assume $\alpha$ is not algebraic.
By Lemma 4.11, $\dim(c_{2}/c_{1}xy)\leq\operatorname{rk}(c_{2}/c_{1}xy)$. We prove the reverse inequality.
Since $\mathcal{M}$ detects tangency, $\operatorname{rk}(c_{2}/c_{1}xy)\leq\operatorname{rk}(c_{2}/xy)-1$. By assumption $c_{2}xy$ is coherent, so $\operatorname{rk}(c_{2}/xy)-1=\dim(c_{2}/xy)-1$. By Lemma 7.25, $\dim(c_{2}/xy)-1=\dim(c_{2}/\alpha)$. Since the $c_{i}$ are independent over $\alpha$, $\dim(c_{2}/\alpha)=\dim(c_{2}/c_{1}\alpha)$. By Lemma 7.20, $\dim(c_{2}/c_{1}\alpha)=\dim(c_{2}/c_{1}xy)$.
Putting together everything in the last paragraph gives $\ rk(c_{2}/c_{1}xy)\leq\dim(c_{2}/c_{1}xy)$, which proves the lemma. ∎
Lemma 7.47.
Let $\alpha$ be a coherent $n_{0}$-slope at $(x,y)$. Let $c$, $c_{1}$, and $c_{2}$ be coherent representatives of $\alpha$, and assume $c$ is independent from $c_{1}c_{2}$ over $\alpha$. Then $\operatorname{stp}_{\mathcal{M}}(c/c_{1}xy)$ and $\operatorname{stp}_{\mathcal{M}}(c/c_{2}xy)$ are parallel (i.e. have a common non-forking extension).
Proof.
By transitivity, we have (1) $c$ is independent from $c_{1}$ over $\alpha$, and (2) $c$ is independent from $c_{2}$ over $c_{1}\alpha$ – equivalently, by Lemma 7.20, over $c_{1}xy$. By (1) and Lemma 7.46, $c$ is coherent over $c_{1}xy$. So, by (2) and Lemma 4.11(4), $c$ is $\mathcal{M}$-independent from $c_{2}$ over $c_{1}xy$. Thus, $\operatorname{stp}_{\mathcal{M}}(c/c_{1}c_{2}xy)$ is a non-forking extension of $\operatorname{stp}_{\mathcal{M}}(c/c_{1}xy)$. By a symmetric argument, $\operatorname{stp}_{\mathcal{M}}(c/c_{1}c_{2}xy)$ is also a non-forking extension of $\operatorname{stp}_{\mathcal{M}}(c/c_{2}xy)$. ∎
We now prove the proposition. Let $\alpha$ be a coherent $n_{0}$-slope at $(x,y)$. Let $c_{1}$ be a coherent representative of $\alpha$. Let $c_{2}$ be an independent realization of $\operatorname{tp}_{\mathcal{K}}(c_{1}/\alpha)$ over $c_{1}$. Let $p=\operatorname{stp}_{\mathcal{M}}(c_{2}/c_{1}xy)$, and let $b=\operatorname{Cb}(p)$. We show that $s=bxy$ is a code of $\alpha$, by verifying (1), (2), and (3) in Definition 7.26:
  1. (1)
    It is obvious that $(x,y)\in\operatorname{acl}_{\mathcal{M}}(s)$, since $(x,y)$ is included in $s$.
  2. (2)
    Let $c$ be a coherent representative of $\alpha$. We show that $s\in\operatorname{acl}_{\mathcal{M}}(cxy)$. It is enough to show that $b\in\operatorname{acl}_{\mathcal{M}}(cxy)$.
    By replacing $c_{2}$ with an independent realization if necessary, we may assume that $c_{2}$ is independent from $c$ over $c_{1}xy$ (note that this does not change $b$). In particular, this easily implies that $c_{2}$ is independent from $c_{1}c$ over $\alpha$. Then by Lemma 7.47, $q=\operatorname{stp}_{\mathcal{M}}(c_{2}/cxy)$ is parallel to $p$, so has canonical base $b$. Since $q$ is a type over $cxy$, this implies that $b\in\operatorname{acl}_{\mathcal{M}}(cxy)$.
  3. (3)
    We now have to show that $s$ and $\alpha$ are interalgebraic. First, by (2) and Lemma 7.20, we have
    $s\in\operatorname{acl}(c_{1}\alpha)\cap\operatorname{acl}(c_{2}\alpha).$
    Since the $c_{i}$ are independent over $\alpha$, this implies $s\in\operatorname{acl}(\alpha)$.
    Next, we show that $s$ is coherent. Indeed, since $p$ is a type over $c_{1}xy$, we have $b\in\operatorname{acl}_{\mathcal{M}}(c_{1}xy)$. But $c_{1}xy$ is coherent by assumption, thus so is $c_{1}bxy$, and thus so is $s$.
    Now toward a contradiction, assume that $\alpha\notin\operatorname{acl}(s)=\operatorname{acl}(bxy)$. By Lemma 7.25, $\dim(\alpha/xy)=1$. So if $s\in\operatorname{acl}(\alpha)$ but $\alpha\notin\operatorname{acl}(s)$, then $b\in\operatorname{acl}(xy)$. Since $s$ is coherent, this implies $b\in\operatorname{acl}_{\mathcal{M}}(xy)$. Thus $p$ does not fork over $xy$, and thus $c_{1}$ and $c_{2}$ are $\mathcal{M}$-independent over $xy$. This contradicts that $\mathcal{M}$ detects tangency.
7.7. Tangent and Multiple Intersections
We now give conditions under which $\mathcal{M}$ detects tangency. The idea is to require that an ‘unusually high slope agreement’ between plane curves is always a ‘topological multiple intersection.’ This is analogous to similar facts regarding intersection multiplicity that were utilized in past trichotomy papers (e.g. Bezout’s theorem in [Rab93] and the argument principle/Rouche’s theorem in [MP90]).
In characteristic $p$ environments, it is possible for any two curves in a family to be first-order tangent at all of their intersection points. However, under mild assumptions, there is always a larger $n$ so that no two independent generic curves are $n$th order tangent at a point. For this reason, the notion of ‘unusually high’ slope agreement is slightly complicated. The informal idea is that for some $n$, the two curves should have the same $n$-slope even though most nearby pairs of curves do not. To make this precise, we will use the machinery of weakly generic intersections developed in section 6. We will specifically work with families of correspondences in $K$:
Definition 7.48.
A $\mathcal{K}(A)$-definable family of correspondences in $K$ is a $\mathcal{K}(A)$-definable family $\mathcal{X}=\{X_{t}:t\in T\}$ of subsets of $K^{2}$, such that each $X_{t}\subset K^{2}$ has dimension 1 and projects finite-to-one to both copies of $K$.
In addition to the terminology on weakly generic intersections developed in section 6, we need the following notions involving slopes:
Definition 7.49.
Let $\mathcal{X}=\{X_{t}:t\in T\}$ and $\mathcal{Y}=\{Y_{u}:u\in U\}$ be $\mathcal{K}(A)$-definable families of correspondences in $K$.
  1. (1)
    If $n\in\mathbb{Z}^{+}$ and $V\subset K^{2}\times T\times U$ is any open set, we say that $\mathcal{X}$ and $\mathcal{Y}$ have $n$-branching in $V$ over $A$ if for all strongly generic $(\mathcal{X},\mathcal{Y})$-intersections $(x,y,t,u)$ over $A$, if $(x,y,t,u)\in V$ then $\alpha_{xy}^{n}(X_{t})\neq\alpha_{xy}^{n}(Y_{u})$.
  2. (2)
    A generic $(\mathcal{X},\mathcal{Y})$-tangency over $A$ is a tuple $(x,y,t,u,n)$ such that $(x,y,t,u)$ is a weakly generic $(\mathcal{X},\mathcal{Y})$-intersection, $n\in\mathbb{Z}^{+}$, $\alpha_{xy}^{n}(X_{t})=\alpha_{xy}^{n}(Y_{u})$, and $X$ and $Y$ have $n$-branching in a neighborhood of $(x,y,t,u)$.
Definition 7.50.
$(\mathcal{K},\tau,\{\mathcal{IA}_{n}\},\{T_{mn}\})$ satisfies TIMI (‘tangent intersections are multiple intersections’) if the following holds: suppose $\mathcal{X}=\{X_{t}:t\in T\}$ and $\mathcal{Y}=\{Y_{u}:u\in U\}$ are $\mathcal{K}(A)$-definable families of correspondences in $K$. If $(x,y,t,u,n)$ is a generic $(\mathcal{X},\mathcal{Y})$-tangency over $A$, then $(x,y,t,u)$ is a generic multiple $(\mathcal{X},\mathcal{Y})$-intersection over $A$.
Remark 7.51.
Recall that at this moment we are working with fixed definable slopes on $(\mathcal{K},\tau)$. In the future, after dropping this assumption, we will use the phrase ‘$(\mathcal{K},\tau)$ has definable slopes satisfying TIMI’ to mean that $(\mathcal{K},\tau)$ has definable slopes $\{\mathcal{IA}_{n}\}$ and $\{T_{mn}\}$ so that $(\mathcal{K},\tau,\{\mathcal{IA}_{n}\},\mathcal{T}_{mn})$ satisfies TIMI.
Recall the notion of detection of multiple intersections, Definition 6.9. In those terms, our main result is:
Theorem 7.52.
Assume that $(\mathcal{K},\tau,\{\mathcal{IA}n\},\{T_{mn}\})$ satisfies TIMI, and $\mathcal{M}$ detects multiple intersections. Then $\mathcal{M}$ detects tangency. In particular, $\mathcal{M}$ interprets a strongly minimal group.
Proof.
Let $\alpha_{0}$ be a non-algebraic coherent $n$-slope at $(x_{0},y_{0})$, and let $(x_{0},y_{0},C_{1},c_{1},\alpha_{0})$ and $(x_{0},y_{0},C_{1},c_{1},\alpha_{0})$ be coherent $n$-slope configurations. We want to show that $c_{1}$ and $c_{2}$ are $\mathcal{M}$-dependent over $(x_{0},y_{0})$. We do this by realizing the $C_{i}$ as generic members of $\mathcal{M}$-definable families of plane curves, $Z_{t_{0}}\in\mathcal{Z}$ and $W_{u_{0}}\in\mathcal{W}$; we then apply TIMI to these families (or rather, their images through our fixed map $\rho:M\rightarrow K$, which we call $\mathcal{X}$ and $\mathcal{Y}$, respectively). The idea is that the non-algebraicity of $\alpha_{0}$ forces each of $\mathcal{X}$ and $\mathcal{Y}$ to realize infinitely many slopes at $(x_{0},y_{0})$, which forces $(x_{0},y_{0},t_{0},u_{0},n)$ to be a generic $(\mathcal{X},\mathcal{Y})$ tangency. By TIMI, we conclude that $(x_{0},y_{0},t_{0},u_{0})$ is a generic multiple $(\mathcal{X},\mathcal{Y})$-intersection; then since $\mathcal{M}$ detects multiple intersections, we get the desired dependence.
Let us proceed. Let $d_{i}=\operatorname{rk}(c_{i})$ for $i=1,2$. By Lemma 7.21, $(x_{0},y_{0})$ is generic in $M^{2}$, and each $\dim(c_{i}/x_{0}y_{0})=d_{i}-1$. Thus $d_{i}\geq 1$. Now by Lemma 6.8, we can realize each $C_{i}$ on a generic curve in a standard family (see Definition 2.4 and Lemma 6.8). That is, we can find standard families of non-trivial plane curves $\mathcal{Z}=\{Z_{t}:t\in T\}$ and $\mathcal{W}=\{W_{u}:u\in U\}$ over $\emptyset$, and generic $t_{0}\in T$ and $u_{0}\in U$, so that $C_{1}$ is almost contained in $Z_{t_{0}}$ and $C_{2}$ is almost contained in $W_{u_{0}}$. For each $t\in T$, let $X_{t}=\rho(Z_{t})$. Similarly, for each $u\in U$, let $Y_{u}=\rho(W_{u})$. Then $\mathcal{X}=\{X_{t}:t\in T\}$ and $\mathcal{Y}=\{Y_{u}:u\in U\}$ are $\mathcal{K}(\emptyset)$-definable families of correspondences in $K$. Note that $\dim(T)=d_{1}$ and $\dim(U)=d_{2}$.
We want to apply TIMI to $\mathcal{X}$ and $\mathcal{Y}$, at the tuple $(\rho(x_{0}),\rho(y_{0}),t_{0},u_{0},n)$. First, note that $(x_{0},y_{0},t_{0},u_{0})$ is a weakly generic $(\mathcal{Z},\mathcal{W})$-intersection over $\emptyset$. It follows easily that $(\rho(x_{0}),\rho(y_{0}),t_{0},u_{0})$ is a weakly generic $(\mathcal{X},\mathcal{Y})$-intersection over $\emptyset$.
Next, by the strong frontier inequality, $Z_{t_{0}}$ and $C_{1}$ agree in a neighborhood of $(x_{0},y_{0})$, as do $W_{u_{0}}$ and $C_{2}$. So $\alpha_{x_{0}y_{0}}^{n}(Z_{t_{0}})=\alpha_{x_{0}y_{0}}^{n}(W_{u_{0}})=\alpha_{0}$. Finally, we check:
Claim 7.53.
$\mathcal{X}$ and $\mathcal{Y}$ have $n$-branching in a neighborhood of $(\rho(x_{0}),\rho(y_{0}),t_{0},u_{0})$ over $\emptyset$.
Proof.
By uniform definability of slopes, $\alpha_{0}\in\operatorname{acl}(t_{0}x_{0}y_{0})$. Since $\alpha_{0}$ is not algebraic, this implies $\dim(t_{0}/\alpha_{0})<\dim(t_{0}/x_{0}y_{0})=\dim(c_{1}/x_{0}y_{0})=d_{1}-1$. So $\dim(t_{0}/x_{0}y_{0})=d_{1}-1$ and $\dim(t_{0}/\alpha_{0})\leq d_{1}-2$.
By uniform definability of slopes, and the definability of dimension, we can find a formula $\phi(x,y,t)\in\operatorname{tp}_{\mathcal{K}}(x_{0},y_{0},t_{0})$ such that whenever $(x,t,y)\in Z$ is generic and $\phi(x,y,t)$ holds, then for $\alpha=\alpha_{xy}^{n}(Z_{t})$, we have $\dim(t/xy)\leq d_{1}-1$ and $\dim(t/\alpha)\leq d_{1}-2$. By the genericity of $(x_{0},y_{0},t_{0})$, $\phi$ holds in a neighborhood of $(x_{0},y_{0},t_{0})$, which we may assume has the form $V_{1}\times V_{2}$, where $V_{1}\subset M^{2}$ and $V_{2}\subset T$. Shrinking if necessary, we may assume that $\rho$ is a local homeomorphism on $V_{1}$, so $\rho(V_{1})\subset K^{2}$ is open. Now let $V=\rho(V_{1})\times V_{2}\times U\subset K^{2}\times T\times U$.
To prove the claim, we show that $\mathcal{X}$ and $\mathcal{Y}$ have $n$-branching in $V$ over $\emptyset$. To do this, let $(\rho(x),\rho(y),t,u)\in V$ be a strongly generic $(\mathcal{X},\mathcal{Y})$-intersection over $\emptyset$, where $(x,y)\in V_{1}$. Let $\alpha=\alpha_{\rho(x)\rho(y)}^{n}(X_{t})=\alpha_{xy}^{n}(Z_{t})$.
Since $\rho$ is finite-to-one, it follows that $(x,y,t)\in Z$ is generic. Thus $\dim(xyt)=\dim(Z)=d_{1}+1$. But by the choice of $V_{1}\times V_{2}$, $\dim(t/xy)\leq d_{1}-1$, thus $\dim(xy)=2$.
Now by uniform definability of slopes,
$\dim(txy\alpha)=\dim(txy)=d_{1}+1.$
Since $\dim(xy)=2$, this implies $\dim(t\alpha/xy)=d_{1}-1$. But by the choice of $V_{1}\times V_{2}$ again, $\dim(t/\alpha xy)\leq d_{1}-2$. Thus $\dim(\alpha/xy)\geq 1$. In particular, $\alpha$ is not algebraic.
Now let $\beta=\alpha_{xy}^{n}(W_{u})$. So $\alpha\in\operatorname{acl}(txy)$ and $\beta\in\operatorname{acl}(uxy)$. Since $t$ and $u$ are independent over $\rho(x)\rho(y)$, they are clearly also independent over $xy$. Thus, $\alpha$ and $\beta$ are independent over $xy$. In particular, $\alpha\notin\operatorname{acl}(\beta)$, and thus $\alpha\neq\beta$. This proves the claim. ∎
Now, by the claim, we have verified that $(\rho(x_{0}),\rho(y_{0}),t_{0},u_{0},n)$ is a generic $(\mathcal{X},\mathcal{Y})$-tangency over $\emptyset$. By TIMI, we conclude that $(\rho(x_{0}),\rho(y_{0}),t_{0},u_{0})$ is a generic multiple $(\mathcal{X},\mathcal{Y})$-intersection over $\emptyset$. Since $\rho$ is a local homeomorphism near $(x_{0},y_{0})$, it follows that $(x_{0},y_{0},t_{0},u_{0})$ is a generic multiple $(\mathcal{Z},\mathcal{Y})$-intersection over $\emptyset$. But then since $\mathcal{M}$ detects multiple intersections, we conclude that $t_{0}$ and $u_{0}$ are $\mathcal{M}$-dependent over $\emptyset$. By $\mathcal{M}$-interalgebraicity, so are $c_{1}$ and $c_{2}$.
Finally, by coherence, we have $\operatorname{rk}(c_{1}/x_{0}y_{0})=\dim(c_{1}/x_{0}y_{0})=d_{1}-1$, and similarly $\operatorname{rk}(c_{2}/x_{0}y_{0})=d_{2}-1$. So if $c_{1}$ and $c_{2}$ are $\mathcal{M}$-independent over $x_{0}y_{0}$, then $\operatorname{rk}(c_{1}c_{2}/x_{0}y_{0})=d_{1}+d_{2}-2$. Since $(x_{0},y_{0})$ is generic in $M^{2}$, this implies $\operatorname{rk}(c_{1}c_{2}x_{0}y_{0})=d_{1}+d_{2}$. Since $c_{1}$ and $c_{2}$ are $\mathcal{M}$-dependent over $\emptyset$, $\operatorname{rk}(c_{1}c_{2})<d_{1}+d_{2}$, so $\operatorname{rk}(x_{0}y_{0}/c_{1}c_{2})>0$. Since $c_{1}$ and $c_{2}$ are canonical bases of strongly minimal sets, this implies $c_{1}=c_{2}$. But then by the assumed $\mathcal{M}$-independence of the $c_{i}$ over $x_{0}y_{0}$,
$10=\operatorname{rk}(c_{1}/c_{2}x_{0}y_{0})=\operatorname{rk}(c_{1}/x_{0}y_{0})=d_{1}-$ 1 .
So $d_{1}=1$. This contradicts that $\dim(t_{0}/\alpha_{0}x_{0}y_{0})\leq d_{1}-2$. We conclude that the $c_{i}$ are independent over $x_{0}y_{0}$, proving that $\mathcal{M}$ detects tangency.
Finally, since $\mathcal{M}$ detects tangency, Proposition 7.45 implies that every coherent $n_{0}$-slope has a code. Then Proposition 7.33 implies that $\mathcal{M}$ interprets a strongly minimal group. ∎
8. Summing Up
We are now done with the abstract setting, and we drop all data that we have fixed up until now. The rest of the paper concerns concrete examples. Before moving on, let us summarize our work to this point:
Theorem 8.1.
Let $(\mathcal{K},\tau)$ be a Hausdorff geometric structure with enough open maps (in particular, it suffices to assume $(\mathcal{K},\tau)$ is either differentiable or has the open mapping property). Assume that $(\mathcal{K},\tau)$ has ramification purity, and definable slopes satisfying TIMI. Let $\mathcal{M}=(M,...)$ be a non-locally modular strongly minimal definable $\mathcal{K}$-relic. Then $\dim(M)=1$, and $\mathcal{M}$ interprets a strongly minimal group.
Proof.
Since non-local modularity is witnessed by a single rank 2 family of plane curves, we may assume the language of $\mathcal{M}$ is finite. If we then add countably many constants, we can assume the language of $\mathcal{M}$ is countable and $\operatorname{acl}_{\mathcal{M}}(\emptyset)$ is infinite.
Next, all of the assumptions on $(\mathcal{K},\tau)$ are invariant under adding a countable set of parameters. Since the language of $\mathcal{M}$ is countable, we may thus assume that every $\emptyset$-definable set in $\mathcal{M}$ is $\emptyset$-definable in $\mathcal{K}$. Thus, we have satisfied all requirements in Assumption 4.1, and so we can apply all the results of sections 4-7.
Now by assumption (or by Proposition 3.33 or Proposition 3.40, depending on which property is assumed), $(\mathcal{K},\tau)$ has enough open maps. So, by Theorem 4.10, $\mathcal{M}$ weakly detects closures. Then, by Theorem 6.10, $\mathcal{M}$ detects multiple intersections. So by Theorem 6.18, $\dim(M)=1$. Then, by Theorem 7.52, $\mathcal{M}$ interprets a strongly minimal group. ∎
Remark 8.2.
One also gets the various additional properties of the group in the statement of Proposition 7.33. We omit them here in order to not overcomplicate the statement. Note, in particular, that changing the language of $\mathcal{M}$ in the proof above does not affect the additional data in Proposition 7.33, because that data only involves the background structure $\mathcal{K}$.
9. Examples of Hausdorff Geometric Structures
In this section, we give various examples of Hausdorff geometric structures satisfying some or all of the additional properties we have studied. The most successful example is algebraically closed valued fields, which satisfy all properties (except differentiability in positive characteristic) we have defined.
9.1. Visceral structrues
In [DG22] Dolich and Goodrick introduce visceral structures as a common generalisation of o-minimality and P-minimality, that – when restricted to the dp-minimal setting – is very similar to the tame uniform structures of Simon and Walsberg ([SW19]) called SW-uniformities in [HHP22]. This formalism also covers the context of 1-h-minimal structures (as pointed out in [CHRV21, Example 2.2.2]). It follows immediately from the results of [DG22] that certain strengthenings of visceral structures considered in that paper are Hausdorff geometric structures.
By [DG22, Proposition 3.10] any visceral structure has uniform finiteness, so any visceral structure satisfying the exchange principle is geometric. A visceral structure is said to have definable finite choice (DFC) if any definable function with finite fibres has a definable section. In particular, any ordered visceral structure has DFC, and so does any visceral structure with definable Skolem functions. We claim that ($\aleph_{1}$-saturated) visceral structures with the exchange property and DFC are Hausdorff geometric structures (see Remark 9.1 below for a discussion of removing the DFC hypothesis).
First, such structures satisfy the strong frontier inequality by Remark 3.9 and [DG22, Corollary 3.35]. The Baire property, as explained in Remark 3.8, would follow from the generic local homeomorphism property and the fact that the topology in Hausdorff visceral structures has no isolated points. So it remains to show the generic local homeomorphism property. So let $Z\subset X\times Y$, all of the same dimension, with $Z\rightarrow X$ and $Z\rightarrow Y$ finite-to-one, and let $(x,y)\in Z$ be generic. By visceral cell decomposition ([DG22], Theorem 3.23, Definition 3.20), we can assume after shrinking that $X$ and $Y$ are open subsets of $K^{n}$, where $n=\dim(X)$. Moreover, by either directly applying DFC or by using the frontier inequality (see the proof of Claim 9.34), we can assume after shrinking again that $Z\rightarrow X$ and $Z\rightarrow Y$ are injective, and thus $Z$ defines a bijection $X\leftrightarrow Y$. The result then follows by generic continuity ([DG22, Theorem 3.19]) applied to each coordinate component of $X\leftrightarrow Y$ in each direction.
Remark 9.1.
In fact, it seems to follow from new results of Johnson ([Joh24b]) that one can remove the DFC assumption above and conclude that all $\aleph_{1}$-saturated visceral structures with exchange are Hausdorff geometric structures. For this, one first deduces the strong frontier inequality from Johnson’s new version ([Joh24b], Theorem 1.10). Now given $Z\subset X\times Y$ as above, one similarly uses the frontier inequality to assume $Z\rightarrow X$ and $Z\rightarrow Y$ are injective (again, as in Claim 9.34), and thus $Z$ defines a bijection $X\leftrightarrow Y$. One then applies the generic continuity clause of Johnson’s cell decomposition ([Joh24b], Theorem 1.14(2), which is enough by the frontier inequality) to each coordinate component of $X\rightarrow Y$ and $Y\rightarrow X$ as above to get the desired homeomorphism.
Hensel minimal structures, introduced in [CHRK22] (for residue characteristic 00 and in [CHRV21] for positive residue characteristic) cover all pure non-trivially valued henselian fields (of characteristic 00) as well as many interesting expansions (covering the $V$-minimal algebraically closed valued fields of [HK06] and the $T$-convex expansions of power-bounded o-minimal fields, [vdDL95]). One sub-class of Hensel minimal structures of special interest (covering all the above examples) is the class of 1-h-minimal fields. As already mentioned, 1-h-minimal structures are visceral, and it is not hard to verify that they are geometric structures (e.g., [AH24, Proposition 2.11]). While 1-h-minimal fields do not necessarily have DFC, every 1-h-minimal field has a 1-h-minimal expansion which does ([CHRV21, Proposition 3.2.3]). Thus, we have:
Corollary 9.2.
Every $\aleph_{1}$-saturated 1-h-minimal field is a Hausdorff geometric structure.
Proof.
By what we have just said, any 1-h-minimal field $\mathcal{K}$ has a DFC expansion $\mathcal{K}^{\prime}$ that is a Hausdorff Geometric Structure.
In both $\mathcal{K}$ and $\mathcal{K}^{\prime}$, the topology is the valuation topology, and it is definable. Since the dimension in both structures is determined by the topology (see, e.g., [AH24, Proposition 2.11]), it follows that for a $\mathcal{K}$-definable set $X$ we have $\dim_{\mathcal{K}}(X)=\dim_{\mathcal{K}^{\prime}}(X)$. Now consider $X$ a $\mathcal{K}$-definable set. Then $\mathrm{Fr}(X)$ is also $\mathcal{K}$-definable, and by the frontier inequality in $\mathcal{K}^{\prime}$ we have that
$\dim_{\mathcal{K}}\mathrm{Fr}(X)=\dim_{\mathcal{K}^{\prime}}\mathrm{Fr}(X)<\dim_{\mathcal{K}^{\prime}}(X)=\dim_{\mathcal{K}}(X),$
implying the frontier inequality in $\mathcal{K}$.
So we only have to verify the Generic Local Homeomorphism Property. If $Z\subseteq X\times Y$ are $\mathcal{K}$-definable of the same dimension, projecting finite-to-one one both components, then for every $\mathcal{K}^{\prime}$-generic $(x,y)\in Z$ we know that $Z$ is, in a neighbourhood of $(x,y)$ the graph of a homeomorphism. As this is a $\mathcal{K}$-definable property of $(x,y)$ (and since dimension in $\mathcal{K}$ and in $\mathcal{K}^{\prime}$ coincide), it must also hold of every $\mathcal{K}$-generic $(x,y)\in Z$. ∎
Remark 9.3.
As stated before, if one uses the new results of Johnson on visceral structures [Joh24b], one does not need DFC, so the above corollary is automatic from the discussion preceding it.
Both o-minimal and 1-h-minimal expansions of fields have a well-developed (and rather similar) basic differential geometry, turning them into differential Hausdorff geometric structures. Let us give the details in the 1-h-minimal case. The situation in the o-minimal setting is analogous and better known.
First, we have to define a notion of smoothness meeting Definition 3.27. We declare the smooth locus of a definable set $X$ to be the set of points $x\in X$ where $X$ is, locally near $x$, a $\dim(X)$-dimensional weak strictly differentiable manifold (see [AH24, Definition 5.3]). By visceral cell decomposition, every definable set is generically locally a topological manifold, and by Proposition 3.12 loc. cit. it is, in fact, generically, a strictly differentiable manifold. This gives all the properties of Definition 3.27 except maybe (5), the preservation of smoothness under preimages of “nice enough” projections. But recall (see the remark following Definition 3.27) that this clause is a generalization of, and would follow from, the ‘submersion theorem’ of differential geometry. Moreover, there is a submersion theorem for 1-h-minimal field ([AH24, Proposition 4.10]). So we simply apply that proposition exactly as in the earlier remark.
So now we have to show that the above notion of smoothness admits a differential structure, as in Definition 3.34. Naturally, in both the 1-h-minimal and the o-minimal settings, the differential structure is given by the tangent space. As before, we focus on the 1-h-minimal setting, as the o-minimal case is similar and more familiar.
We use the usual notion of a tangent space associated with the notion of (weak) strictly differentiable manifolds ([AH24, Definition 5.4]). All properties (1)-(7) of the definition are automatic, and we don’t dwell on them. Clause (8) of the definition, asserting that a definable morphism inducing an isomorphism on the level of tangent spaces is open (locally, near the point in question) is an immediate consequence of the Inverse Mapping Theorem ([AH24, Proposition 4.4]). So it only remains to check that if $f:(X,x)\to(Y,y)$ is such that $x\in X$ and $y\in Y$ are generic, then the induced map between the tangent spaces is surjective. This, in turn, is an immediate result of the corresponding version of Sard’s Theorem [AH24, Proposition 5.21], asserting that the set of singular values of $f$ is non-generic.
So we conclude:
Theorem 9.4.
Every $\aleph_{1}$-saturated o-minimal expansion of a real closed field, as well as every $\aleph_{1}$-saturated 1-h-minimal valued field, is a differential Hausdorff geometric structure.
9.2. éz Expansions of Fields
We turn now toward ACVF in all characteristics. We will work with a general setting of topological fields. In fact, the ensuing results could serve as an alternate proof of Theorem 9.4 for RCVF and for pure characteristic zero Henselian fields.
Our abstract setting is inspired by éz fields. We briefly recall this notion. First, recall that a field $K$ is large if, whenever $V$ is a smooth variety over $K$ and $V(K)\neq\emptyset$, then $V(K)$ is infinite (equivalently, $V(K)$ is Zariski dense in some irreducible component). In [JTWY23], the authors place a canonical non-discrete topology on the set $V(K)$ for every (irreducible) variety $V$ over a large field $K$. The topology is called the étale open topology, and is defined as the weakest topology where étale maps are open (so it is generated by the images $f(W(K))$ of all étale morphisms of $K$-varieties $f:W\rightarrow V$). It is shown in [JTWY23, Section 6] and  [JWY23, DWY22] that the étale open topology often coincides with existing topologies of interest when the field is close to being henselian (particularly, the order topology in real closed fields, and the valuation topology in characteristic zero Henselian fields).
One says that the étale open topology on $K$ is induced by a field topology if there is a field topology $\tau$ on $K$ such that, for every affine variety $V$ over $K$, the étale open topology agrees on $V(K)$ with the topology induced on $V(K)$ by $\tau$ (via the product and subspace topologies). In general, the étale open topology is induced by a field topology if it respects products (i.e. the étale open topology on $K^{n}$ is the product topology induced by the étale open topology on $K$). This is not always the case (the main counterexample being pseudofinite fields, or more generally PAC fields). See [JTWY23, Proposition 4.9 and Section 8] or [DWY22] for more details.
The follow-up work [WY23] then defined éz fields as large fields $K$ such that every definable set $X\subset K^{n}$ (in the pure field language) is a union of finitely many definable étale-open subsets of Zariski closed sets. The idea is that éz fields have ‘local quantifier elimination’ – where the word ’local’ is in the sense of the étale open topology. Éz fields include algebraically closed fields, real closed fields, pseudofinite fields, and characteristic zero Henselian fields (e.g. $p$-adically closed fields). It is shown in [WY23] that many tameness properties of definable sets in these examples (e.g. generic continuity of functions) can be adapted to the general éz setting.
It is well-known that algebraically closed valued fields and real closed valued fields satisfy a similar local quantifier elimination: every definable set is a union of finitely many valuation-open subsets of Zariski closed sets. In fact, this description of definable sets holds in all Henselian valued fields of characteristic 0 [vdD89]. Our goal is to introduce a single notion capturing such a structural decomposition and incorporating éz fields. Since we have used the product topology freely throughout the paper, we will restrict to those éz fields whose étale open topology is induced by a field topology.
Before making the general definition, we establish some conventions:
Notation 9.5.
[…]
  1. (1)
    By a variety over $K$, we mean a reduced, separate scheme of finite type over $K$. Varieties are denoted with letters such as $V$, $W$, etc. Note that varieties are not definable objects in a first-order sense.
  2. (2)
    If $V$ is a quasi-projective variety over $K$, we use $V(K)$ to denote the $\mathcal{K}$-interpretable set of $K$-rational points of $V$.
  3. (3)
    If $X\subset K^{n}$ is definable, the Zariski closure of $X$ is the smallest affine $K$-variety $V$ such that $X\subset V(K)$.
  4. (4)
    On the other hand, we still use the term Zariski closed for subsets of $K^{n}$: $X\subset K^{n}$ is Zariski closed if $X=V(K)$ for some affine variety $V$ over $K$.
  5. (5)
    Similarly, if $V$ is a variety over $K$, and $X\subset V(K)$, we say that $X$ is Zariski dense in $V$ if for every proper closed subvariety $W\subset V$, there is some $x\in X\cap V(K)-W(K)$.
  6. (6)
    If $V$ is a quasi-affine variety over $K$, and $A\subset K$, we distinguish between $V$ being defined over $A$ (i.e. defined by polynomials with coefficients in $A$) and $V(K)$ being definable over $A$ (definable in the structure $\mathcal{K}$ with parameters $A$).
  7. (7)
    If $V$ is a quasi-affine variety over $K$, the notation $\dim(V)$ refers to the dimension of $V$ as an algebraic variety. The notation $\dim(V(K))$ will only be used if $\mathcal{K}$ is a geometric structure; and in this case, it refers to dimension in the sense of geometric structures.
We will use freely the following:
Lemma 9.6.
[…]
  1. (1)
    $Z(K)$ is definable over $A$ (regardless of whether the variety $Z$ is over $A$).
  2. (2)
    If $Z^{S}$ is the smooth locus of $Z$, then $Z^{S}(K)$ is definable over $A$.
Proof.
These statements are unaffected by passing to an elementary extension of $\mathcal{K}$. In particular, as opposed to topological statements about Hausdorff geometric structures, the current lemma only references definable objects.
So, let us assume that $\mathcal{K}$ is a ‘monster model’ – i.e. $\kappa$-saturated and $\kappa$-strongly homogenous for some large $\kappa$. In this case, one can detect the parameters in a definition using automorphism invariance. Thus, to prove the lemma, it suffices to note that $Z(K)$ and $Z^{S}(K)$ are invariant under field automorphisms fixing $X$ setwise. ∎
We also need a general notion of an invariant topology on a structure:
Definition 9.7.
[…]
  1. (1)
    Each $X\in\mathcal{B}$ is definable.
  2. (2)
    Suppose $\phi(x,y)$ is a formula, where $x$ is a single variable and $y$ is a tuple. Let $a$ and $b$ be tuples in the arity of $y$ with $\operatorname{tp}(a)=\operatorname{tp}(b)$. Let $X$ and $Y$ be the sets defined by $\phi(x,a)$ and $\phi(x,b)$, respectively. Then $X\in\mathcal{B}$ if and only if $Y\in\mathcal{B}$.
Note that, if $(K,+,\cdot,\tau)$ is a topological field, then $\tau$ induces a natural topology on $V(K)$ for every variety $V$ over $K$ (the affine case is given by the subspace topology from $K^{n}$, and the general case is obtained by gluing; see [Mum88, Page 57, Chapter I Section 10], for example). Using this, and inspired by the above discussion, we now define:
Definition 9.8.
[…]
  1. (1)
    $(K,+,\cdot,\tau)$ is a non-discrete topological field.
  2. (2)
    Every definable set $X\subset K^{n}$ is a finite union of definable $\tau$-open subsets of Zariski closed sets.
  3. (3)
    For each étale morphism $f:V\rightarrow W$ of varieties over $K$, the induced map $V(K)\rightarrow W(K)$ is $\tau$-open.
Remark 9.9.
Suppose $(\mathcal{K},\tau)$ is an éz topological field expansion. Note that since $\tau$ is invariant, every elementary extension of $\mathcal{K}$ is canonically a field with an invariant topology satisfying (1) and (3) of Definition 9.8. However, (2) is not first-order unless an additional uniformity condition is assumed. Thus, it is not clear that elementary extensions of $\mathcal{K}$ are also éz topological field expansions (though this holds in most natural examples).
Note that if $(K,+,\cdot)$ is an $\aleph_{1}$-saturated éz field whose étale open topology is induced by a field topology, then $K$ is (trivially) an éz topological field expansion when endowed with the étale open topology.
It also follows from well-known facts that $\aleph_{1}$-saturated models of ACVF and RCVF are éz topological field expansions. We do not elaborate on RCVF, because it is 1-h-minimal (thus covered by the previous subsection). On the other hand, let us sketch the argument for ACVF.
The main fact we need is:
Fact 9.10 ([MB12a]).
Let $(K,v)$ be an algebraically closed valued field. Let $f:V\rightarrow W$ be a universally open morphism of $K$-varieties. Then the induced map $V(K)\rightarrow W(K)$ is open in the valuation topology.
Lemma 9.11.
Let $\mathcal{K}=(K,v)$ be an $\aleph_{1}$-saturated model of ACVF. Then $(K,v)$ is an éz topological field expansion.
Proof.
Let $\tau$ be the valuation topology. As is well-known, $\tau$ is a non-discrete Hausdorff field topology. Moreover, as the collection of balls is a $\emptyset$-definable basis, $\tau$ is invariant.
So it remains to check (2) and (3) in Definition 9.8.$(\mathcal{K},\tau)$. (2) is a well-known consequence of quantifier elimination (see also [vdD89]). For (3), recall that étale morphisms are universally open ([Sta24, Lemmas 03WT and 02GO]), and thus apply Fact 9.10. ∎
Assumption 9.12.
For the rest of section 9, we fix an éz topological field expansion $(\mathcal{K},\tau)$, with underlying field $(K,+,\cdot)$.
Remark 9.13.
Note that, since $\tau$ is a field topology, it refines the Zariski topology: for every affine variety $V$ over $K$, the set $V(K)$ is $\tau$-closed. Similarly, every morphism of $K$-varieties induces a $\tau$-continuous map on $K$-points. We will use these facts throughout.
9.3. Basic Properties
Our general goal is to show that $(\mathcal{K},\tau)$ is a Hausdorff geometric structure. First, we note that many of the facts from [WY23] transfer (that is, there is no dependence on the particular language or topology). Thus, we now prove a series of basic lemmas about $(\mathcal{K},\tau)$.
We begin with:
Lemma 9.14.
Let $X\subset K^{n}$ be definable and Zariski dense in $\mathbb{A}^{n}$. Then $X$ has non-empty interior in the sense of $\tau$. In particular, every infinite definable subset of $K$ has non-empty interior.
Proof.
By Definition 9.8(2), we can write $X=\bigcup_{i=1}^{m}U_{i}$, where each $U_{i}$ is a definable open subset of $Z_{i}(K)$ for some closed subvariety $Z_{i}$ of $\mathbb{A}^{n}$. Since $X$ is Zariski dense in $\mathbb{A}^{n}$, there is some $i$ such that $U_{i}$ is Zariski dense in $\mathbb{A}^{n}$. It follows that $Z_{i}=\mathbb{A}^{n}$, so that $Z_{i}(K)=K^{n}$. But then $U_{i}$ is open in $K^{n}$ (and since it is Zariski dense, it is non-empty). ∎
The next two lemmas are consequences of Lemma 9.14:
Lemma 9.15.
  1. (1)
    $K$ is perfect.
  2. (2)
    If $\operatorname{char}(K)=p>0$ then the Frobenius map $x\mapsto x^{p}$ is a homeomorphism.
Proof.
  1. (1)
    Suppose $\operatorname{char}(K)=p>0$ and let $K^{p}$ be the set of $p$th powers. By Lemma 9.14, $K^{p}$ has non-empty interior. Let $U\subset K^{p}$ be a non-empty definable open set. Since $K^{p}$ is a subfield, we may translate and assume $0\in U$.
    Now let $a\in K$. We show $a\in K^{p}$. If $a=0$ this is clear. If $a\neq 0$, then $U\cap\frac{1}{a}U$ is a neighborhood of 0, and since $\tau$ is non-discrete, there is some $b\in U\cap\frac{1}{a}U$ with $b\neq 0$. By the choice of $U$, there are $x,y\in K$ with $x^{p}=b$ and $y^{p}=ab$. Then $(\frac{y}{x})^{p}=a$.
  2. (2)
    By (1), $x\mapsto x^{p}$ is a continuous bijection. We show it is open. Since $\tau$ is a field topology, and $x\mapsto x^{p}$ respects addition, it will suffice to show that if $X\subset K$ is open and contains 0, then $X^{p}$ contains a neighborhood of 0. So, fix such an $X$. We may assume $X$ is definable. Since subtraction is continuous, there is an open definable $Y$ containing 0 such that $Y-Y\subset X$. Since $\tau$ is not discrete, $Y$ is infinite, and thus so is $Y^{p}$. So by Lemma 9.14, $Y^{p}$ contains an infinite open set $Z$. Let $a\in Z$. Then $Z-a$ is a neighborhood of 0, and
    $Z-a\subset Z-Z\subset Y^{p}-Y^{p}\subset X^{p},$
    as desired.
Lemma 9.16.
$\mathcal{K}$ is geometric.
Proof.
It was shown in [JY23] that if an expansion of a field satisfies exchange, it also eliminates $\exists^{\infty}$. Thus, we only need to show that $\mathcal{K}$ satisfies exchange. But if not, then there is a definable $X\subset K^{2}$ so that (1) the left projection $X\rightarrow K$ is finite-to-one and (2) the right projection $X\rightarrow K$ has infinitely many infinite fibers. By (2), $X$ is Zariski dense in $\mathbb{A}^{2}$, so by Lemma 9.14, $X$ has non-empty interior. Thus $X$ contains a set $B_{1}\times B_{2}$, where each $B_{i}\subset K$ is non-empty and open. Since $\tau$ is non-discrete, each $B_{i}$ is infinite. This contradicts (1). ∎
Now that we know $\mathcal{K}$ is geometric, we will use the notation $\dim$ on definable sets. We next show:
Lemma 9.17.
Let $V$ be a smooth variety over $K$. Let $X$ be a non-empty open definable subset of $V(K)$. Then $\dim(X)=\dim(V)$. In particular, $K$ is a large field.
Proof.
It is easy to see that $\dim(X)\leq\dim(V)$. We show the reverse inequality. Let $a\in X$, and $n=\dim(V)$. Using that $a$ is a smooth point of $V$, one can show there is a rational map $f:V\rightarrow\mathbb{A}^{n}$ which is defined and étale at $a$. Let $U\subset V$ be the (necessarily relatively Zariski open) étale locus of $f$. Without loss of generality, we may assume $X\subset U(K)$. Now by Definition 9.8(3), the image $\pi(X)\subset K^{n}$ is non-empty and open, so contains a non-empty definable box $B_{1}\times...\times B_{n}$, where each $B_{i}\subset K$ is open. Since $\tau$ is non-discrete, each $B_{i}$ is infinite. Thus $\dim(B_{1}\times...\times B_{n})=n$. Finally, since $\pi$ is étale on $U$ it is finite-to-one, and thus preserves dimension. So
$\dim(X)=\dim(f(X))\geq\dim(B_{1}\times...\times B_{n})=n.$
The next three lemmas give the main technical tools we will need to show that $(\mathcal{K},\tau)$ is a Hausdorff geometric structure. We will use the following notion:
Definition 9.18.
Let $X\subset K^{n}$ be definable, with Zariski closure $Z$. Let $Z^{S}$ be the smooth locus of $Z$. Let $x\in Z(K)$. We call $x$ polished for $X$ if there is an open neighborhood $U$ of $x$ in $Z(K)$ such that:
  1. (1)
    $U\subset Z^{S}(K)$.
  2. (2)
    $U$ is either contained in $X$ or disjoint from $X$.
Lemma 9.19.
Let $X\subset K^{n}$ be definable, with Zariski closure $Z$. Then there is a closed subvariety $T\subset Z$ with $\dim(T)<\dim(Z)$, such that every $x\in Z(K)-T(K)$ is polished for $X$.
Proof.
By Definition 9.8(2), we can write $X=\bigcup_{i=1}^{m}U_{i}$ and $Z(K)=\bigcup_{j=1}^{k}V_{j}$, where each $U_{i}$ (resp. $V_{j}$) is a definable open subset of $Z_{i}(K)$ (resp. $W_{j}(K)$) for some affine varieties $Z_{i}$ and $W_{j}$ over $K$. Refining the decomposition if necessary, one shows easily that we may assume each $Z_{i}$ and $W_{j}$ is irreducible over $K$ and contained in $Z$ (note that irreducibility is only ensured over $K$, so we cannot assume absolute irreducibility).
Now, since $K$ is perfect, varieties over $K$ are generically smooth. That is, if $Z^{S}$ is the (Zariski open) smooth locus of $Z$, then $\dim(Z-Z^{S})<\dim(Z)$. We now define $T$ to be the union of $Z-Z^{S}$ with all those $Z_{i}$ and $W_{j}$ of dimension less than $\dim(Z)$. Clearly, then, $\dim(T)<\dim(Z)$.
To see this works, let $x\in Z(K)-T(K)$. We show that $x$ is polished for $X$. Without loss of generality, assume $x\in X$ (the case $x\in Z-X$ is symmetric). So $x\in U_{i}$ for some $i$. By the choice of $T$, $\dim(Z_{i})\geq\dim(Z)$. But by assumption, $Z_{i}$ is an irreducible closed subvariety of $Z$. It follows that $Z_{i}$ is an irreducible component of $Z$.
Recall that to show $x$ is polished, we need to find a neighborhood of $x$ (in $Z(K)$) which is contained in $X\cap Z^{S}(K)$. We claim that $U=U_{i}\cap Z^{S}(K)$ works. Clearly $U\subset X\cap Z^{S}(K)$, so all we need to show is that $U$ is relatively open in $Z(K)$.
For this, we establish the following three claims:
Claim 9.20.
$U$ is relatively open in $Z_{i}(K)\cap Z^{S}(K)$.
Proof.
By assumption, $U_{i}$ is open in $Z_{i}(K)$. Thus, automatically, $U=U_{i}\cap Z^{S}(K)$ is open in $Z_{i}(K)\cap Z^{S}(K)$. ∎
Claim 9.21.
$Z_{i}(K)\cap Z^{S}(K)$ is relatively open in $Z^{S}(K)$.
Proof.
Note that (by generic smoothness and the fact that $\dim(Z_{i})=\dim(Z)$) $Z_{i}\cap Z^{S}$ is an irreducible component of the smooth variety $Z^{S}$. Since smooth Zariski-connected varieties are irreducible, the irreducible components of $Z^{S}$ are relatively Zariski open. Thus $Z_{i}\cap Z^{S}$ is Zariski open in $Z^{S}$; and since $\tau$ refines the Zariski topology, $Z_{i}(K)\cap Z^{S}(K)$ is relatively open in $Z^{S}(K)$. ∎
Claim 9.22.
$Z^{S}(K)$ is relatively open in $Z(K)$.
Proof.
Since $\tau$ refines the Zariski topology, it is enough to note that $Z^{S}$ is automatically Zariski open in $Z$. ∎
By the three claims, $U$ is relatively open in $Z(K)$, completing the proof of Lemma 9.19. ∎
We now use Lemma 9.19 to characterize the dimension function of $\mathcal{K}$:
Lemma 9.23.
Let $X\subset K^{n}$ be definable, with Zariski closure $Z$. Then $\dim(X)=\dim Z(K)=\dim Z$.
Proof.
It is easy to see that $\dim(X)\leq\dim(Z(X))\leq\dim(Z)$. We show that $\dim(X)\geq\dim(Z)$.
Let $T$ be a relatively closed subvariety as in Lemma 9.19. By the choice of $Z$, $X$ is Zariski dense in $Z$. In particular, there is some $x\in X-T(K)$. So, by the choice of $T$, $x$ is polished for $X$. It follows by definition that $X$ contains a non-empty open subset of $Z^{S}(K)$, where $Z^{S}$ is the smooth locus of $Z$. So by Lemma 9.17, $\dim(X)\geq\dim(Z^{S})$. But by generic smoothness over perfect fields, $\dim(Z^{S})=\dim(Z)$, completing the proof. ∎
Finally, we deduce our main technical lemma:
Lemma 9.24.
Let $X\subset K^{n}$ be definable over a countable set $A$. Let $Z$ be the Zariski closure of $X$. If $x\in Z(K)$ is generic over $A$, then $x$ is polished for $X$.
Proof.
Let $T$ be a closed subvariety of $Z$ as in Lemma 9.19. Let $B\supset A$ be countable so that $T(K)$ is definable over $B$.
Let $y$ be an independent realization of $\operatorname{tp}(x/A)$ over $B$. So $y\in Z(K)$, and by Lemma 9.23, we have
$\dim(y/B)=\dim(x/A)=\dim(Z)>\dim(T)\geq\dim(T(K)),$
so that $y\notin T(K)$. Thus $y$ is polished for $X$. Let $\phi(z,b)$ (for $z$ in the arity of $Z(K)$) be a formula defining a neighborhood of $y$ witnessing that $y$ is polished for $X$. Since $\operatorname{tp}(x/A)=\operatorname{tp}(y/A)$, there is $a$ with $\operatorname{tp}(xa/A)=\operatorname{tp}(yb/A)$. Then by the invariance of $\tau$ (Definition 9.7), it follows that $\phi(z,a)$ defines a neighborhood of $x$ witnessing that $x$ is polished for $X$. ∎
Remark 9.25.
Lemma 9.24 says that, near generic points, we can treat definable sets as smooth varieties. More precisely, if $x\in X$ is generic over $A$, then $x$ is also generic in $S(K)$ for some smooth variety $S$ having the same dimension and germ as $X$ at $x$. Thus, if we are only interested in d-local properties of $(X,x)$, we may replace $X$ with $S(K)$. If we are also willing to pass from $A$ to $\operatorname{acl}(A)$ (which does not affect the genericity of $x$), then we may further assume $S$ is irreducible over $A$. This will be our main strategy moving forward.
9.4. Homeomorphisms and Reduction to Generically Smooth Maps
One of the difficulties of the trichotomy in the positive characteristic case is the existence of finite everywhere-ramified morphisms. As has become standard, we will get around this problem by decomposing an arbitrary map into a universal homeomorphism composed with a generically smooth map. Then, in studying topological properties, we can pay attention only to the generically smooth part.
Such a decomposition was given explicitly in [WY23], using the reduced relative Frobenius construction. This is not quite good enough for us, because we need the universal homeomorphism to preserve normality. We thus give a more intricate argument:
Lemma 9.26.
Let $V\rightarrow W$ be a dominant morphism of irreducible varieties over $K$. Then there are an irreducible variety $V^{\prime}$, and a factorization $V\rightarrow V^{\prime}\rightarrow W$, such that:
  1. (1)
    $V\rightarrow V^{\prime}$ is a universal homeomorphism, and $V(K)\rightarrow V^{\prime}(K)$ is a $\tau$-homeomorphism.
  2. (2)
    $V^{\prime}\rightarrow W$ is generically smooth.
  3. (3)
    If $V$ is normal then so is $V^{\prime}$.
Proof.
It is shown in [WY23, Lemma 1.12, Fact 3.3 and 3.5] that there is such a factorization satisfying (1) and (2), given by a sufficient iterate of the reduced relative Frobenius of $V\rightarrow W$ (the fact that this construction gives a $\tau$-homeomorphism follows from Lemma 9.15(2)). Call this factorization $V\rightarrow V_{1}\rightarrow W$. Now suppose $V$ is normal (and note that $V_{1}$ might not be). We show how to refine the construction to satisfy (3).
Let $N\rightarrow V_{1}$ be the normalization of $V_{1}$. So $N\rightarrow V_{1}$ is surjective and birational. Since $V$ is normal, and by the universal property of normalizations [Sta24, Lemma 035Q(4)], $V\rightarrow V_{1}$ factors as $V\rightarrow N\rightarrow V_{1}$. Since dominant quasi-finite morphisms to normal varieties are universally open [Sta24, Lemma 0F32], the image of $V$ in $N$ is an open (so also normal) subvariety, say $U$.
We now have $V\rightarrow U\rightarrow N\rightarrow V_{1}\rightarrow W$, where
  1. (1)
    $V\rightarrow U$ is surjective and universally open.
  2. (2)
    $U\rightarrow N$ is an open immersion.
  3. (3)
    $N\rightarrow V_{1}$ is surjective and birational.
  4. (4)
    $V_{1}\rightarrow W$ is generically smooth.
Since $V\rightarrow V_{1}$ is a universal homeomorphism, it is universally injective. Thus $V\rightarrow U$ is also universally injective. Note that base surjections are automatically universally surjective [Sta24, Lemma 01S1]. Thus, $V\rightarrow U$ is universally open, universally injective, and universally surjective, and so is a universal homeomorphism.
We now apply the reduced relative Frobenius again, this time to $V\rightarrow U$. We obtain a factorization $V\rightarrow V^{\prime}\rightarrow U$ again satisfying (1) and (2) of the lemma (for $V\rightarrow U$). Thus, we have dominant morphisms
$V\rightarrow V^{\prime}\rightarrow U\rightarrow N\rightarrow V_{1}\rightarrow W.$
To complete the proof of the lemma, we note:
Claim 9.27.
$V\rightarrow V^{\prime}$ is a universal homeomorphism and a $\tau$-homeomorphism.
Proof.
This is because $V\rightarrow V^{\prime}\rightarrow U$ satisfies (1) and (2). ∎
Claim 9.28.
$V^{\prime}\rightarrow W$ is generically smooth.
Proof.
Because $V^{\prime}\rightarrow U\rightarrow N\rightarrow V_{1}\rightarrow W$ is a chain of generically smooth maps. ∎
And, the main point:
Claim 9.29.
$V^{\prime}$ is normal.
Proof.
We have that $V\rightarrow U$ and $V\rightarrow V^{\prime}$ are both universal homeomorphisms, which implies that $V^{\prime}\rightarrow U$ is also a universal homeomorphism [Sta24, Lemma 0H2M], and thus (by generic smoothness) a birational universal homeomorphism. But then, as a birational universal homeomorphism to a normal variety, $V^{\prime}\rightarrow U$ is an isomorphism [Sta24, Lemma 0AB1]. Thus $V^{\prime}$ is normal because $U$ is. ∎
Remark 9.30.
Suppose, in the setup of Lemma 9.26, the morphism $V(K)\rightarrow W(K)$ is defined over $A$. We note that each step in the proof of the lemma was given by a functorial construction. It then follows that the whole sequence $V(K)\rightarrow V^{\prime}(K)\rightarrow W(K)$ is still defined over $A$. We will use this implicitly.
9.5. Proofs of the Axioms
Armed with the machinery from the previous subsection, we now show that $(\mathcal{K},\tau)$ is a Hausdorff geometric structure. We also give conditions under which $(\mathcal{K},\tau)$ has enough open maps.
Theorem 9.31.
$(\mathcal{K},\tau)$ is a Hausdorff geometric structure.
Proof.
We already showed that $\mathcal{K}$ is geometric. Thus, it suffices to establish the strong frontier inequality, the Baire category axiom, and the generic local homeomorphism property. These are given in the next three claims.
Claim 9.32.
Let $X\subset K^{n}$ be definable over $A$, and let $a\in\overline{\operatorname{Fr}(X)}$. Then $\dim(a/A)<\dim(X)$.
Proof.
Let $Z$ be the Zariski closure of $X$. Note that $a\in Z(K)$, since $a\in\overline{X}$ and $\tau$ refines the Zariski topology. By Lemma 9.23, $\dim(Z(K))=\dim(X)$. Now assume toward a contradiction that $\dim(a/A)=\dim(X)$. Then $\dim(a/A)=\dim(Z(K))$, so $a$ is generic in $Z(K)$. By Lemma 9.24, $a$ is thus polished for $X$. But the definition of being polished for $X$ implies that $a\notin\overline{\operatorname{Fr}(X)}$, a contradiction. ∎
Claim 9.33.
Let $X\subset K^{n}$ be definable over $A$, and let $a\in X$ be generic over $A$. Let $B\supset A$ be countable. Then every neighborhood of $a$ contains a generic of $X$ over $B$.
Proof.
Let $Z$ be the Zariski closure of $X$, and let $Z^{S}$ be its smooth locus. By Lemma 9.23, $\dim(X)=\dim(Z(K))$. So since $a$ is generic in $X$ over $A$, it is also generic in $Z(K)$ over $A$. So by Lemma 9.24, $a$ is polished for $X$. Let $U\subset Z(K)$ be a neighborhood of $a$ witnessing this. That is, $U$ is open in $Z(K)$, and $U$ is contained in $X\cap Z^{S}(K)$.
To prove the claim, let $V$ be any neighborhood of $a$ in $X$. Shrinking if necessary, we may assume $V$ is definable and contained in $U$, so in particular $V$ is relatively open in $Z^{S}(K)$. Adding to $B$ if necessary, we may assume $V$ is definable over $B$.
Now by Lemmas 9.17 and 9.23, and generic smoothness over perfect fields, we conclude that
$\dim(V)=\dim(Z^{S})=\dim(Z)=\dim(X).$
In particular, if we choose any $b\in V$ generic over $B$, then $b$ is also generic in $X$ over $B$, proving the claim. ∎
Claim 9.34.
Let $Z\subset X\times Y$ be definable over $A$ and all of dimension $d$, with $Z\rightarrow X$ and $Z\rightarrow Y$ finite-to-one. Let $(x,y)\in Z$ be generic over $A$. Then $Z$ restricts to a homeomorphism between neighborhoods of $x$ and $y$ in $X$ and $Y$, respectively.
Proof.
It is enough to show the projection $Z\rightarrow X$ is locally a homeomorphism near $(x,y)$, since by symmetry the same will apply to $Z\rightarrow Y$. For this, it suffices to show separately that $Z\rightarrow X$ is locally injective, locally continuous, and locally open at $(x,y)$. Moreover, local continuity is automatic, since $Z\rightarrow X$ is a projection. So we show local injectivity and local openness.
First, suppose $Z\rightarrow X$ is not locally injective at $(x,y)$. Then $(x,y,y)\in\operatorname{Fr}(T)$, where $T$ is the set of $(u,v,w)$ with $(v\neq w)$ and $(u,v),(u,w)\in Z$. Since $Z\rightarrow X$ is finite-to-one, $\dim(T)\leq d$. Then by the strong frontier inequality, we get $\dim(xy/A)<d$, a contradiction. So $Z\rightarrow X$ is locally injective near $(x,y)$.
Now we show local openness at $(x,y)$. The idea is that, as outlined in the beginning of this subsection, we can reduce to the case that $Z\rightarrow X$ is generically smooth – thus étale – and then apply Definition 9.8(3). More precisely, by Lemma 9.24, there are smooth d-dimensional varieties $W$ and $V\subset\mathbb{A}^{n}\times W$ so that:
  1. (1)
    $V(K)$ and $W(K)$ are definable over $A$.
  2. (2)
    $V(K)$ and $W(K)$ have the same germs as $Z$ and $X$ at $(x,y)$ and $x$, respectively.
We may assume, without loss of generality, that $A=\operatorname{acl}(A)$. Thus, we may also assume $V$ and $W$ are irreducible. Since $Z\rightarrow X$ is finite-to-one, it follows that $V\rightarrow W$ is dominant.
Now let $V\rightarrow V^{\prime}\rightarrow W$ be a factorization as in Lemma 9.26, and let $z$ be the image of $(x,y)$ in $V^{\prime}(K)$. It follows from Remark 9.30 that $z$ is generic in $V^{\prime}(K)$. Moreover, since $V(K)\rightarrow V^{\prime}(K)$ is a homeomorphism, it suffices to show that $V^{\prime}(K)\rightarrow W(K)$ is open near $z$. But by the choice of $V^{\prime}$ (and the fact that $\dim(V^{\prime})=\dim(W)$, $V^{\prime}\rightarrow W$ is generically étale; and since $z\in V^{\prime}$ is generic, it follows that $V^{\prime}\rightarrow W$ is étale in a Zariski neighborhood of $z$. Local openness then follows as promised from Definition 9.8(3). ∎
9.6. Smoothness and Enough Open Maps
We now proceed to give conditions under which $(\mathcal{K},\tau)$ is either differentiable or has the open mapping property. First, we define a canonical notion of smoothness.
Definition 9.35.
For definable $X\subset K^{n}$, we let $X^{S}$ be the set of $x\in X$ which are polished for $X$ – that is, the relative interior of $X$ in the smooth locus of its Zariski closure.
In other words, the smooth points of $X$ are those points where we can view $X$ d-locally as a smooth variety. We now show:
Lemma 9.36.
The map $X\mapsto X^{S}$ is a notion of smoothness on $(\mathcal{K},\tau)$.
Proof.
Most of the properties in Definition 3.27 follow from standard facts about smooth varieties (e.g. closure under products and isomorphisms). The fact that generic points are smooth follows from Lemma 9.24. We show only Definition 3.27(5) (smoothness of preimages under appropriate maps).
To recall the setting: we have a projection $f:X\rightarrow Y$ of definable sets over $A$, and a point $x\in X$ generic over $A$, such that $y=f(x)$ is generic in $Y$ over $A$. We then have a set $Z$, definable over $B$, so that (1) $y$ is generic in $Z$ over $B$ and (2) the germ of $Z$ at $y$ is contained in the germ of $Y$ at $y$. Our goal is to show that $x\in(f^{-1}(Z\cap Y))^{S}$.
To do this, one follows exactly the same procedure as in the proof of Claim 9.34 to reduce everything d-locally to smooth varieties and smooth morphisms. In the end, one only needs to show the following:
Claim 9.37.
Let $g:V\rightarrow W$ be a smooth morphism of smooth varieties over $K$, and let $U\subset W$ be a smooth subvariety. Then $g^{-1}(U)$ is smooth.
Proof.
Recall that smooth morphisms are stable under composition and base change [Sta24, Lemma 01VA,01VB]. Moreover, since $K$ is perfect, a variety over $K$ is smooth if and only if its structure morphism to $\operatorname{Spec}(K)$ is smooth. We proceed to repeatedly use these facts.
First, the map $g^{-1}(U)\rightarrow U$ is isomorphic to the base change of $V\rightarrow W$ by $U\rightarrow W$. Since $V\rightarrow W$ is smooth, so is $g^{-1}(U)\rightarrow U$. But $U\rightarrow\operatorname{Spec}(K)$ is smooth by assumption, so the composition $g^{-1}(U)\rightarrow\operatorname{Spec}(K)$ is smooth, and thus $g^{-1}(U)$ is smooth. ∎
We now, automatically, have:
Lemma 9.38.
Suppose that whenever $g:V\rightarrow W$ is a quasi-finite morphism of smooth $d$-dimensional varieties over $K$, the induced map $g:V(K)\rightarrow W(K)$ is open. Then $(\mathcal{K},\tau)$ has the open mapping property.
Proof.
Let $f:X\rightarrow Y$ be a finite-to-one definable projection, where $\dim(X)=\dim(Y)$, and let $x\in X^{S}$ with $y=f(x)\in Y^{S}$. We need to show that $f$ is open near $x$. But this is automatic: the assumption that $x\in X^{S}$ and $y\in Y^{S}$ lets us reduce d-locally to a projection of smooth varieties, and we apply the hypothesis of the lemma. ∎
Remark 9.39.
We note, for example, that the hypothesis of Lemma 9.38 holds of the complex field with the analytic topology (though technically, this structure is not an éz topological field expansion, because it has no invariant basis). Indeed, in this case, the induced map $g:V(K)\rightarrow W(K)$ is a holomorphic map of $d$-dimensional complex manifolds with all fibers discrete, and one can apply a usual open mapping theorem from complex analysis (e.g. [GR84], p. 107). This was one of the key geometric properties used in [Cas23]), and is the inspiration for Lemma 9.38. Indeed, we will see later (Corollary 9.60) that the hypothesis of the lemma holds in ACVF in all characteristics.
We now move on to differentiability. In fact, quite generally, we have:
Lemma 9.40.
If $K$ has characteristic zero, then $(\mathcal{K},\tau)$ is differentiable.
Proof.
Given a definable $X$ and $x\in X^{S}$, the germ of $X^{S}$ at $x$ agrees with a smooth variety $S(K)$, where $\dim(S)=\dim(X)$. We then set $T_{x}(X)$ to be the Zariski tangent space $T_{x}(S)$. All of the required properties of Definition 3.34 follow easily. Note that the weak inverse function theorem follows since étale maps are open (Definition 9.8(3)), and Sard’s theorem follows since dominant morphisms in characteristic zero are generically smooth. ∎
9.7. The Algebraically Closed Case and Purity of Ramification
We now address the special case that the underlying field $K$ is algebraically closed. In this case, we will prove a ‘topologized’ purity of ramification statement for morphisms of varieties, and subsequently deduce that $(\mathcal{K},\tau)$ has ramification purity in the sense of Definition 6.16.
Assumption 9.41.
Throughout this subsection, we assume that $K$ is algebraically closed.
The main advantages of the algebraically closed case are summarized by the following fact and corollary:
Fact 9.42.
Let $V$ be an irreducible variety over $K$, and let $U\subset V(K)$ be definable, relatively open, and non-empty. Then $U$ is Zariski dense in $V$. In particular, $\dim(U)=\dim(V)$.
Proof.
A proof can be found in [MB12b, Proposition 2.1.1]. ∎
Fact 9.42 implies that $\tau$ behaves identically to the Zariski topology on the level of constructible sets:
Corollary 9.43.
Let $V$ be a variety over $K$ and let $X\subset V$ be constructible (equivalently, definable in the pure field language). Then the relative Zariski and $\tau$-closures of $X$ in $V(K)$ coincide.
We will also use the following fact about rings:
Proposition 9.44.
Let $A\leq B$ be a flat extension of integral domains, and suppose the extension of fraction fields $\operatorname{Frac}(A)\leq\operatorname{Frac}(B)$ is finite and separable. Then the tensor product $B\otimes_{A}B$ is a reduced ring.
Proof.
For convenience, let $K_{A}$ and $K_{B}$ denote the respective fraction fields. The idea of the proof is to use flatness to embed $B\otimes_{A}B$ into $K_{B}\otimes_{K_{A}}K_{B}$, so that we can assume $A$ and $B$ are already fields. The next claim follows easily from flatness:
Claim 9.45.
$B\otimes_{A}B$ embeds into $K_{B}\otimes_{A}K_{B}$.
Let $S=A-\{0\}$, so that $K_{A}$ is the localization $S^{-1}A$. Note that $K_{B}$ is already a $K_{A}$-module, so $S^{-1}K_{B}\cong K_{B}$. Moreover, the following is standard:
Claim 9.46.
$S^{-1}(K_{B}\otimes_{A}K_{B})\cong K_{B}\otimes_{A}K_{B}$.
Finally, we conclude:
Claim 9.47.
$B\otimes_{A}B$ embeds into $K_{B}\otimes_{K_{A}}K_{B}$. In particular, in proving Proposition 9.44, we may assume $A$ and $B$ are fields.
Proof.
Recall [Sta24, Lemma 00DL] that we have a canonical isomorphism
$S^{-1}K_{B}\otimes_{S^{-1}A}S^{-1}K_{B}\cong S^{-1}(K_{B}\otimes_{A}K_{B}),$
which simplifies to
$K_{B}\otimes_{K_{A}}K_{B}\cong S^{-1}(K_{B}\otimes_{A}K_{B}).$
By Claim 9.46, we moreover obtain an isomorphism $K_{B}\otimes_{K_{A}}K_{B}\cong K_{B}\otimes_{A}K_{B}$. Then combining with Claim 9.45 gives an embedding of $B\otimes_{A}B$ into $K_{B}\otimes_{K_{A}}K_{B}$.
Now to reduce to the case that $A$ and $B$ are fields, simply note that if $B\otimes_{A}B$ had nilpotents, then so would $K_{B}\otimes_{K_{A}}K_{B}$. ∎
Now assume that $A$ and $B$ are fields, so that $A\leq B$ is a finite separable extension. By the primitive element theorem, we have $B=A(b)$ for some $b$. Let $p(x)$ be the minimal polynomial of $b$ over $A$. One then sees easily that
$B\otimes_{A}B\cong A[x,y]/(p(x),p(y))\cong B[y]/(p(y)).$
By separability, $p$ factors over $B$ into a product of distinct irreducibles $p_{i}$. Then by the Chinese Remainder Theorem, our tensor product is now $\Pi_{i}B[y]/p_{i}(y)$, which is a product of fields, and thus reduced. ∎
We now move toward our version of purity of ramification. First, recall the classical purity of ramification in algebraic geometry (see, e.g., [Sta24, Lemma 0EA4]):
Fact 9.48.
Let $f:V\rightarrow W$ be a dominant morphism of irreducible varieties over $K$. Assume that $V$ is normal and $W$ is smooth. Then every irreducible component of the non-étale locus of $f$ has dimension at least $\dim(V)-1$.
Recall that a function $f:X\rightarrow Y$ between topological spaces is topologically unramified at $x\in X$ if $f$ is injective on some neighborhood of $x$. Our goal is to prove an analog of Fact 9.48 for $(\mathcal{K},\tau)$, where we replace ‘étale’ with ‘topologically unramified’ (for the topology $\tau$).
The proof will consist, essentially, of two steps. First, we use Fact 9.26 to reduce to the case that $f$ is generically smooth. Then we show that, in the generically smooth case, the topological ramification locus coincides with the non-étale locus up to a codimension 2 error. Using these two steps, one can reduce our statement to Fact 9.48.
The first of the two steps above is automatic. For the second step, we use the following technical fact:
Proposition 9.49.
Let $V$ and $W$ be smooth irreducible varieties over $K$, and let $f:V\rightarrow W$ be a dominant, generically smooth morphism. Then the following are equivalent:
  1. (1)
    $f$ is étale.
  2. (2)
    The restricted map $f:V(K)\rightarrow W(K)$ is topologically unramified.
Proof.
Throughout, let $Z$ be the fiber product $V\times_{W}V$ (as schemes). Let $Z_{red}$ be the reduced scheme associated to $Z$ (i.e. the fiber product as varieties). So $Z_{red}(K)$ is the set of $(x,y)\in V(K)^{2}$ with $f(x)=f(y)$. Moreover, let $\Delta$ be the diagonal (viewed as a subvariety of $Z_{red}$ – equivalently the image of $V$ in $Z_{red}$). Finally, let $\delta:V\rightarrow Z$ be the diagonal morphism. Since $V$ is a variety (thus reduced), $\delta$ factors as $V\rightarrow Z_{red}\rightarrow Z$.
First suppose $f$ is étale. Then $f$ is unramified, so $\delta$ is an open immersion, and thus $\Delta$ is Zariski open in $Z_{red}$. By Corollary 9.43, $\Delta(K)$ is $\tau$-open in $Z_{red}(K)$. This is a restatement of the assertion that $f:V(K)\rightarrow W(K)$ is topologically unramified.
Now assume $f:V(K)\rightarrow W(K)$ is topologically unramified. As above, we get that $\Delta(K)$ is $\tau$-open in $Z_{red}(K)$, and thus $\Delta$ is Zariski open in $Z_{red}$. We first conclude:
Lemma 9.50.
$f$ is quasifinite and flat.
Proof.
Since $\Delta$ is Zariski open in $Z_{red}$, it follows that for each $x\in V(K)$, $\{(x,x)\}$ is open in $(Z_{red})_{x}(K)=\{y:(x,y)\in Z_{red}(K)\}$. Equivalently, $\{x\}$ is open in the fiber $f^{-1}(f(x))$. Thus $f$ has discrete fibers, and so is quasifinite. Flatness then follows by Miracle Flatness [Sta24, Lemma 00R4] by looking at the induced maps of local rings between points. ∎
By Lemma 9.50, we have:
Lemma 9.51.
$Z$ is reduced, and thus $Z_{red}\rightarrow Z$ is an isomorphism.
Proof.
After reducing to affine opens, this reduces exactly to Proposition 9.44. In particular, the fact that $V=\mathrm{Spec}B\rightarrow\mathrm{Spec}A=W$ is quasifinite and generically smooth guarantees that the field extension $\operatorname{Frac}(A)\leq\operatorname{Frac}(B)$ appearing in Proposition 9.44 is finite separable. ∎
It remains to show that $f$ is unramified, that is, that $\delta$ is an open immersion. But $\delta$ is automatically a (locally closed) immersion [Sta24, Lemma 01KJ], namely it factorizes as $j\circ i$ where $i$ is a closed immersion and $j$ is an open immersion. And we have shown it has open image; By the irreducibility and the fact that the only surjective closed immersion to a reduced scheme is an isomorphism, we have that we can assume $i$ is an isomorphism. Thus, by Lemma 9.51, we are done. ∎
Finally, we are ready to ‘topologize’ Fact 9.48. In what follows, it will be convenient to extend the notion of local dimension to non-definable sets:
Definition 9.52.
Let $V$ be a variety over $K$, and $X\subset V(K)$ arbitrary. Let $x\in X$, and $d$ a non-negative integer. We say that $X$ has local dimension at least $d$ at $x$ if every relative neighborhood of $x$ in $X$ contains a definable set of dimension $d$.
We now show:
Theorem 9.53.
Let $f:V\rightarrow W$ be a dominant morphism of irreducible varieties over $K$. Assume that $V$ is normal and $W$ is smooth. Let $R$ be the topological ramification locus of $f$. Then $R$ has local dimension at least $\dim(V)-1$ at every point.
Proof.
First note that, by Lemma 9.26, we may assume $f$ is generically smooth (since homeomorphisms preserve the topological ramification locus). In particular, the variety $V^{\prime}$ obtained by Lemma 9.26 is still normal, so the hypotheses of the theorem still hold.
Thus, moving forward, we assume $f$ is generically smooth. As in Proposition 9.49, let $Z$ be the fiber product $V\times_{W}V$ (as schemes), and let $Z_{red}$ be the associated reduced scheme. Then let $\Delta$ be the diagonal subvariety of $Z_{red}$, and $\delta:V\rightarrow Z$ the diagonal morphism, which factors as $V\rightarrow Z_{red}\rightarrow Z$.
Toward a proof of the theorem, let $x\in R$, and let $U_{1}$ be any relative neighborhood of $x$ in $R$. So $U_{1}=U\cap R$ for some open $U\subset V(K)$. Shrinking $U_{1}$ and $U$ if necessary, we may assume $U$ is definable (in fact so are $U_{1}$ and $R$, but this won’t be directly used). Our goal is to find a definable $X\subset U_{1}$ with $\dim(X)=\dim(V)-1$.
First, since $f$ topologically ramifies at $x$, one sees that $(x,x)$ is not in the $\tau$-interior of $\Delta(K)$ in $Z_{red}(K)$. By Corollary 9.43, $(x,x)$ is not in the Zariski interior of $\Delta$ in $Z_{red}$. It follows easily that $V\rightarrow Z_{red}$ is not an open map in any neighborhood of $x$. But $Z_{red}\rightarrow Z$ is a Zariski homeomorphism, so $\delta$ is also not an open map in any neighborhood of $x$. Thus, $\delta$ is not locally an open immersion at $x$, and thus $f$ ramifies at $x$ [Sta24, Lemma 02GE].
So $f$ is not étale at $x$. By purity of ramification (Fact 9.48), there is an irreducible closed codimension 1 subvariety $T\subset V$ such that $x\in T$ and $f$ is not étale on any point of $T$. Let $V^{S}$ be the smooth locus of $V$. Let $T^{\prime}$ be the open subvariety $T\cap V^{S}\subset T$. Finally, let $X=U\cap T^{\prime}(K)$. So $X$ is definable because $U$ is. To conclude, we need the following two claims:
Claim 9.54.
$f$ topologically ramifies at every point of $X$, and thus $X\subset U_{1}$.
Proof.
By Proposition 9.49 applied to the restricted map $V^{S}(K)\rightarrow W(K)$, it suffices to observe (by the choice of $T$) that $f$ is non-étale at every point of $T$. ∎
Claim 9.55.
$\dim(X)=\dim(V)-1$.
Proof.
Since normal varieties are smooth outside codimension 2, we have
$\dim(T-V^{S})\leq\dim(V-V^{S})\leq\dim(V)-2<\dim(V)-1=\dim(T),$
and thus $\dim(T^{\prime})=\dim(T)$. So $T^{\prime}$ is an open subvariety of $T$ of the same dimension; since $T$ is irreducible, this implies $T^{\prime}$ is Zariski dense in $T$. Then, by Corollary 9.43, $T^{\prime}(K)$ is also $\tau$-dense in $T(K)$. In particular, since $U\cap T(K)$ is non-empty (witnessed by $x$), it follows that $U\cap T^{\prime}(K)$ is also non-empty, so that (by Fact 9.42)
$\dim(X)=\dim(U\cap T^{\prime}(K))=\dim(T^{\prime})=\dim(V)-1.$
Finally, we proceed to the proof of ramification purity. The first step is the following approximation of the full statement:
Lemma 9.56.
Let $\mathcal{Y}=\{Y_{t}:t\in U\}$ and $\mathcal{Z}=\{Z_{u}:u\in U\}$ be $\mathcal{K}(A)$-definable families of subsets of a $\mathcal{K}(A)$-definable set $X$, with graphs $Y\subset X\times T$ and $Z\subset X\times U$. Let $I$ be the set of $(x,t,u)$ with $(x,t)\in Y$ and $(x,u)\in Z$, and let $R$ be the topological ramification locus of $I\rightarrow T\times U$. Suppose $(x_{0},t_{0},u_{0})$ is a weakly generic $(\mathcal{Y},\mathcal{Z})$-multiple intersection over $A$ which is strongly approximable over $A$. Then $R$ has local dimension at least $d-1$ at $(x_{0},t_{0},u_{0})$, where $d=\dim(Y)+\dim(Z)-\dim(X)$.
Proof.
The idea is to reduce all sets and maps in the statement to smooth varieties and smooth morphisms, and then apply Theorem 9.53.
First, note that $x_{0}\in X^{S}$, $t_{0}\in T^{S}$, $u_{0}\in U^{S}$, $(x_{0},t_{0})\in Y^{S}$, and $(x_{0},u_{0})\in Z^{S}$ (since each point is generic in the corresponding set). Moreover, we are only interested in d-local data of $X$, $Y$, $Z$, and $I$ near $(x_{0},t_{0},u_{0})$. So it is harmless to assume each of $X$, $Y$, $Z$, $T$, and $U$ is the set of $K$-points of a smooth variety. It is also harmless to replace $A$ with $\operatorname{acl}(A)$, so we may assume these smooth varieties are all irreducible. For ease of notation, let us call these varieties $X$, $Y$, $Z$, $T$, and $U$, and rename the original definable sets as $X(K)$, $Y(K)$, $Z(K)$, $T(K)$, and $U(K)$.
The assumption that $(x,t)\in Y$ and $x\in X$ are generic over $A$ gives that $Y\rightarrow X$ is dominant. Similarly, $Z\rightarrow X$ is dominant. Now, it is harmless to apply a Frobenius power to each of $T$ and $U$, since we are only interested in topological data. So by Lemma 9.26, we may assume the morphisms $Y\rightarrow X$ and $Z\rightarrow X$ are generically smooth. In particular, by the genericity of $(x,t)\in Y$ and $(x,u)\in Z$, each of $Y\rightarrow X$ and $Z\rightarrow X$ is smooth in a Zariski neighborhood of the relevant point. It is harmless to shrink $Y$ and $Z$ so that $Y\rightarrow X$ and $Z\rightarrow X$ are each smooth.
In this case, $I\rightarrow X$ is the fiber product of the two smooth morphisms $Y\rightarrow X$ and $Z\rightarrow X$. Thus $I\rightarrow X$ is also smooth. Since $X$ is smooth, this implies $I$ is also smooth. Let $I^{\prime}$ be an irreducible component of $I$ containing $(x_{0},t_{0},u_{0})$. Since we are assuming $A=\operatorname{acl}(A)$, $I^{\prime}(K)$ is definable over $A$.
We now consider the morphism $I^{\prime}\rightarrow T\times U$ of irreducible smooth varieties. We claim this morphism is dominant. To see this, note that since $(x_{0},t_{0},u_{0})$ is strongly approximable, $I^{\prime}(K)$ contains a strongly generic $(\mathcal{Y},\mathcal{Z})$-intersection $(x,t,u)$ over $A$. By definition of strong genericity, $(t,u)$ is generic in $T(K)\times U(K)$ over $A$. This shows that $I^{\prime}(K)\rightarrow T(K)\times U(K)$ has generic image, and this implies dominance.
We have now set up Theorem 9.53 for the morphism $I^{\prime}\rightarrow T\times U$. Let $R^{\prime}$ be the topological ramification locus of $I^{\prime}\rightarrow T\times U$ (and note that $R^{\prime}\subset R$). Theorem 9.53 tells us that $R^{\prime}$ has local dimension at least $\dim(I^{\prime})-1$ at $(x_{0},t_{0},u_{0})$. Since $R^{\prime}\subset R$, we conclude that $R$ also has local dimension at least $\dim(I^{\prime})-1$ at $(x_{0},t_{0},u_{0})$.
Finally, to prove the lemma, we show that $\dim(I^{\prime})\geq d=\dim(Y)+\dim(Z)-\dim(X)$. For this, we again use our strongly generic intersection $(x,t,u)$. Indeed, an easy computation, using strong genericity, yields that $\dim(xtu/A)=d$. Meanwhile, since $(x,t,u)\in I^{\prime}(K)$, we have $\dim(xtu/A)\leq\dim(I^{\prime})$, and thus $\dim(I^{\prime})\geq d$ as desired. ∎
We next deduce a slightly stronger statement:
Lemma 9.57.
Let $\mathcal{Y}=\{Y_{t}:t\in T\}$ and $\mathcal{Z}=\{Z_{u}:u\in U\}$ be $\mathcal{K}(A)$-definable families of subset of a $\mathcal{K}(A)$-definable set $X$, with graphs $Y\subset X\times T$ and $Z\subset X\times U$. Let $(x_{0},t_{0},u_{0})$ be a generic $(\mathcal{Y},\mathcal{Z})$-multiple intersection over $A$ which is strongly approximable over $A$. Then there is a generic $(\mathcal{Y},\mathcal{Z})$-multiple intersection $(x,t,u)$ satisfying $\dim(xtu/A)\geq d-1$, where $d=\dim(Y)+\dim(Z)-\dim(X)$.
Proof.
Let $I$ be the set of $(x,t,u)$ with $(x,t)\in Y$ and $(x,u)\in Z$, and let $R$ be the topological ramification locus of $I\rightarrow T\times U$. We are asked to find $(x,t,u)$ satisfying:
  1. (1)
    $(x,t,u)\in R$.
  2. (2)
    $x$, $t$, $u$, $(x,t)$, and $(x,u)$ are generic in $X$, $T$, $U$, $Y$, and $Z$ over $A$, respectively.
  3. (3)
    $\dim(xtu/A)\geq d-1$.
Since $\mathcal{K}$ is $\aleph_{1}$-saturated, it suffices to realize any finite fragment of (1)-(3) above. In particular, we will find $(x,t,u)$ realizing (1), (3), and a finite fragment of (2).
Let us briefly elaborate. For the purposes of this proof, given $\mathcal{K}$-definable sets $D_{1}\subset D_{2}$, let us call $D_{1}$ large in $D_{2}$ if $\dim(D_{2}-D_{1})<\dim(D_{2})$. It is easy to see that, if $D$ is $A$-definable, then the set of $A$-generic elements of $D$ is the intersection of all $A$-definable large subsets of $D$. In particular, to show some partial type is consistent with the $A$-generic locus of $D$, one can show that it is consistent with each $A$-definable large subset of $D$. So our plan is to satisfy (1), (3), and an approximation of (2) replacing each $\mathcal{K}(A)$-definable set in the statement with a $\mathcal{K}(A)$-definable large subset.
Now let us give the details. Suppose $X^{\prime}$, $Y^{\prime}$, $Z^{\prime}$, $T^{\prime}$, and $U^{\prime}$ are any $\mathcal{K}(A)$-definable large subsets of $X$, $Y$, $Z$, $T$, and $U$, respectively. Then, precisely, it suffices to find $(x,t,u)$ satisfying (1), (2’), and (3), where (2’) is the conjunction of $x\in X^{\prime}$, $(x,t)\in Y^{\prime}$, $(x,u)\in Z^{\prime}$, $t\in T^{\prime}$, and $u\in U^{\prime}$.
We now construct such $(x,t,u)$. First, let $I^{\prime}$ be the set of $(x,t,u)\in X^{\prime}\times T^{\prime}\times U^{\prime}$ with $(x,t)\in Y^{\prime}$ and $(x,u)\in Z^{\prime}$. Let $R^{\prime}$ be the topological ramification locus of the projection $I^{\prime}\rightarrow T^{\prime}\times U^{\prime}$. Note that any $(x,t,u)\in R^{\prime}$ automatically satisfies (1) and (2’). Thus, to additionally satisfy (3), it suffices to find an element $(x,t,u)\in R^{\prime}$ with $\dim(xtu/A)\geq d-1$.
Now by Lemma 3.20, the sets $X$ and $X^{\prime}$ have the same germ at $x_{0}$ (that is, by genericity of $x_{0}$, they both realize the germ of $\operatorname{tp}_{\mathcal{K}}(x_{0}/A)$); and the analogous statements hold for $Y$, $Z$, $T$, and $U$. Since they are defined analogously, it follows that $I$ and $I^{\prime}$ have the same germ at $(x_{0},t_{0},u_{0})$ and thus that $R$ and $R^{\prime}$ have the same germ at $(x_{0},t_{0},u_{0})$. So $R^{\prime}$ contains some relative neighborhood of $(x_{0},t_{0},u_{0})$ in $R$. But by Lemma 9.56, every relative neighborhood of $(x_{0},t_{0},u_{0})$ in $R$ contains a definable set of dimension $d-1$. Thus, in particular, $R^{\prime}$ contains a definable set – say $D$ – of dimension $d-1$. Let $D$ be definable over $B\supset A$, and let $(x,t,u)\in D$ be generic over $B$. Then $(x,t,u)\in R^{\prime}$, and
$\dim(xtu/A)\geq\dim(xtu/B)=\dim(D)=d-1,$
as desired. ∎
Remark 9.58.
Note that in the above lemma, we cannot assume $R$ and $R^{\prime}$ are definable (this would require a uniformly definable basis for $\tau$, and we do not assume uniformity). This is the reason we needed the more general notion of local dimension in Definition 9.52.
Finally, we show:
Theorem 9.59.
$(\mathcal{K},\tau)$ has ramification purity.
Proof.
Let $\mathcal{Y}=\{Y_{t}:t\in T\}$ and $\mathcal{Z}=\{Z_{u}:u\in U\}$ be $\mathcal{K}(A)$-definable subsets of a $\mathcal{K}(A)$-definable set $X$, with graphs $Y\subset X\times T$ and $Z\subset X\times U$. Let $(x_{0},t_{0},u_{0})$ be a generic $(\mathcal{Y},\mathcal{Z})$-multiple intersection over $A$ which is strongly approximable over $A$. We want to find a generic $(\mathcal{Y},\mathcal{Z})$-multiple intersection over $A$ which has codimension at most 1 over $A$.
By Lemma 9.57, there is a generic $(\mathcal{Y},\mathcal{Z})$-multiple intersection $(x,t,u)$ over $A$ with $\dim(xtu/A)\geq d-1$, where $d=\dim(Y)+\dim(Z)-\dim(X)$. We claim that $\operatorname{codim}_{A}(xtu)\leq 1$, which will complete the proof. Indeed, we have
  • $\dim(x/A)=\dim(X)$
  • $\dim(xt/A)=\dim(Y)$,
and thus by additivity, $\dim(t/Ax)=\dim(Y)-\dim(X)$. Similarly, we get $\dim(u/Ax)=\dim(Z)-\dim(X)$.
Meanwhile, we have
  • $\dim(x/A)=\dim(X)$
  • $\dim(xtu/A)\geq d-1$,
and thus by additivity, $\dim(tu/Ax)\geq d-1-\dim(X)$. So, putting everything together, we have
$\operatorname{codim}_{A}(xtu)=\dim(t/Ax)+\dim(u/Ax)-\dim(tu/Ax)$
$\leq\dim(Y)-\dim(X)+\dim(Z)-\dim(X)-(d-1-\dim(X))$
$\dim(Y)+\dim(Z)-\dim(X)-(d-1)=1.$
Let us now drop the ambient assumptions on $(\mathcal{K},\tau)$. We end Section 9 by collecting our results for ACVF and pure éz fields:
Corollary 9.60.
Every $\aleph_{1}$-saturated algebraically closed valued field, equipped with the valuation topology, is a Hausdorff geometric structure with the open mapping property and ramification purity.
Proof.
Let $(\mathcal{K},\tau)$ be such a field. By Lemma 9.11, $(\mathcal{K},\tau)$ is a Hausdorff geometric structure. By Theorem 9.59, $(\mathcal{K},\tau)$ has ramification purity. Finally, to prove the open mapping property, it suffices to verify the hypothesis of Lemma 9.38. So, let $f:V\rightarrow W$ be a quasi-finite morphism of smooth $d$-dimensional $K$-varieties. Then by [Sta24, Lemma 0F32], $f$ is universally open, and thus by Fact 9.10, $V(K)\rightarrow W(K)$ is open. ∎
Corollary 9.61.
Let $(K,+,\cdot)$ be an $\aleph_{1}$-saturated éz field, whose étale open topology is induced by a field topology. Then $(K,+,\cdot)$, equipped with the étale open topology, is a Hausdorff geometric structure, and if $K$ has characteristic zero then $(K,\tau)$ is differentiable.
Proof.
By Theorem 9.31 and Lemma 9.40. ∎
10. Definable Slopes in ACVF
Our goal now is to show that algebraically closed valued fields have definable slopes satisfying TIMI (Definition 7.10 and Definition 7.50). Throughout this section, we fix an $\aleph_{1}$-saturated algebraically closed valued field $(K,v)$. In particular, unlike the rest of the paper to this point, our topology now has a uniformly definable basis. This means that we can definably speak about germs of functions at a point (because the equivalence relation of having the same germ is definable). We will do this throughout.
Otherwise, we use very little from the theory ACVF, so we do not elaborate here. Interested readers may refer to [HHM06, §2.1], [HK06, §2.1] and references therein for more background. Section 2.1 of the more recent [HOP24] contains also some background (and relevant references) on the analytic theory of functions definable in complete models of ACVF that is needed in a couple of points below.
10.1. Taylor Groupoids
We will work with approximate versions of TIMI, successively generalizing until we get the full statement. To streamline the presentation, we work with ind-definable sets. For us, an $A$-ind-definable set is a countable union of $A$-definable sets; a set is ind-definable if it is $A$-ind-definable for some $A$; and a function is ($A$)-ind-definable if its graph is. Thus, an $A$-ind-definable group is a group whose underlying set, composition operation, and inverse operation, are all $A$-indefinable.
We want to work more generally with ind-definable groupoids:
Definition 10.1.
A $\emptyset$-ind-definable groupoid is a groupoid $\mathcal{C}$ such that (i) the set of objects of $\mathcal{C}$, (ii) the set of morphisms of $\mathcal{C}$, (iii) the composition operation on morphisms of $\mathcal{C}$, and (iv) the inverse operation on morphisms of $\mathcal{C}$, are all $\emptyset$-ind-definable.
Example 10.2.
Define the groupoid $\mathcal{LDH}$ (local definable homeomorphisms) whose objects are the elements of $K$, and whose morphisms $x\rightarrow y$ are germs of definable homeomorphisms between neighborhoods of $x$ and $y$. Then $\mathcal{LDH}$ is a $\emptyset$-ind-definable groupoid.
Our first goal is to introduce, in the language of ind-definable groupoids, an abstract notion of Taylor series.
Definition 10.3.
A weak Taylor groupoid consists of the following data and requirements:
  1. (1)
    A $\emptyset$-ind-definable groupoid $\mathcal{C}$, which is a sub-groupoid of $\mathcal{L}DH$.
  2. (2)
    A $\emptyset$-ind-definable group $G$.
  3. (3)
    A $\emptyset$-ind-definable homomorphism $a$ from morphisms in $\mathcal{C}$ to $G$. This means that if $g\circ f=h$ as morphisms in $\mathcal{C}$, then $a(g)a(f)=a(h)$ in $G$.
  4. (4)
    For each $n\geq 1$, a $\emptyset$-ind-definable map $f\mapsto c_{n}(f)$ from morphisms in $\mathcal{C}$ to $K$, such that $c_{1}(f)\neq 0$ for all $f$.
For ease of notation, we often denote a weak Taylor groupoid by the underlying groupoid $\mathcal{C}$, omitting the additional data.
Definition 10.4.
Let $\mathcal{C}$ be a weak Taylor groupoid, and let $f:x\rightarrow y$ be a morphism in $\mathcal{C}$. For each $n\geq 1$, we define the $n$th truncation of $f$, dented $T_{n}(f)$, to be the tuple $(x,y,a(f),c_{1}(f),...,c_{n}(f))$.
Definition 10.5.
Let $\mathcal{C}$ be a weak Taylor groupoid. Then $\mathcal{C}$ is a Taylor groupoid if the following hold:
  1. (1)
    If $g\circ f=h$ as morphisms in $\mathcal{C}$, then for each $n$, the $n$th truncation of $h$ is determined by the $n$th truncations of $f$ and $g$.
  2. (2)
    If $f$ is a morphism in $\mathcal{C}$, then $f$ is determined by its sequence of $n$th truncations for $n\geq 1$.
  3. (3)
    For fixed $n\geq 1$, every $(x,y,a,c_{1},...,c_{n})\in K^{2}\times G\times K^{n}$ with $c_{1}\neq 0$ is the $n$th truncation of a morphism in $\mathcal{C}$.
Taylor groupoids are very close to asserting Taylor’s Theorem about a certain class of functions in an analytic context: each function is definably assigned a countable sequence of ‘coefficients’ at each point, which determine the function in a neighborhood of the point, and these coefficients carry definable composition and inverse operations which are well-defined on finite truncations. The main difference, and the reason for the term ‘generalized’, is the group $G$. One should think of elements of $G$ as powers of the Frobenius map, which are necessary to cover all the functions we need in positive characteristic.
Since the various properties required in the two notions are quite similar, it is not hard to see that a ‘big enough’ Taylor groupoid implies that $(K,v)$ has definable slopes:
Lemma 10.6.
Let $\mathcal{C}$ be a Taylor groupoid. Assume that the set of objects of $\mathcal{C}$ is $K$, and that every basic invertible arc (in the sense of Definition 7.5) is a morphism in $\mathcal{C}$. Then $(K,v)$ has definable slopes.
Proof.
Definition 10.5(1) implies that for each $n$, multiplication and inverse in $\mathcal{C}$ induce a groupoid structure on the $n$th truncations of all morphisms in $\mathcal{C}$. One then defines $n$-slopes as $n$th truncations, and the axioms of definable slopes follow easily. We point out that the uniform definability of slopes, and the definability of composition and inverse of $n$-slopes, follow from the $\emptyset$-ind-definability of truncation, composition, and inverse in $\mathcal{C}$; while the fact that $\mathcal{C}$ is the inverse limit of the truncation categories follows from compactness and Definition 10.5 (2) and (3). ∎
Note that the groupoids and functors constructed in Lemma 10.6 are canonically determined by $\mathcal{C}$. Thus we define:
Definition 10.7.
If $\mathcal{C}$ is a Taylor groupoid as in Lemma 10.6, we will call the system of groupoids and truncation maps constructed in Lemma 10.6 the definable slopes on $(K,v)$ induced by $\mathcal{C}$.
10.2. Etale Functions at 0
Our goal is to construct increasingly large Taylor groupoids until we reach one as in Lemma 10.6. We start with étale functions at $(0,0)$.
Definition 10.8.
Let $f:0\rightarrow 0$ be a morphism in $\mathcal{LDH}$ (i.e. a germ of a definable homeomorphism of neighborhoods of 0).
  1. (1)
    We say that $f$ is a basic étale function at $(0,0)$ if there is a polynomial $P(x,y)\in K[x,y]$ such that $P(0,0)=0$, both partial derivatives of $P$ are non-zero at $(0,0)$, and $f$ is given by the restriction of $P(x,y)=0$ to a neighborhood of $(0,0)$.
  2. (2)
    We say that $f$ is an étale function at $(0,0)$ if it is a composition of finitely many basic étale functions at $(0,0)$.
It is evident from the definition that étale functions are closed under composition and inverse, and thus define a $\emptyset$-ind-definable groupoid with 0 as its only object.
Notation 10.9.
We let $\mathcal{EFO}$ (étale functions at the origin) denote the $\emptyset$-ind-definable groupoid of étale functions at $(0,0)$.
Our first goal is to extend $\mathcal{EFO}$ to a Taylor groupoid, with the group $G$ interpreted as the trivial group.
Lemma 10.10.
Let $f$ be an étale function, and fix $n\geq 1$. Then there is exactly one polynomial $P(x)\in K[x]$ of degree at most $n$ such that $\lim_{x\rightarrow 0}\frac{P(x)-f(x)}{x^{n}}=0$.
Proof.
The lemma can be expressed using infinitely many first order sentences. In particular, it suffices to carry out the proof in a complete model. So let $\mathbb{K}$ be a complete model of the same characteristics as $K$. It follows from the implicit function theorem that every basic étale function at $(0,0)$ is analytic, i.e. is given by a convergent power series in a neighborhood of 0. Thus, so is $f$. Write $f(x)=\sum_{i=1}^{\infty}c_{i}x^{i}$. Then the desired polynomial $P$ is $\sum_{i=1}^{n}c_{i}x^{i}$. ∎
Notation 10.11.
Let $f$ be an étale function at $(0,0)$, and $n\geq 1$. We define the $n$th coefficient of $f$, denoted $c_{n}(f)$, to be the coefficient of $x^{n}$ in the polynomial provided in Lemma 10.10.
We now show:
Proposition 10.12.
With the maps $c_{n}$ defined above, the groupoid $\mathcal{EFO}$ becomes a Taylor groupoid.
Proof.
It is clear from the statement of Lemma 10.10 that the maps $c_{n}$ are $\emptyset$-ind-definable, so (interpreting the group $G$ as the trivial group) we obtain a weak Taylor groupoid. We now verify (1)-(3) in Definition 10.5:
  1. (1)
    As in the proof of Lemma 10.10, the statement is first-order, so we can work in a complete model. Now let $g\circ f=h$ be étale functions at $(0,0)$. Then the $n$th truncations of $f,g,h$ are just the truncations of their power series expansions at $(0,0)$. In other words, (1) just says that composition and inverse are well-defined on truncated polynomials. This is well-known and easy to check.
  2. (2)
    Suppose $f$ and $g$ have the same $n$th truncations for all $n$. It follows that, if we let $h=f-g$, then $\lim_{x\rightarrow 0}\frac{h(x)}{x^{n}}=0$ for all $n$. If $h$ is identically 0, we are done. Otherwise, in a small enough ball, we may assume that all fibers of $h$ have size at most some integer $d$, and that $h(x)=0$ only holds for $x=0$. Now let $\Gamma$ be the value group, and for $r\in\Gamma$, define $s(r)$ to be the supremum of all $v(h(x))$ for $x$ with $v(x)=r$. If $s(r)=\infty$ for some $r$, it follows by the Swiss Cheese decomposition that the image under $h$ of the elements of value $r$ contains all non-zero points in some open ball at 0. In particular, since $h$ is at most $d$-to-one, there are only finitely many such $r$. So, passing to a smaller neighborhood of 0, we may assume $s(r)<\infty$ for all $r$. But for each $n$, since $\lim_{x\rightarrow 0}\frac{h(x)}{x^{n}}=0$, it follows that for sufficiently large $r\in\Gamma$ we have $s(r)\geq nr$. In particular, $s:\Gamma\rightarrow\Gamma$ grows faster than any linear function. This contradicts that $\Gamma$ is a pure divisible ordered abelian group.
  3. (3)
    Equivalently, we want to show that every $(c_{1},...,c_{n})\in K^{n}$ with $c_{1}\neq 0$ is the $n$th truncation of an étale function. This is trivial, since $(c_{1},...,c_{n})$ is the $n$th truncation of the polynomial $\sum_{i=1}^{n}c_{i}x^{i}$, which is clearly an étale function since $c_{1}\neq 0$.
10.3. Generalized Etale Functions at 0
We next generalize $\mathcal{EFO}$ to include Frobenius powers.
Definition 10.13.
Let $G$ denote the Frobenius group of maps $K\rightarrow K$, defined as follows: if $K$ has characteristic zero, $G=\{\mathrm{id}\}$. If $K$ has characteristic $p>0$, $G$ is the group of maps $x\mapsto x^{p^{n}}$.
We view $G$ as a $\emptyset$-ind-definable group acting $\emptyset$-ind-definably on $K$. The following are the key properties we need, which are evident from the definition:
Fact 10.14.
For the group $G$ defined above:
  1. (1)
    Every element of $G$ is a $\emptyset$-definable automorphism of $(K,v)$, so in particular also a homeomorphism.
  2. (2)
    If $a\in G$ is an étale function, then $a=\mathrm{id}$.
Definition 10.15.
A generalized étale function at $(0,0)$ is a morphism at $(0,0)$ in $\mathcal{LDH}$ of the form $f\circ a$, where $a\in G$ and $f$ is an étale function at $(0,0)$.
Since $G$ is $\emptyset$-ind-definable, the class of generalized étale functions at $(0,0)$ is also $\emptyset$-ind-definable. We want to show that this class forms a Taylor groupoid in a natural way. Everything will follow from the ensuing three lemmas:
Lemma 10.16.
Each generalized étale function at $(0,0)$ has exactly one expression as $f\circ a$, where $a\in G$ and $f$ is an étale function at $(0,0)$.
Proof.
Assume that $f\circ a=g\circ b$ are two such expressions. Simplifying yields $g^{-1}\circ f=b\circ a^{-1}$. But $g^{-1}\circ f$ is an étale function at $(0,0)$, so then $b\circ a^{-1}\in G$ is also an étale function at $(0,0)$. By Fact 10.14, $b\circ a^{-1}=\mathrm{id}$. Thus $a=b$, and thus $f=g$. ∎
Notation 10.17.
Let $a\in G$, and let $f$ be an étale function at $(0,0)$. By $f^{a}$ we mean the function whose graph is the image of $f$ under the automorphism $a$: that is, the map sending $a(x)$ to $a(y)$ whenever $f(x)=y$.
Lemma 10.18.
Let $f$ be an étale function at $(0,0)$, and let $a\in G$. Then:
  1. (1)
    $f^{a}$ is an étale function at $(0,0)$.
  2. (2)
    The truncations of $f^{a}$ are the images of the truncations of $f$ under $a$.
  3. (3)
    $a\circ f=f^{a}\circ a$.
Proof.
Both (1) and (2) follow from the fact that $a$ is an automorphism, since the class of étale functions at $(0,0)$ and their truncation maps are $\emptyset$-ind-definable. (3) is immediate from the definition of $f^{a}$. ∎
Lemma 10.19.
Let $f$ and $g$ be étale functions at $(0,0)$, and let $a,b\in G$.
  1. (1)
    $(g\circ b)\circ(f\circ a)=(g\circ f^{b})\circ(b\circ a)$.
  2. (2)
    $(f\circ a)^{-1}=f^{a^{-1}}\circ a^{-1}$.
  3. (3)
    In particular, the set of generalized étale functions is closed under composition and inverse, so forms a $\emptyset$-ind-definable groupoid.
Proof.
(1) and (2) follow from Lemma 10.18(3). (3) follows from (1) and (2). ∎
Now let $\mathcal{GEFO}$ be the $\emptyset$-ind-definable groupoid of generalized étale functions at $(0,0)$. If $f\circ a$ is a morphism in $\mathcal{GEFO}$, we define its $n$th truncation to be $(0,0,a,c_{1},...,c_{n})$, where $(0,0,\mathrm{id},c_{1},...,c_{n})$ is the $n$th truncation of $f$. By Lemma 10.16, this is well-defined.
Proposition 10.20.
With the Frobenius group $G$ and the truncation maps defined above, $\mathcal{GEFO}$ is a Taylor groupoid.
Proof.
The $\emptyset$-ind-definability of the truncation maps follows from Lemma 10.16, the $\emptyset$-ind-definability of $G$, and the $\emptyset$-ind-definability of truncation in $\mathcal{EFO}$.
Next, it follows from Lemma 10.19(1) that composition in $\mathcal{GEFO}$ induces composition in $G$. We have now shown that $\mathcal{GEFO}$ is a weak Taylor groupoid. We proceed to verify (1)-(3) in the definition of Taylor groupoids (Definition 10.5:
  1. (1)
    It follows from Lemmas 10.19 and 10.18(2) that composition and inverse are well-defined on truncations.
  2. (2)
    Suppose $f\circ a$ and $g\circ b$ are morphisms in $\mathcal{GEFO}$ with the same $n$th truncations for all $n$. By definition of the truncations in $\mathcal{GEFO}$, this means that $a=b$ and $f$ and $g$ have the same $n$th truncations for all $n$. Since $\mathcal{EFO}$ is a Taylor groupoid, this implies $f=g$.
  3. (3)
    For any $a\in G$ and any $c_{1},...,c_{n}\in K$ with $c_{1}\neq 0$, we want to find a morphism in $\mathcal{GEFO}$ with $n$th truncation $(0,0,a,c_{1},...,c_{n})$. Let $P(x)=\sum_{i=1}^{n}c_{i}x^{i}$. Then the desired truncation is achieved by $P\circ a$.
10.4. Generalized Étale Functions
Finally, we now extend $\mathcal{GEFO}$ to include maps at arbitrary points in $K^{2}$, not just the origin.
Definition 10.21.
Let $(x_{0},y_{0})\in K^{2}$. A generalized étale function at $(x_{0},y_{0})$ is a morphism at $(x_{0},y_{0})$ in $\mathcal{LDH}$ of them form $f(x-x_{0})+y_{0}$, where $f$ is a morphism in $\mathcal{GEFO}$. A generalized étale function is a generalized étale function at some point.
It is clear that each generalized étale function at $(x_{0},y_{0})$ can be expressed as $f(x-x_{0})+y_{0}$ for a unique morphism $f$ from $\mathcal{GEFO}$.
Notation 10.22.
We denote the generalized étale function $f(x-x_{0})+y_{0}$ by $f_{x_{0}}^{y_{0}}$.
The following properties are immediate:
Lemma 10.23.
Let $f_{x}^{y}$ and $g_{y}^{z}$ be generalized étale functions.
  1. (1)
    $g_{y}^{z}\circ f_{x}^{y}=(g\circ f)_{x}^{z}$.
  2. (2)
    $(f_{x}^{y})^{-1}=(f^{-1})_{y}^{x}$.
  3. (3)
    In particular, the generalized étale functions form the morphisms of a $\emptyset$-ind-definable groupoid whose set of objects is $K$.
Notation 10.24.
Let $\mathcal{GEF}$ be the $\emptyset$-definable groupoid of generalized étale functions.
We now point out that $\mathcal{GEFO}$ extends canonically to a Taylor groupoid on $\mathcal{GEF}$.
Definition 10.25.
If $f_{x}^{y}$ is a generalized étale function and $n\geq 1$, we define the $n$th truncation of $f_{x}^{y}$ to be $(x,y,a,c_{1},...,c_{n})$, where $(0,0,a,c_{1},...,c_{n})$ is the $n$th truncation of $f$.
This time, there is really nothing to check. We leave the details to the reader, and just state:
Proposition 10.26.
With the group $G$ and the truncation maps given above, $\mathcal{GEF}$ forms a Taylor groupoid.
We conclude:
Theorem 10.27.
$(K,v)$ has definable slopes, induced as in Lemma 10.6 by the Taylor groupoid $\mathcal{GEF}$.
Proof.
By Lemma 10.6, we only need to show that every basic invertible arc is a morphism in $\mathcal{GEF}$. So, let $X\subset K^{2}$ be $A$-definable of dimension 1, with both projections $X\rightarrow K$ finite-to-one. Let $(x_{0},y_{0})\in K^{2}$ be generic over $A$. By genericity, we may assume that $X$ is defined in a neighborhood of $(x_{0},y_{0})$ by an irreducible polynomial equation $P(x,y)=0$ with coefficients in $\operatorname{acl}(A)$.
Note that the map $x\mapsto x^{p}$ defines a generalized étale function at every point. Indeed, for any $a$, we have $x^{p}=(x-a)^{p}+a^{p}$, which is clearly a generalized étale function at $(a,a^{p})$. In particular, we may freely replace $X$ with its image under applying any Frobenius power to either coordinate. Since the Frobenius is $\emptyset$-definable, this will send $(x_{0},y_{0})$ to a generic point of the resulting set; and since the Frobenius is a homeomorphism, the local behavior of $X$ is preserved.
Now, by the above paragraph, we may assume that $P(x,y)$ does not belong to either $K[x^{p},y]$ or $K[x,y^{p}]$. Let $V$ be the variety defined by $P$. It follows that each projection of $V$ to $K$ induces a separable extension of function fields, which implies that each such projection is generically étale. Since $(x_{0},y_{0})$ is generic in $X$, each projection $V\rightarrow K$ is étale at $(x_{0},y_{0})$. In particular, the partial derivatives of $P$ at $(x_{0},y_{0})$ are both non-zero.
Now, at this point, it is harmless to translate $X$ and assume $x_{0}=y_{0}=0$. Then since both partial derivatives of $P$ are non-zero at $(0,0)$, $P$ defines a basic étale function at $(0,0)$. ∎
10.5. TIMI in ACVF
It remains to show that with respect to the notion of definable slopes we have constructed, ACVF satisfies TIMI. We need some preliminaries. First, we note the following, which is clear:
Lemma 10.28.
Let $P(x,y,\overline{z})$ be a polynomial over $K$ in $m+2$ variables, and fix $(x_{0},y_{0},\overline{z}_{0})$ with $P(x_{0},y_{0},\overline{z}_{0})=0$. Assume the partial derivatives of $P$ with respect to $x$ and $y$ are non-zero at $(x_{0},y_{0},\overline{z}_{0})$. Then the set $\{(x,y):P(x,y,\overline{z}_{0})=0\}\subset K^{2}$ restricts to a generalized étale function $f$ at $(x_{0},y_{0})$, such that $a(f)$ (the associated element of the Frobenius group $G$) is the identity.
Notation 10.29.
Let $P(x,y,\overline{z})$ be a polynomial whose graph near some point $(x_{0},y_{0},\overline{z}_{0})$ gives a generalized étale function at $(x_{0},y_{0})$ in the sense of the above lemma. We denote this generalized étale function by $P\restriction(x_{0},y_{0},\overline{z}_{0})$.
Note that if the partial derivatives are non-zero at $(x_{0},y_{0},\overline{z}_{0})$ as in Lemma 10.28, then they are non-zero in a neighborhood of $(x_{0},y_{0},\overline{z}_{0})$. Thus, the notation $P\restriction(x_{0},y_{0},\overline{z}_{0})$ is well-defined in a neighborhood of $(x_{0},y_{0},\overline{z}_{0})$.
Now our main tool for proving TIMI is the following:
Lemma 10.30.
Let $P(x,y,\overline{z})$ and $Q(x,y,\overline{w})$ be polynomials in $m_{P}+2$ and $m_{Q}+2$ variables, respectively, and let $I\subset K^{2}\times K^{m_{P}}\times K^{m_{Q}}$ be the set of $(x,y,\overline{z},\overline{w})$ with $P(x,y,\overline{z})=Q(x,y,\overline{w})=0$. Fix $(x_{0},y_{0},\overline{z}_{0},\overline{w}_{0})\in I$ so that all partial derivatives of $P$ and $Q$ in the $x$ and $y$ variables are non-zero at $(x_{0},y_{0},\overline{z}_{0},\overline{w}_{0})$. Assume for some $n$ that the $n$th truncations of $P\restriction(x_{0},y_{0},\overline{w}_{0})$ and $Q\restriction(x_{0},y_{0},\overline{z}_{0})$ coincide, but every neighborhood of $(x_{0},y_{0},\overline{z}_{0},\overline{w}_{0})$ contains some $(x,y,\overline{z},\overline{w})$ so that the $n$th truncations of $P\restriction(x,y,\overline{w})$ and $Q\restriction(x,y,\overline{z})$ do not coincide. Then the projection $I\rightarrow K^{m_{P}}\times K^{m_{Q}}$ is not injective in any neighborhood of $(x_{0},y_{0},\overline{z}_{0},\overline{w}_{0})$.
Proof.
The lemma can be expressed by infinitely many first-order sentences, so it suffices to carry out the proof in a complete model. In this case, by the Implicit Function Theorem, the graph of $P(x,y,\overline{z})=0$ is, in a neighborhood of $(x_{0},y_{0},\overline{z}_{0})$, the graph of an analytic function $(x,\overline{z})\mapsto y_{P}(x,\overline{z})$. Similarly, the graph of $Q$ is given locally by an analytic function $y_{Q}(x,\overline{w})$. Let $g$ be the analytic function $(x,\overline{w},\overline{z})\mapsto y_{P}(x,\overline{z})-y_{Q}(x,\overline{w})$. For given $\overline{z},\overline{w}$, we let $g_{\overline{z}\overline{w}}$ be the analytic function in one variable given by restricting $g$ to tuples with the values $\overline{z}$ and $\overline{w}$ in the $K^{m_{P}}$ and $K^{m_{Q}}$ coordinates.
Now the fact that the $n$th truncations of $P\restriction(x_{0},y_{0},\overline{w}_{0})$ and $Q\restriction(x_{0},y_{0},\overline{z}_{0})$ coincide, equivalently stated, gives that $x_{0}$ is a root of multiplicity $n$ of $g_{\overline{z}_{0}\overline{w}_{0}}$. By the continuity of roots (this follows easily, say, from [CW02]), the number of roots of $g_{zw}$ including multiplicity, in any sufficiently small neighborhood $U$ of $(\overline{x}_{0},\overline{y}_{0})$, is constant in a neighborhood of $(\overline{z}_{0},\overline{w}_{0})$. But if $(x,y,\overline{z},\overline{w})\in I$ is sufficiently close to $(x_{0},y_{0},\overline{z}_{0},\overline{w}_{0})$ such that the $n$th truncations do not coincide on $(x,y,\overline{z},\overline{w})$, then $(x,y)$ is a root of multiplicity less than $n$ of $g_{zw}$, implying that $g_{zw}$ has at least one other root in $U$. Such a root corresponds to another point $(x^{\prime},y^{\prime},\overline{z},\overline{w})\in I$, violating local injectivity of $I\rightarrow K^{m_{P}}\times K^{m_{Q}}$. ∎
We now conclude:
Theorem 10.31.
With the definable slopes induced by $\mathcal{GEF}$, $(K,v)$ satisfies TIMI.
Proof.
We are given two $A$-definable families of correspondences in $K$, $X\subset K^{2}\times T$ and $Y\subset K^{2}\times U$, and generic points over $A$, $(x_{0},y_{0},t_{0})\in X$ and $(x_{0},y_{0},u_{0})\in Y$. We assume that the $n$-slopes of $X_{t_{0}}$ and $Y_{u_{0}}$ coincide at $(x_{0},y_{0})$, but that in some neighborhood of $u_{0}$, if $u\models\operatorname{tp}(u_{0}/x_{0}y_{0})$ with $u$ independent from $t_{0}$ over $x_{0}y_{0}$, then the $n$th slopes of $X_{t_{0}}$ and $Y_{u}$ do not coincide at $(x_{0},y_{0})$. We want to show that $(x_{0},y_{0},t_{0},u_{0})$ is a topological ramification point of the projection $I\rightarrow T\times U$, where $I$ is the set of $(x,y,t,u)$ with $(x,y)\in X_{t}\cap Y_{u}$.
Now, $X_{t_{0}}$ and $Y_{u_{0}}$ define generalized étale functions at $(x_{0},y_{0})$. It is harmless to apply any Frobenius power to any copy of $K$, since the Frobenius is $\emptyset$-definable (so doesn’t change which points are generic), is a homeomorphism (so doesn’t change the topological ramification locus), and is a generalized étale function at every point (so doesn’t change the relation of two curves having the same $n$-slope). In particular, we may assume the Frobenius map associated to the generalized étale functions $X_{t_{0}}$ and $Y_{u_{0}}$ is the identity.
It is also harmless to restrict to any $A$-definable neighborhood of $(x_{0},y_{0},t_{0},u_{0})$, and to replace $T$ and $U$ by any sets $A$-definably homeomorphic to $T$ and $U$. In particular, since $t_{0}\in T$ and $u_{0}\in U$ are generic, we may assume $T$ and $U$ are open subsets of powers of $K$.
Now by the genericity of $(x_{0},y_{0},t_{0})\in X$ and $(x_{0},y_{0},u_{0})\in Y$, after restricting to small enough neighborhoods we can assume $X$ and $Y$ are defined by polynomial equations $P(x,y,t)=0$ and $P(x,y,u)=0$. Since the generalized étale functions of $X_{t_{0}}$ and $Y_{u_{0}}$ have no Frobenius power, it follows that the partial derivatives of $P$ in $Q$ in the $x$ and $y$ variables are non-zero at $(x_{0},y_{0},t_{0},u_{0})$. It follows that we are in the situation of Lemma 10.30: indeed, by compactness, any neighborhood of $u_{0}$ contains some $u\models\operatorname{tp}(u/x_{0}y_{0})$ with $u$ independent from $t_{0}$ over $x_{0}y_{0}$ – thus, by assumption, $(x_{0},y_{0},t_{0},u_{0})$ belongs to the closure of points $(x,y,t,u)\in I$ where the $n$th slopes of $X_{t}$ and $Y_{u}$ do not coincide. ∎
11. Proofs of the Main Theorems
We now prove our main theorems for definable relics of ACVF. For this, we need to recall the main result of the recent preprint [HOP24] (Fact 11.3 below). We use some ad hoc terminology to recall their setting most easily.
Recall the notion of almost embeddability (Definition 2.1).
Definition 11.1.
Let $(K,v)$ be an $\aleph_{1}$-saturated algebraically closed valued field. Let $(G,\oplus_{G})$ and $(H,\oplus_{H})$ be interpretable groups in $(K,v)$ over some parameter set $A$. Assume that each of $G$ and $H$ is almost embeddable into $K$. We say that $G$ and $H$ are locally equivalent if there are $g_{12},g_{23},g_{13}\in G$ and $h_{12},h_{23},h_{23}\in H$ such that:
  1. (1)
    $\dim(g_{12}g_{23}/A)=\dim(h_{12}h_{23}/A)=2$.
  2. (2)
    $g_{12}\oplus_{G}g_{23}=g_{13}$ and $h_{12}\oplus_{H}h_{23}=h_{13}$.
  3. (3)
    For each $1\leq i<j\leq 3$, $g_{ij}$ and $h_{ij}$ are interalgebraic over $A$.
Remark 11.2.
The definition of local equivalence in [HOP24] is more general. It does not assume almost embeddability into $K$, and uses the more general $dp$-rank (discussed in the next section) instead of dimension. However, we will only encounter groups almost embeddable into $K$; and for these groups, the two notions coincide (see [Sim14, Theorem 0.3], and use [DGL11] to apply it).
Fact 11.3.
[…]
Let us proceed with our main results. First, Theorem 11.4 is a composite of our results on ACVF thus far, which is independent of Fact 11.3:
Theorem 11.4.
Let $\mathcal{K}=(K,v)$ be an $\aleph_{1}$-saturated algebraically closed valued field. Let $\mathcal{M}$ be a non-locally modular strongly minimal definable $\mathcal{K}$-relic. Then $\mathcal{M}$ interprets a strongly minimal group that is almost embeddable into $K$ and locally equivalent to either $(K,+)$ or $(K^{\times},\times)$.
Proof.
By Corollary 9.60, $(K,v)$ is a Hausdorff geometric structure with the open mapping property and ramification purity. Moreover, by Theorems 10.27 and 10.31, $(K,v)$ has definable slopes induced by $\mathcal{GEF}$ and satisfying TIMI. Now apply Theorem 8.1. The result is that $\dim(M)=1$ and $\mathcal{M}$ interprets a strongly minimal group satisfying all of the conclusions of Proposition 7.33. To recall, this gives us the following (where $n_{0}$ is the smallest $n$ such that there is a non-algebraic coherent $n$-slope):
  • A parameter set $B$,
  • An $\mathcal{M}(B)$-interpretable group $(G,\cdot)$ which is strongly minimal as an interpretable set in $\mathcal{M}$,
  • A generic element $y\in K$ over $\emptyset$,
  • For each $i,j\in\{1,2,3\}$ with $i<j$, an element $g_{ij}\in G$, and
  • For each $i,j\in\{1,2,3\}$ with $i<j$, a morphism $f_{ij}:y\rightarrow y$ in $\mathcal{IA}_{n_{0}}$,
such that:
  1. (1)
    $y$ is $\mathcal{K}(B)$-definable (that is, $y\in\operatorname{dcl}_{\mathcal{K}}(B)$).
  2. (2)
    The $(n_{0}-1)$-th truncation of each $f_{ij}$ is the identity at $y$.
  3. (3)
    $g_{12}$ and $g_{23}$ are independent generics in $G$ over $B$.
  4. (4)
    $g_{13}=g_{23}\cdot g_{12}$ and $f_{13}=f_{23}\circ f_{12}$.
  5. (5)
    For each $i,j\in\{1,2,3\}$ with $i<j$, $g_{ij}$ and $f_{ij}$ are interalgebraic over $B$.
Our goal is to show that $(G,\cdot)$ is almost embeddable into $K$ and locally equivalent to either the additive or multiplicative group of $K$. The first clause is easy: By strong minimality, $G$ is in finite correspondence with $M$; but $M$ admits a finite-to-one map to $K$ (by using that $\dim(M)=1$ and taking piecewise projections) – so composing gives a finite correspondence between $G$ and a subset of $K$. Thus $G$ is almost embeddable into $K$.
Now we move toward local equivalence with the additive or multiplicative group. Recalling that our definable slopes in ACVF are induced by the groupoid $\mathcal{GEF}$ from the previous section, we can express each $f_{ij}$ as an $n_{0}$-truncated morphism $y\rightarrow y$ in $\mathcal{GEF}$ with trivial $(n_{0}-1)$-th truncation. Recall that such truncations are notated $(y,y,a,c_{1},...,c_{n_{0}})$, where $a$ is a Frobenius power and $c_{1},...,c_{n_{0}}$ are the coefficients of an $n_{0}$-truncated polynomial sending 0 to 0. Since our $(n_{0}-1)$-th truncations are trivial, all but the last entry in such an expression are determined. Precisely, we have the following options for $f_{ij}$:
  • Case 1: $n_{0}=1$. Then we have
    $f_{ij}=(y,y,\textrm{id},h_{ij})$
    for some $h_{ij}\in K$.
  • Case 2: $n_{0}>1$. Then we have
    $f_{ij}=(y,y,\textrm{id},1,0,0,...,0,h_{ij})$
    for some $h_{ij}\in K$.
The first case corresponds to scaling maps $x\mapsto h_{ij}x$, and the second corresponds to maps of the form $x+h_{ij}x^{n_{0}}$. In any case, we can fix $h_{ij}\in K$ as above for each $i<j$. Since $y\in\operatorname{dcl}_{\mathcal{K}}(B)$, each $h_{ij}$ is interdefinable with $f_{ij}$ over $B$, and so is also interalgebraic with $g_{ij}$ over $B$. One thus easily concludes (1) and (3) from Definition 11.1, interpreting $H$ as either $(K,+)$ or $(K^{\times},\times)$ (and replacing $A$ from Definition 11.1 with our $B$). So to complete the proof, we must show that either $h_{13}=h_{12}+h_{23}$ or $h_{13}=h_{12}h_{23}$. But $f_{13}=f_{12}\circ f_{23}$, and composition of truncations in $\mathcal{GEF}$ corresponds to composition of truncated polynomials – so we just need to check the composition operation on the truncated polynomials in the two cases above:
  • Case 1: $n_{0}=1$. Then $h_{13}x$ is the composition of $h_{12}x$ and $h_{23}x$, which is $h_{12}h_{23}x$, and thus $h_{13}=h_{12}h_{23}$.
  • Case 1: $n_{0}=2$. Then $x+h_{13}x^{n_{0}}$ is the $n_{0}$-truncated composition of $x+h_{12}x^{n_{0}}$ and $x+h_{23}x^{n_{0}}$. One easily checks this truncated composition to be $x+(h_{12}+h_{23})^{n_{0}}$, and thus $h_{13}=h_{12}+h_{23}$.
Now combining our work with Fact 11.3, we obtain the full restricted trichotomy for definable relics of ACVF:
Theorem 11.5.
Let $\mathcal{K}=(K,v)$ be an algebraically closed valued field. Let $\mathcal{M}$ be a non-locally modular strongly minimal definable $\mathcal{K}$-relic. Then $\mathcal{M}$ interprets a field $F$, $(K,v)$-definably isomorphic to $K$. Moreover, the structure induced on $F$ from $\mathcal{M}$ is a pure algebraically closed field.
Proof.
First, we check that we can assume $(K,v)$ is $\aleph_{1}$-saturated. This is a straightforward reduction, and similar arguments are given in [Cas23, Theorem 9.6] and [HS24, Lemma 4.8]. We give a sketch. Let $(K_{1},v_{1})$ be an $\aleph_{1}$-saturated elementary extension of $(K,v)$. Then the same formulas defining $\mathcal{M}$ in $(K,v)$ also define a definable $(K_{1},v_{1})$-relic, say $\mathcal{M}_{1}$. It is easy to see that $\mathcal{M}_{1}$ is an elementary extension of $\mathcal{M}$, so is strongly minimal and not locally modular.
Now suppose $\mathcal{M}_{1}$ interprets a field definably isomorphic to $K_{1}$, say $F$. Then there are formulas (with parameters from $K_{1}$) defining the set $F$, the field structure on $F$, the interpretation of $F$ in $\mathcal{M}_{1}$, and a field isomorphism $F\leftrightarrow K_{1}$. But all of these properties are first-order, so they can be pulled down to $(K,v)$ and $\mathcal{M}$. Thus, using different instances of the same formulas, one can find a definably isomorphic copy of $K$ in $\mathcal{M}$.
So we assume $(K,v)$ is $\aleph_{1}$-saturated. By Theorem 11.4, $\mathcal{M}$ interprets a strongly minimal group $G$ locally equivalent to $(K,+)$ or $(K^{\times},\times)$. Now apply Fact 11.3 to the structure induced on $G$ from $\mathcal{M}$. ∎
Finally, we deduce a full solution of the restricted trichotomy for pure algebraically closed fields:
Corollary 11.6 (Restricted Trichotomy Conjecture).
Let $K$ be an algebraically closed field, and let $\mathcal{M}$ be a strongly minimal $K$-relic. If $\mathcal{M}$ is not locally modular, then $\mathcal{M}$ interprets the field $K$.
Proof.
By elimination of imaginaries, we may assume $\mathcal{M}$ is a definable $K$-relic (that is, its universe is a subset of some $K^{n}$). By an identical argument to Theorem 11.5 above, we may assume $K$ is $\aleph_{1}$-saturated (really, $\aleph_{0}$ is enough here), and thus there is a non-trivial valuation on $K$. Let $v$ be such a valuation. Then $\mathcal{M}$ is still interpreted in $(K,v)$. One can now finish in two ways: either apply Theorem 11.5 directly (thus using [HOP24]); or apply Theorem 8.1 to show $\dim(M)=1$, and then conclude with the main result of [HS24]. ∎
12. Imaginaries in ACVF
In the final section, we discuss strongly minimal structures interpreted, rather than defined, in ACVF. That is, we allow the universe of the structure to be in an imaginary sort. Our main result is a reduction to the real sorts in residue characteristic zero, which works in residue characteristic $p$ assuming a conjectural condition about the distinguished sort $K/\mathcal{O}$ which we expect to be true.
We follow the strategy used in [HHP22] for classifying interpretable fields in various valued fields by reducing them to four distinguished sorts (the valued field, the residue field, the value group, and the set of closed balls of radius zero). The guiding principle is that because of the ‘richness’ and ‘uniformity’ provided by families of plane curves in non-locally modular strongly minimal structures, such a structure should be (1) only able to ‘live’ in one of these four sorts, and (2) not able to live in the two ‘non-field’ sorts. We thus hope to reduce to structures genuinely coming from either the valued field or the residue field.
12.1. The Setting
Throughout this section, $\mathcal{K}=(K,v)$ is an $\aleph_{1}$-saturated model of ACVF. We let $\Gamma$ denote the value group, $\mathcal{O}$ the valuation ring, and k the residue field. $\mathcal{M}=(M,...)$ is a non-locally modular strongly minimal $\mathcal{K}$-relic, whose universe $M$ need not be definable. Absorbing parameters into $(K,v)$ and $\mathcal{M}$, we assume that $\operatorname{acl}_{\mathcal{M}}(\emptyset)$ is infinite, and that every $\emptyset$-definable set in $\mathcal{M}$ is $\emptyset$-definable in $(K,v)$.
Convention 12.1.
Recall (see Definition 2.1) that a finite correspondence between two interpretable sets $X$ and $Y$ is an interpretable set $C\subset X\times Y$ such that the projections of $C$ to both factors are finite-to-one and surjective.
12.2. Ranks
It was shown in [DGL11] that ACVF is dp-minimal, and it thus follows that any ACVF relic has finite dp-rank (if the universe of the relic is $X/E$ for some definable set $X\subseteq K^{n}$ and definable equivalence relation $E$ then $\operatorname{dprk}(X^{n}/E)<n$). See [HHP22, §2.1] for more relevant details on dp-rank. We remind the readers that dp-rank (as opposed to dimension) is sub-additive, but need not be additive:
$\operatorname{dprk}(ab/A)\leq\operatorname{dprk}(a/A)+\operatorname{dprk}(b/Aa).$
We use $\operatorname{dprk}$ for dp-rank in $(K,v)$, and $\operatorname{MR}$ for Morley rank in $\mathcal{M}$. Since $M$ is infinite, $\operatorname{dprk}(M)\geq 1$.
We will repeatedly use the following:
Lemma 12.2.
Let $X\subset M$ be infinite and interpretable over $A$. Let $a=(a_{1},...,a_{n})\in X^{n}$ with $\operatorname{dprk}(a/A)=n\cdot\operatorname{dprk}(X)$. Then $a$ is $\mathcal{M}$-generic in $M^{n}$ over $A$.
Proof.
If not, then some $a_{m}\in\operatorname{acl}_{\mathcal{M}}(Aa_{1}...a_{m-1})$. But then for that $m$ we have
$\operatorname{dprk}(a_{m}/Aa_{1}...a_{m-1})=0,$
and so sub-additivity implies that
$\operatorname{dprk}(a/A)\leq\sum_{j=1}^{n}\operatorname{dprk}(a_{j}/Aa_{1}...a_{j-1})\leq 0+(n-1)\cdot\operatorname{dprk}(X)<n\cdot\operatorname{dprk}(X),$
a contradiction. ∎
It follows from the main result of [Joh22] that finiteness is definable in ACVFeq. In such situations, dp-rank is also definable in definable families, but this is not needed in the sequel.
12.3. Local Interalgebraicity
In [HHP22], Halevi, Hasson, and Peterzil classify the interpretable fields in various valued fields of interest. Their main tool is the notion of an interpretable set $Y$ being locally almost strongly internal to an interpretable set $D$: this means that some infinite interpretable subset of $Y$ admits an interpretable finite-to-one map to $D$. In particular, the authors showed that every interpretable set is locally almost strongly internal to one of four distinguished sets (the valued field, the value group, the residue field, and the set of closed balls of radius zero). They then classify the fields locally almost strongly internal to each of these four sorts.
We will use a very similar strategy. However, for our purposes, it is more convenient to work with finite correspondences instead of finite-to-one maps (roughly, because without assuming a field, one cannot always generate enough definable functions). Thus, we will introduce a nearly identical, but slightly weaker, notion. Note that the reduction to the four distinguished sorts will remain valid.
Definition 12.3.
Given interpretable sets $Y$ and $D$, we define the $D$-critical number of $Y$, $\operatorname{Crit}_{D}(Y)$, to be the largest dp rank of an interpretable subset $X\subset Y$ which is in interpretable finite correspondence with a subset of some $D^{n}$. We moreover call any such $X\subset Y$ $D$-critical in $Y$. Finally, we say that $Y$ is locally interalgebraic with $D$ if $\operatorname{Crit}_{D}(Y)\geq 1$.
By compactness, we have:
Fact 12.4.
Let $Y$ and $D$ be interpretable, and let $X$ be $D$-critical in $Y$. Let $A$ be a parameter set, let $y\in Y$, and let $d_{1},...,d_{n}\in D$. If $y$ is interalgebraic with $d_{1},...,d_{n}$ over $A$, then $\operatorname{dprk}(y/A)\leq\operatorname{dprk}(X)$.
The following is also clear:
Fact 12.5.
Suppose $Y$, $Z$, and $D$ are interpretable sets, and $Y$ is in finite correspondence with a subset of $Z$. Then $\operatorname{Crit}_{D}(Y)\leq\operatorname{Crit}_{D}(Z)$.
The following key observation is [HHP22, Proposition 5.5], and forms the basis for our approach:
Fact 12.6.
Every infinite interpretable set is locally interalgebraic with at least one of $K$, k, $\Gamma$, and $K/\mathcal{O}$.
We will need the following stronger result (see [HOP24, Lemma 2.8]):
Fact 12.7.
Let $Y$ be an infinite interpretable set which is not locally interalgebraic with any of k, $\Gamma$, or $K/\mathcal{O}$. Then $Y$ definably embeds into $K^{n}$ for some $n$.
12.4. Richness
Our main goal is to ‘rule out’ the two non-field sorts $\Gamma$ and $K/\mathcal{O}$, by showing in some sense, that they are not complex enough to admit a non-locally modular strongly minimal structure. Our tool for showing this is richness. Roughly, we want to say that a sort $D$ is ‘rich’ if it admits arbitrarily large families of infinite subsets of a fixed $D^{n}$, parametrized by tuples from $D$, satisfying an appropriate ‘non-redundancy’ condition.
Definition 12.8.
Let $A$ be a parameter set, let $k$ be a non-negative integer, and $a$ and $b$ finite tuples. We call $(a,b,A)$ a $k$-rich configuration if the following hold:
  1. (1)
    $a\notin\operatorname{acl}(Ab)$.
  2. (2)
    Whenever $\overline{b}=\{b_{1},b_{2},...\}$ is an infinite sequence of distinct realizations of $\operatorname{tp}_{(K,v)}(b/Aa)$, then $a\in\operatorname{acl}(A\overline{b})$.
  3. (3)
    $\operatorname{dprk}(b/A)\geq k$.
Definition 12.9.
An interpretable set $D$ is rich if there is $n$ such that for arbitrarily large $k$, there are $m$, elements $a\in D^{n}$ and $b\in D^{m}$, and a parameter set $A$, such that $(a,b,A)$ is a $k$-rich configuration.
Example 12.10.
The motivating example of the above notions is given by a family of plane curves in our strongly minimal relic $\mathcal{M}$. Namely, let $\{C_{t}:t\in T\}$ be an almost faithful $\mathcal{M}(A)$-definable family of plane curves in $\mathcal{M}$, where $T\subset M^{k}$ has Morley rank $k$. Let $t\in T$ with $\operatorname{dprk}(t/A)=\operatorname{dprk}(T)$, and let $x\in C_{t}$ with $\operatorname{dprk}(x/At)=\operatorname{dprk}(C_{t})$. Then one shows easily that $(x,t,A)$ is a $k$-rich configuration. In particular, since $\mathcal{M}$ is not locally modular (and thus $k$ can be arbitrarily large), it follows that the universe $M$ is rich (where in the notation of Definition 12.9, we use $n=2$).
The notion of a $k$-rich configuration may seem a bit obscure. Its main benefit to us is that it is preserved under finite correspondences – a useful feature given the definition of local interalgebraicity:
Lemma 12.11.
Assume that $(a,b,A)$ is a $k$-rich configuration, and let $c$ and $d$ be finite tuples with $\operatorname{acl}(Aa)=\operatorname{acl}(Ac)$ and $\operatorname{acl}(Ab)=\operatorname{acl}(Ad)$. Then $(c,d,A)$ is a $k$-rich configuration.
Proof.
Interalgebraicity gives automatically that $\operatorname{dprk}(d/A)\geq k$ and $c\notin\operatorname{acl}(Ad)$. Now suppose $\overline{d}=(d_{1},d_{2},...)$ are infinitely many realizations of $\operatorname{tp}_{(K,v)}(d/Ac)$. For each $i$, let $(a_{i},b_{i})$ be such that $\operatorname{tp}(a_{i}b_{i}cd_{i}/A)=\operatorname{tp}(abcd/A)$. Since $a\in\operatorname{acl}(Ac)$, we may assume after passing to a subsequence that each $a_{i}=a_{1}$. Since each $d_{i}\in\operatorname{acl}(Ab_{i})$, and the $d_{i}$ are distinct, we may assume after passing to a further subsequence that the $b_{i}$ are all distinct (here we are using that all of the tuples involved are finite). Now since $\operatorname{tp}(a_{1}b_{1}/A)=\operatorname{tp}(ab/A)$, $(a_{1},b_{1},A)$ is a $k$-rich configuration. It follows that $a_{1}\in\operatorname{acl}(Ab_{2}b_{3}...)$. But each $b_{i}$ is interalgebraic with $d_{i}$ over $A$, and $a_{1}$ is interalgebraic with $c$ over $A$. Thus $c\in\operatorname{acl}(Ad_{2},d_{3},...)$. ∎
Corollary 12.12.
Let $X$ and $Y$ be interpretable sets. Assume that $X$ is in definable finite correspondence with a subset of some $Y^{n}$. If $X$ is rich, then so is $Y$.
12.5. Main Proposition
We now give the main technical result of this section. This result should be thought of as saying that if $M$ has any interaction at all with some sort $D$, then it has a large amount of interaction with $D$.
Proposition 12.13.
Suppose $M$ is locally interalgebraic with an interpretable set $D$. Assume that $X$ is $D$-critical, and that $X$, as well as some finite correspondence witnessing it, are $A$-definable. Let $\mathcal{C}=\{C_{t}:t\in T\}$ be an $\mathcal{M}(A$)-definable almost faithful family of plane curves with $T\subset M^{2}$ of Morley rank 2. Let $(u_{0},v_{0})\in M^{2}$ be $\mathcal{M}$-generic over $A$, and let $(x_{0},y_{0})\in X^{2}$ be such that $\operatorname{dprk}(x_{0}y_{0}/Au_{0}v_{0})=2\operatorname{dprk}(X)$. Then:
  1. (1)
    There is at least one, but only finitely many, $t\in T$ with $(u_{0},v_{0}),(x_{0},y_{0})\in C_{t}$.
  2. (2)
    If $(u_{0},v_{0}),(x_{0},y_{0})\in C_{t}$ for some $t$ then $C_{t}\cap X^{2}$ is infinite.
Proof.
It follows from Lemma 12.2 that $(x_{0},y_{0})$ is $\mathcal{M}$-generic in $M^{2}$ over $Au_{0}v_{0}$, so that $(u_{0},v_{0})$ and $(x_{0},v_{0})$ are $\mathcal{M}$-independent $\mathcal{M}$-generics in $M^{2}$ over $A$. Since $M^{2}$ is stationary, this easily implies (1).
Now we show (2). Absorbing parameters, we assume $A=\emptyset$. Let $t$ be such that $(u_{0},v_{0}),(x_{0},y_{0})\in C_{t}$. By (1), $t\in\operatorname{acl}(u_{0}v_{0}x_{0}y_{0})$. Now assume toward a contradiction that $C_{t}\cap X^{2}$ is finite. Then $(x_{0},y_{0})\in\operatorname{acl}(t)$, so $(x_{0},y_{0})$ is interalgebraic with $t$ over $(u_{0},v_{0})$. Now $t$ belongs to the set of $t^{\prime}$ with $(u_{0},v_{0})\in C_{t^{\prime}}$, so it follows that $\operatorname{MR}(t/u_{0}v_{0})\leq 1$. In particular, by strong minimality, $t$ is interalgebraic with an element of $M$ over $(u_{0},v_{0})$. But since $x_{0},y_{0}\in X$, $(x_{0},y_{0})$ is interalgebraic with a tuple $\bar{d}$ of elements of $D$. Putting everything together, the tuples $t$, $(x_{0},y_{0})$, and $\bar{d}$ are interalgebraic over $(u_{0},v_{0})$. Since $X$ is $D$-critical, Lemma 12.4 now implies that $\operatorname{dprk}(x_{0}y_{0}/u_{0}v_{0})\leq\operatorname{dprk}(X)$. But by assumption $\operatorname{dprk}(x_{0}y_{0}/u_{0}v_{0})=2\cdot\operatorname{dprk}(X)$, forcing $\operatorname{dprk}(X)=0$, and thus contradicting that $M$ is locally interalgebraic with $D$. ∎
12.6. Corollaries of the Main Proposition
We now deduce two useful corollaries from Proposition 12.13.
Corollary 12.14.
Suppose $M$ is locally interalgebraic with an interpretable set $D$. Then $D$ is rich.
Proof.
Let $X\subset M$ be $D$-critical. We show that $X$ is rich, which implies by Corollary 12.12 that $D$ is rich.
Fix $k$. We will find a $k$-rich configuration $(a,b,A)$ where $a\in X^{2}$ and $b$ is a tuple from $X$. First let $j=k+2$. Then let $\mathcal{C}=\{C_{t}:t\in T\}$ be an almost faithful family of plane curves in $\mathcal{M}$, where $T\subset M^{j+2}$ is generic (see [Cas23], Fact 2.27 and Lemma 2.42). Adding parameters, we assume $\mathcal{C}$ is $\mathcal{M}$-definable over $\emptyset$. Given $z\in M^{j}$, we let $\mathcal{C}(z)$ be the family with graph $\{(x,y,s)\in M\times M\times M^{2}:(x,y)\in C_{(s,z)}\}$. If $z$ is $\mathcal{M}$-generic in $M^{j}$, then $\mathcal{C}(z)$ is an almost faithful family indexed by a generic subset of $M^{2}$.
Let us fix $z\in X^{j}$ with $\operatorname{dprk}(z)=j\cdot\operatorname{dprk}(X)$. It follows from Lemma 12.2 that $z$ is $\mathcal{M}$-generic in $M^{j}$.
Now let $(u,v)\in X^{2}$ with $\operatorname{dprk}(uv/z)=2\operatorname{dprk}(X)$, and let $(x,y)\in X^{2}$ with $\operatorname{dprk}(xy/uvz)=2\operatorname{dprk}(X)$. It follows from Lemma 12.2 that $(u,v)$ is $\mathcal{M}$-generic in $M^{2}$ over $z$.
We are in the situation of Proposition 12.13. We conclude that there is $s\in M^{2}$ such that for $t=(s,z)$, we have $(u,v),(x,y)\in C_{t}$ and $C_{t}\cap X^{2}$ is infinite. Now since $C_{t}\cap X^{2}$ is infinite, there is $a=(a_{1},a_{2})\in C_{t}\cap X^{2}$ with $\operatorname{dprk}(a/uvxyt)\geq 1$. We conclude:
Claim 12.15.
$(a,t,ux)$ is a $k$-rich configuration.
Proof.
Since $\operatorname{dprk}(a/uvxyt)\geq 1$, $a\notin\operatorname{acl}(uxt)$. Now $\operatorname{dprk}(t)\geq\operatorname{dprk}(z)=j\cdot\operatorname{dprk}(X)$. Since $(u,x)\in X^{2}$, $\operatorname{dprk}(ux)\leq 2\operatorname{dprk}(X)$. So by subadditivity (remember that $j=k+2$), $\operatorname{dprk}(t/ux)\geq k\cdot\operatorname{dprk}(X)\geq k$.
Finally, let $\bar{t}=t_{1},t_{2},...$ be distinct realizations of $\operatorname{tp}_{(K,v)}(t/uxa)$. Since $\mathcal{C}$ is almost faithful, there is $N$ such that $C_{t_{1}}\cap C_{t_{N}}$ is finite, which shows that $a\in\operatorname{acl}(ux\bar{t})$. ∎
To prove the corollary, we need to modify $(a,t,ux)$ by replacing $t$ by a tuple from $X$. The key to doing this is:
Claim 12.16.
$t$ is interalgebraic with $zvy$ over $ux$. In particular, $(a,zvy,ux)$ is a $k$-rich configuration.
Proof.
Clearly $z\in\operatorname{acl}(t)$ because $z$ is a subtuple of $t$. Now it follows from the non-triviality of $C_{t}$ that $v\in\operatorname{acl}(tu)$ and $y\in\operatorname{acl}(tx)$. On the other hand, by Proposition 12.13, there are only finitely many $C_{t^{\prime}}$ containing both $(u,v)$ and $(x,y)$, which shows that $t\in\operatorname{acl}(uvxy)$. The claim now follows by Lemma 12.11. ∎
Finally, note that $a\in X^{2}$ and $zvy\in X^{j+2}$. So since $k$ was arbitrary, we conclude that $X$ is rich, and thus so is $D$. ∎
Remark 12.17.
The above proof of Corollary 12.14 would have been easier if, in the definition of richness of $X$, the ‘parameter’ tuple $b$ was not required to be a sequence of elements of $X$. Indeed, in this case we could have taken $(a,t,\emptyset)$ as our $k$-rich configuration. The problem is that our ultimate goal is to show that $\Gamma$ and $K/\mathcal{O}$ are not rich; and if we adopted this weaker version of richness, then $K/\mathcal{O}$ would be rich in positive characteristic (indeed, taking $\mathcal{O}$-linear combinations of Frobenius powers gives arbitrarily large families of subsets of $(K/\mathcal{O})^{2}$ indexed by powers of $\mathcal{O}$). Recall, however, that $K/\mathcal{O}$ is not stably embedded – and we expect that it is impossible to find such large families in $(K/\mathcal{O})^{n}$ without using external parameters.
Corollary 12.18.
Suppose $M$ is locally interalgebraic with an interpretable set $D$. Then $M$ is almost internal to $D$.
Proof.
Let $X\subset M$ be $D$-critical. Fix a rank 2 family $\mathcal{C}=\{C_{t}:t\in T\}$ of non-trivial plane curves as in Proposition 12.13, and assume it is $\emptyset$-definable in $\mathcal{M}$. Now let $v$ be any element of $M-\operatorname{acl}(0)$. Let $u\in X$ be such that $\operatorname{dprk}(u/v)=\operatorname{dprk}(X)$, and let $(x,y)\in X^{2}$ be such that $\operatorname{dprk}(xy/uv)=2\operatorname{dprk}(X)$. So $(u,v)$ is $\mathcal{M}$-generic in $M^{2}$. Then by Proposition 12.13, there is $t\in T$ with $(u,v),(x,y)\in C_{t}$ and $C_{t}\cap X^{2}$ infinite. Let $(x^{\prime},y^{\prime})\in C_{t}\cap X^{2}$ with $\operatorname{dprk}(x^{\prime}y^{\prime}/txy)\geq 1$. So $(x^{\prime},y^{\prime})$ is $\mathcal{M}$-generic in $C_{t}$ over $txy$, and thus there are only finitely many curves from $\mathcal{C}$ containing both $(x^{\prime},y^{\prime})$ and $(x,y)$ (see [Cas23], Lemma 2.38). Since $C_{t}$ is one such curve, $t\in\operatorname{acl}(xyx^{\prime}x^{\prime})$. But since $C_{t}$ is non-trivial, we also have $v\in\operatorname{acl}(tu)$. Thus in total we have $v\in\operatorname{acl}(uxyx^{\prime}y^{\prime})$. Now $u,x,y,x^{\prime},y^{\prime}\in X$. By compactness, all but finitely many elements of $M$ are in $\operatorname{acl}(X)$. Adding finitely many parameters to cover the remaining elements, we get that $M$ is almost internal to $X$. But by definition $X$ is almost internal to $D$, so it follows that $M$ is almost internal to $D$. ∎
12.7. Ruling out the Non-field Sorts
Our next goal is to show that $\Gamma$ and $K/\mathcal{O}$ are not rich. For $K/\mathcal{O}$, our approach will only work in residue characteristic 0.
Let $(D,+,...)$ be either $\Gamma$ or $K/\mathcal{O}$. We can equip $D$ with a non-discrete group topology: for $\Gamma$, we use the order topology, and for $K/\mathcal{O}$, we use the topology generated by balls of negative radius (equivalently, the topology induced by valuation on $K/\mathcal{O}$, identifying $v(\mathcal{O})$ with 0). It was noted in [HHP22] that, with this topology, $D$ forms a uniform structure in the sense of Simon and Walsberg [SW19] (the o-minimal case appears already in the work of Simon and Walsberg; for $K/\mathcal{O}$ see [HHP22, Lemma 5.13]) We do not define this notion here. Instead, we give the following facts that we will use:
Fact 12.19.
Let $X\subset D^{n}$ be $A$-definable with $\operatorname{dprk}(X)=k$. Let $a\in X$ with $\operatorname{dprk}(a/A)=k$. Then there are a projection $\pi:X\rightarrow D^{k}$, a set $B\supset A$, and a $B$-neighborhood $U$ of $a$ in $X$, such that $\operatorname{dprk}(a/B)=k$ and $\pi$ restricts on $U$ to a homeomorphism with an open subset of $D^{k}$.
Fact 12.20.
Let $X\subset D^{n}$ be $A$-definable, and let $a\in X$ be an isolated point. Then $a\in\operatorname{acl}(A)$.
For the first of the above facts see [SW19, Proposition 4.6] (and [HHP22, Corollary 3.13] to ensure $\operatorname{dprk}(a/B)=k$), and for the second see [SW19, Lemma 3.6].
The following notion was studied in [HHP22]:
Definition 12.21.
$D$ is locally linear if whenever $f:U\rightarrow D$ is an $A$-definable function on an open subset of $D^{n}$, and $a\in U$ with $\operatorname{dprk}(a/A)=n$, there is a neighborhood of $a$ on which $f$ agrees with a map of the form $x\mapsto g(x)+b$, where $g:D^{n}\rightarrow D$ is a definable group homomorphism and $b\in D$ is a constant.
One can quickly improve on this definition:
Lemma 12.22.
Assume $D$ is locally linear. Let $X\subset D^{n}$ be $A$-definable, and let $a\in X$ with $\operatorname{dprk}(a/A)=\operatorname{dprk}(X)$. Then there are $B\supset A$ and a $B$-definable neighborhood $U$ of $a$ in $X$, such that $\operatorname{dprk}(a/B)=\operatorname{dprk}(X)$ and $X$ agrees with a coset of a definable subgroup of $D^{n}$.
Proof.
Let $k=\operatorname{dprk}(X)$. By Fact 12.19, we can assume there is a projection $\pi:X\rightarrow D^{k}$ which is a homeomorphism onto an open set. We thus view $X$ locally as the graph of a function $D^{k}\rightarrow D^{n-k}$. Now apply the definition of local linearity to each coordinate component of this function. ∎
The main fact we need is:
Fact 12.23.
$\Gamma$ is locally linear. $K/\mathcal{O}$ is locally linear if $K$ has residue characteristic zero.
For $\Gamma$ this is well known, since $\Gamma$ is an o-minimal pure divisible ordered abelian group. For $K/\mathcal{O}$ this is [HHP22, Proposition 6.24].
We do not know if $K/\mathcal{O}$ is locally linear in residue characteristic $p$, but we suspect that it is.
Continuing, we now show:
Proposition 12.24.
Assume $D$ is locally linear. Then $D$ is not rich.
Proof.
Let $a\in D^{n}$, $b\in D^{m}$, $A$, and $k$ be such that $(a,b,A)$ is a $k$-rich configuration (for now, we do not assume anything about $k$). Let $r=\operatorname{dprk}(ab/A)$, and let $X\subset D^{n}\times D^{m}$ be $A$-definable with $(a,b)\in X$ and $\operatorname{dprk}(ab/A)=r$. Since $(a,b,A)$ is a $k$-rich configuration, and using compactness, we can assume that the intersection of any infinitely many distinct fibers $X_{b^{\prime}}\subset D^{n}$ is finite.
By Lemma 12.22, there are $B\supset A$, a $B$-definable neighborhood $U$ of $a$ in $X$, and a subgroup $H\leq D^{n+m}$, such that $\operatorname{dprk}(ab/B)=r$ and $X$ agrees with the coset $H+(a,b)$ on $U$. Note, then, that whenever $(a^{\prime},b^{\prime})\in U\cap X$, we have $H+(a,b)=H+(a^{\prime},b^{\prime})$, and thus $X$ also agrees with $H+(a^{\prime},b^{\prime})$ on $U$.
Now let $G\times\{0\}$ be the kernel of the projection $H\rightarrow D^{m}$ – so $G$ is a subgroup of $D^{n}$. By the above, whenever $(a^{\prime},b^{\prime})\in U\cap X$, the fiber $X_{b^{\prime}}$ agrees with $G+a^{\prime}$ in a neighborhood of $a^{\prime}$. Indeed, in some small enough neighborhood of $a^{\prime}$, we have
$x\in X_{b^{\prime}}\iff(x,b^{\prime})\in X\iff(x,b^{\prime})\in H+(a^{\prime},b^{\prime})\iff(x,b^{\prime})-(a^{\prime},b^{\prime})\in H$
$\iff(x-a^{\prime},0)\in H\iff x-a^{\prime}\in G\iff x\in G+a^{\prime}.$
Now in particular, this implies that for any $b,b^{\prime}$ with $(a,b),(a,b^{\prime})\in X\cap U$, the fibers $X_{b}$ and $X_{b^{\prime}}$ agree in a neighborhood of $a$ (because they both agree with $G+a$).
Now assume toward a contradiction that $D$ is rich. This says that, in the situation constructed above, we can make $\operatorname{dprk}(ab/A)=\operatorname{dprk}(ab/B)$ arbitrarily large, while leaving $\operatorname{dprk}(a/A)$ (and thus also $\operatorname{dprk}(a/B)$) bounded. In particular, we may fix all of the data in such a way that $b\notin\operatorname{acl}(Ba)$. Thus, we can find $\bar{b}=b_{1},b_{2},...$, an infinite sequence of distinct realizations of $\operatorname{tp}_{(K,v)}(b/Ba)$. So each $(a,b_{i})\in U$, and thus (by the previous paragraph) any finitely many of the fibers $X_{b_{i}}$ coincide on a neighborhood of $a$.
Finally, since $(a,b,A)$ is a $k$-rich configuration, we have $a\notin\operatorname{acl}(Ab)$. So by Fact 12.20, $a$ is not isolated in $X_{b}$, and thus $a$ is not isolated in any $X_{b_{i}}$. Thus, any finitely many $X_{b_{i}}$ have infinite intersection. By compactness, the intersection of all of the $X_{b_{i}}$ is infinite, a contradiction. ∎
Corollary 12.25.
$M$ is not locally interalgebraic with $\Gamma$. If k has characteristic zero, $M$ is not locally interalgebraic with $K/\mathcal{O}$. If k has characteristic $p>0$ and $K/\mathcal{O}$ is locally linear, $M$ is not locally interalgebraic with $K/\mathcal{O}$.
12.8. Concluding
We now show our main result:
Theorem 12.26.
Assume that $K$ has residue characteristic 0, or that k has characteristic $p>0$ and $K/\mathcal{O}$ is locally linear. Then $M$ definably embeds into a power of $K$ or k.
Proof.
By Corollary 12.25, $M$ is not locally interalgebraic with $\Gamma$ or $K/\mathcal{O}$.
First, suppose $M$ is locally interalgebraic with $k$. Then by Corollary 12.18, $M$ is almost internal to k. It follows by an easy type-counting argument that $M$ is stable and stably embedded. But all such sets are internal to k ([HHM06], Proposition 3.4.11). Combining with stable embeddedness and elimination of imaginaries in k, we then embed $M$ into some $k^{n}$.
Now assume $M$ is not locally interalgebraic with k. Then by Fact 12.7, $M$ definably embeds into a power of $K$. ∎
We now drop the ambient assumptions and end the paper by collecting all results on ACVF:
Theorem 12.27.
Let $\mathcal{K}=(K,v)$ be an algebraically closed valued field, and let $\mathcal{M}$ be a non-locally modular strongly minimal $\mathcal{K}$-relic. Assume the residue field k has characteristic zero, or $K/\mathcal{O}$ is locally linear. Then $\mathcal{M}$ interprets a field $(F,\oplus,\otimes)$, $(K,v)$-definably isomorphic to either $K$ or k. Moreover, the induced structure on $F$ from $\mathcal{M}$ is a pure algebraically closed field.
Proof.
By Theorem 12.26, we may assume that $M\subset K^{n}$ or $M\subset\textbf{k}^{n}$ for some $n$. If $M\subset\textbf{k}^{n}$, we reduce to Corollary 11.6. If $M\subset K^{n}$, we reduce to Theorem 11.5. ∎
Acknowledgements: This project started while the first and third authors attended the 2021 Thematic Program on Trends in Pure and Applied Model Theory at the Fields Institute. We thank the Fields Institute for their hospitality. We would also like to thank Yatir Halevi for discussions around the model theory of valued fields, Will Johnson for his help with visceral theories, Amnon Yekutily and Ilya Tyomkin for their help with algebro-geometric aspects of the paper, and Martin Hils for pointing out a mistake in an earlier version of the paper.